### v1

Abstract
We formulate the Knapsack Contracts problem—a strategic version of the classic Stochastic Knapsack problem, which builds upon the inherent randomness shared by stochastic optimization and contract design. In this problem, the principal incentivizes agents to perform jobs with stochastic processing times, the realization of which depends on the agents’ efforts.
Algorithmically, we show that Knapsack Contracts can be viewed as Stochastic Knapsack with costs and multi-choice, features that introduce significant new challenges. We identify a crucial and economically meaningful parameter—the Return on Investment (ROI) value. We show that the Inverse of ROI (or IOR for short) precisely characterizes the extent to which the approximation guarantees for Stochastic Knapsack extend to its strategic counterpart.
For IOR of $\alpha$, we develop an algorithm that finds an $O(\alpha)$-approximation policy that does not rely on adaptivity. We establish matching $\Omega(\alpha)$ lower bounds, both on the adaptivity gap, and on what can be achieved without full distributional knowledge of the processing times. Taken together, our results show that IOR is fundamental to understanding the complexity and approximability of Knapsack Contracts, and bounding it is both necessary and sufficient for achieving non-trivial approximation guarantees. Our results highlight the computational challenges arising when stochasticity in optimization problems is controlled by strategic effort.
1 Introduction
The study of how incentives reshape classic algorithmic problems has been at the heart of algorithmic game theory research since its inception. In their seminal work, Nisan and Ronen [24] famously conjectured that adding incentives for truthfulness to the optimization problem of scheduling would make the problem extremely hard to approximate. Two decades later, intensive research has culminated in a proof of this conjecture [9]. Many additional works explore the intersection of algorithm design with truthfulness constraints [27, 25, 10, 6, 8, 22, 23].
Over recent years, the algorithmic game theory community has begun to explore incentive constraints beyond truthfulness. A recent focus is on an agent’s earnestness in exerting effort to complete a task, where the task is delegated to the agent by a principal. The main economic tool for incentivizing earnest effort is performance-based payment schemes, a.k.a. contracts.
Contracts are designed to create incentives that counteract moral hazard: When an agent exerts effort to carry out a job on behalf of a principal, this leads to a stochastic outcome that only partially reflects the agent’s effort. This stochasticity opens the possibility for the agent to claim they worked hard despite exerting a low level of effort, by blaming poor outcomes on “bad luck”. This phenomenon is known as moral hazard. A contract uses outcome-based payments and distributional knowledge to incentivize earnest effort, thus mitigating moral hazard. The goal of this work is to initiate the study of the intersection of algorithm design with contractual incentives for earnest effort.
1.1 Problem Formulation
To achieve our goal, we seek an algorithmic optimization problem with a natural connection to contractual incentives—similar to the connection between scheduling and truthfulness [24]. Our starting point is the observation that the field of contract design has a natural affinity to stochastic optimization, due to the nondeterminism inherent to both. Building upon this connection, we formulate a model of Knapsack Contracts, in which jobs that need to fit into a given time frame are delegated to agents. In our model, the agents have costs for carrying out the jobs, and the completion times (a.k.a. job lengths or sizes) are a stochastic outcome of the level of effort. The latter makes our model a generalization of Stochastic Knapsack [11]. We show that the agents’ costs lead to incentive constraints that make the problem significantly harder than classic Stochastic Knapsack. Since Stochastic Knapsack has a central place in stochastic optimization [3, 19, 20, 18, 28, 5, 2, 21], we view its study through the prism of contract design as a potential gateway to deeper exploration of the intersection between the two areas.
Example 1.
A PI (principal) wishes to run a set of experiments on the cloud before the deadline of a top AI conference. Each experimental task has a fixed value for the paper, which does not depend on how fast the cloud providers (agents) finish running the experiments. The PI does not benefit from experiments that finish running after the submission deadline. Despite this, she still needs to pay the providers for running them. The PI must spend substantial funds from her research grant to incentivize the agents to use powerful but costly resources when processing the experiments, resulting in shorter running times. This leads to an interesting tension between the PI’s benefit from an experiment, and how much they are willing to pay to incentivize an agent to run it quickly. The problem of optimizing the experiments thus has a mixed-sign objective.
Generalizing Example 1, we consider a single principal and multiple agents with jobs that must be executed sequentially. Each agent $i$ is associated with a fixed known value $v_{i}$ to the principal for completing their job before the deadline expires. Moreover, each agent can choose one of $m$ costly actions, and the chosen action $j\in[m]$ determines a distribution $F_{i}^{j}$ over the possible completion times of the job. The principal cannot directly observe the action taken by an agent, and can only inspect the completion times. To incentivize agent $i$, the principal designs a contract $t_{i}$ that maps the completion time $s_{i}$ to a monetary transfer. Each agent $i$ wants to maximize their own utility: the transfer received minus the cost $c_{i}^{j}$ exerted by them when taking action $j$. The principal wants to maximize the profit, which is the sum of values over all jobs completed before the deadline, minus the sum of transfers made to the agents (i.e., a mixed-sign objective). See Section 2 for a formal problem description.
Relation to vanilla contract design.
Usually in contract settings, the value created for the principal by the agent’s effort is stochastic. Our model captures settings in which the principal’s value for the job delegated to the agent is fixed, but the time it takes the agent to complete the job is a random variable, and if the job doesn’t finish on time then the resulting value is zero. This seemingly simple twist on the standard principal-agent model leads to new challenges.
Relation to Stochastic Knapsack.
In the standard Stochastic Knapsack problem, there are $n$ jobs to execute, each with a fixed value but stochastic size. A policy-maker executes jobs sequentially, gaining the value of any job that terminates before the deadline. This process ends once a job terminates after the deadline. The optimal policy for this problem can be highly complex, adaptive, and intractable. Yet, Dean et al. [11] show that there exists a simple non-adaptive policy that approximates, up to a constant factor, the reward of the optimal adaptive policy.
In contrast, in our model, even if a job ends after the deadline, the principal must pay the agent according to the contract, for processing it. As we show in Section 2.2, our model corresponds to a version of Stochastic Knapsack with job costs and multi-choice, and we show this makes a substantial difference relative to [11]. A first indication of the difference arises from the linear programming approach of [11], which they use to establish their main result. In their setting, the integrality gap of the LP is bounded; however, in Appendix A we demonstrate that the LP approach can grossly overshoot the true optimum in our setting. In fact, as we shall see, in our model there is no constant-approximation non-adaptive policy; any constant-approximation algorithm must adaptively choose the next job to execute. Intuitively, adaptivity is necessary to avoid the danger of incurring a very high cost, so high it overshadows any value the algorithm may obtain.
1.2 Our Results and Techniques
Return on Investment.
The example in Appendix A is based on a “hard instance” of Stochastic Knapsack with job costs. We show this instance (and in fact any other hard instance) has the following property: For some of the jobs, their cost almost “covers” their value. Thus, in order to incentivize such jobs, the principal has to pay the agent their cost, and therefore, they are left with a small return on investment (ROI). I.e., the ratio between the obtained profit and the transfer made is small. Such jobs are intuitively problematic—if their cost must be reimbursed by the principal, but their value is not gained since they finish after the deadline, the loss can be substantial. We essentially show that these are the only problematic jobs.
In order to parameterize this, we inspect the Inverse of the ROI (or IOR) of a job (as the ROI becomes small, the IOR grows). We introduce an economically-meaningful parameter $\alpha$, which represents the worst IOR among all jobs—see Definition 2.6. Our main finding is that IOR is precisely the parameter that determines the loss in approximation due to moving from the classic Stochastic Knapsack problem, to a contract design problem subject to incentive constraints:
  • We show that the principal can efficiently and non-adaptively compute a set of contracts, which give an $O(\alpha)$-approximation to the profit of the (intractable and highly adaptive) optimal policy of the principal. See Theorem 3.1.
  • We show the existence of an instance such that any non-adaptive policy of the principal is bounded away by a factor of $\Omega(\alpha)$ in terms of profit from the optimal adaptive policy. This establishes a tight $\Theta(\alpha)$ adaptivity gap for the Knapsack Contracts problem. See Theorem 4.1.
  • We establish an information gap: Consider any policy (adaptive or not) that has limited information about the jobs’ size distributions in the form of finitely many moments, as well as access to the probability of knapsack overflow at every step. No such policy can guarantee an approximation factor better than $O(\alpha)$ with respect to the profit of the optimal policy with full access to the size distributions. See Theorem 5.2.
We now describe our results in more detail as well as the techniques that enable them.
Positive results.
In Section 3, we provide an efficient algorithm that finds a non-adaptive set of contracts (see Section 2 for adaptivity notions), which give an $O(\alpha)$-approximation to the optimal profit the principal can obtain, where $\alpha$ is the IOR parameter (Theorem 3.1). Our algorithm needs to choose in advance the order in which agents are to execute their jobs and which effort level to incentivize each agent. Each effort level induces a different expected transfer needed in order to incentivize it. As a first step, in Lemma 2.1, we reduce the Knapsack Contract problem to a variant of Stochastic Knapsack that we term the Multi-Choice Stochastic Knapsack Problem with Costs (MSKC). In this purely algorithmic problem, each job is assigned one of multiple combinations of cost and size distribution. When a policy attempts to execute a job, it has to pay the cost up front.
Our algorithm formulates an LP (see Eq. (2)) for the MSKC problem, where the weight of each job $i$ and effort level $j$ corresponds to the profit of executing job $i$ at effort level $j$, and the size is the expected size of the job at effort level $j$. In Dean et al. [11], an LP formulation of the problem gives an upper bound on the reward of the optimal adaptive policy, which is approximated by the reward of the best non-adaptive policy. In our case, the LP is still an upper bound on the optimal profit of an adaptive policy, but this upper bound can be a factor $\alpha$ away (see Appendix A for an example). We show that there exists a non-adaptive set of contracts that yields an $O(\alpha)$-approximation to the LP’s value.
The way in which we construct the set of contracts is the following. We first efficiently find a well-structured solution to the LP where each job is taken integrally at up to one effort level, except for at most a single job that can be taken as a convex combination of two effort levels. This is done by taking an arbitrary solution of the LP, and then shifting weights along the Pareto curves of jobs in order to iteratively move jobs to be taken integrally, while keeping the same solution value. Once we obtain the well-structured solution, we either use a standard bang-per-buck ordering of the jobs selected, or take a single job if its expected profit is larger than a predefined threshold. This threshold is set in a way that will (i) prevent a large loss for the bang-per-buck order in case the IOR parameter $\alpha$ has a large value; and (ii) ensure a large profit when there’s a single job that gives such profit, but might have a low probability of being selected when jobs are executed in the bang-per-buck order.
Negative results.
In Section 4 and Section 5 we give further evidence for the role that IOR has in Knapsack Contracts. In Section 4 we inspect the adaptivity gap, considering three types of policies:
  1. 1.
    Fully adaptive policies that can choose the next job to execute (if any) and at which effort level based on the entire history.
  2. 2.
    Stopping-time adaptive policies that chooses a fixed set of contracts (job/effort level pairs) up front at a fixed order, but stops once any future sequence of jobs has a negative expected profit. These types of policies avoid a large loss that the principal can incur due to the transfer component of the contract problem and are used in the algorithm of Section 6.
  3. 3.
    Non-adaptive policies that chooses a fixed set of contracts at a fixed order, and the process terminates either when the last job in the order has been executed, or when there’s a job that is finished after the deadline.
Dean et al. [11] show that in the setting without costs, the reward of the optimal non-adaptive policy is within a factor 4 of the optimal adaptive policy. We show that when introducing costs, the ratio between these can be unbounded. Namely, for every IOR parameter $\alpha>1$, we show:
  • An instance where the ratio between the profit of stopping-time adaptive policies and non-adaptive policies is $\Omega(\alpha)$ (Section 4.1).
  • An instance where the ratio between the profit of fully adaptive policies and stopping-time adaptive policies is $\Omega(\alpha^{1/3})$ (Section 4.2).
Our instances are ones where each job has only a single effort level, demonstrating that the cost aspect by itself leads to a significant change relative to the setting without cost.
Information barrier.
In Section 5 we inspect how much information is needed to get an $O(\alpha)$ approximation in the purely algorithmic Stochastic Knapsack problem where there is an inherent cost for executing jobs (this corresponds to having a single effort level for each agent). In the purely algorithmic setting, our algorithm from Section 3 only uses first moment information (the expectation) of jobs’ sizes, along with the probability each job’s size will exceed the deadline, and obtains an $O(\alpha)$ approximation. We show that this information does not suffice if one wants to obtain a better approximation. Namely, for every finite $k$, there exists an instance where any (possibly adaptive) policy that can only access the $k$ moments of the size distribution, along with the probability that each job will exceed the deadline, is bounded away by a factor of $\Omega(\alpha)$ from the optimal policy (Theorem 5.2).
Bounded-size jobs.
In Section 6, we consider the case where the IOR parameter can be arbitrarily large, but each job’s probability to take more than $1-\delta$ of the length of the time budget is small for $\delta\in(0,1)$. We give an efficient algorithm that finds a stopping-time adaptive set of contracts, which guarantees an $O(1/\delta)$ approximation to the optimal profit of a fully-adaptive set of contracts.
In order to obtain this result, we first consider a variant of the problem where, when a job finishes after the deadline, the principal still obtains its reward, and then the process stops. For this variant, the algorithm from Section 3 gives a constant approximation to the LP value. We then consider running this algorithm while imagining the deadline is $\delta$-fraction of the actual deadline. We show that in this case, having bounded-size jobs is equivalent to obtaining the reward of the first job that finishes after the deadline. Moreover, halving the deadline only causes a factor $\delta$ loss in the objective. Note that (moderate) adaptivity is required for this variant as well, as non-adaptive policies lose a factor that grows with $\alpha$ (see Appendix B).
Ordered-Adaptive Contracts.
Motivated by real-world examples where jobs must be executed in a given order, in Appendix D we study the model where the order is given, but for each job in the order, the principal can decide whether to execute it, and if so, at which effort level. In Theorem D.1 we give a resource augmentation-type result: given a time budget larger by a factor of $1+\epsilon$ from the original time budget, one can compute an adaptive policy efficiently (in polynomial time in the number of jobs, effort levels, and $1/\epsilon$), which obtains a reward larger than the optimal adaptive policy for the original time budget. The policy, similar to [11, 5], defines a dynamic programming solution on carefully discretized rounded-down processing times. This result hints that our problem boils down to ordering the jobs; i.e., one may hope to find a nearly-optimal adaptive policy for our non-ordered model if one can efficiently find a near-optimal ordering of the jobs.
1.3 Related Work
The study of sequential contracts in [14] investigates the problem of incentivizing a single agent to perform multiple jobs in a sequential manner. In contrast, our work considers a sequence of contracts in which the jobs are executed sequentially by different agents, with no interdependence between the jobs. The works of [26, 15, 1, 16] consider contract design subject to a limited monetary budget constraining the transfers from principal to agent. In our work, we study a multi-agent setting under a time budget constraint on the agents’ actions, where time is limited and the principal must incentivize the agents to complete their jobs on time.
Classic contract theory.
Classic contract theory has a long-standing tradition in economics, focusing on incentive design under asymmetric information. A result by Carroll [7] shows that linear contracts are often robust across a range of environments, providing a foundation for the study of simpler contract forms. These classical models emphasize moral hazard, adverse selection, and trade-offs between risk and incentives.
Algorithmic contract theory.
In recent years, a series of studies has emerged that study contracts through the lens of the theory of computation, initiated by [12, 4]. The authors of [12] raise concerns regarding the complexity and unintuitiveness of optimal contracts, while exploring the approximation ratios of simpler contract forms, such as linear contracts. The multi-agent setting was initially explored by Babaioff et al. [4]. A recent survey paper [13] provides an overview of developments in algorithmic contract theory.
2 Model
In Section 2.1 we present our contract setting; we reduce this setting to a purely combinatorial setting in Section 2.2.
2.1 Contract Design Setting
An instance $I$ of the knapsack contracts problem consists of a principal, $n$ agents, and a knapsack of size $S$. Without loss of generality, let $S=1$, where $S$ represents a budget of some limited resource. As our running example, we assume $S$ is the time remaining until a set deadline. Each agent $i$ has a job which they can carry out to the benefit of the principal. Agent $i$ is represented by the following:
  • $v_{i}$: The reward which the principal obtains if agent $i$ completes their job before the deadline.
  • $\{(F_{i}^{j},c_{i}^{j})\}_{j\in[m]}$: The possible effort levels of agent ii. [^1] All agents have the same number of effort levels without loss of generality, since we can add “dummy” effort levels that would never be incentivized by the principal. Each effort level $j\in[m]$ determines a distribution $F_{i}^{j}$ over the time it takes the agent to complete the job, as well as a cost $c_{i}^{j}\geq 0$. If the agent chooses to exert effort level $j$, they incur cost $c_{i}^{j}$, and complete their job within time $s_{i}$ drawn (independently) from distribution $F_{i}^{j}$. We assume non-decreasing costs $c_{i}^{1}\leq\dots\leq c_{i}^{m}$. Let $\mathbf{F}_{i}=\left(F_{i}^{1},\ldots,F_{i}^{m}\right)$. Let $\mathrm{supp}(F_{i}^{j})$ denote the support of $F_{i}^{j}$, and $\mathrm{supp}(\mathbf{F}_{i})=\cup_{j\in[m]}\mathrm{supp}(F_{i}^{j}).$ We assume the support of each distribution is finite and is explicitly given.
The principal knows $I=(v_{i},\{F_{i}^{j},c_{i}^{j}\}_{j\in[m]})_{i\in[n]}$ in advance. After an agent completes a job, the principal observes the time it took the agent to complete it, but it does not observe the effort level chosen by the agent. The principal offers each agent $i$ a contract $t_{i}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, mapping $i$’s observed processing time to a transfer paid by the principal to the agent. Given contract $t_{i}$, agent $i$ chooses an effort level
$j\in\arg\max_{\ell}\operatorname{\mathbb{E}}_{s_{i}\sim F_{i}^{\ell}}[t_{i}(s_{i})]-c_{i}^{\ell}.$
Definition 2.1 (Implementable effort level).
For agent $i$, an effort level $j$ is implementable if there exists a contract $t_{i}$ for which this action maximizes the agent’s expected utility among all actions given the contract. In this case, we say contract $t_{i}$ incentivizes action $j$.
Any implementable effort level $j$ for agent $i$ has an optimal contract that incentivizes it, which can be obtained by solving the following LP:
$\min\quad$ $\sum_{s\in\mathrm{supp}(F_{i}^{j})}t_{i}(s)\cdot F_{i}^{j}(s)$
s.t. $\sum_{s\in\mathrm{supp}(F_{i}^{j})}t_{i}(s)\cdot F_{i}^{j}(s)-c_{i}^{j}\geq\sum_{s\in\mathrm{supp}(F_{i}^{j^{\prime}})}t_{i}(s)\cdot F_{i}^{j^{\prime}}(s)-c_{i}^{j^{\prime}}$ $\forall j^{\prime}\in[m]$
$t_{i}(s)\geq t_{i}(s^{\prime})$ $\forall s<s^{\prime}\in\mathrm{supp}(\mathbf{F}_{i})$ (1)
$t_{i}(s)\geq 0$ $\forall s\in\mathrm{supp}(\mathbf{F}_{i})$
Constraint (1) ensures monotonicity of payments. This is required as without it, an agent might have an incentive to hover around after completing the job and get a higher payment.
The principal’s problem can then be described as: when hiring agent $i$, the principal chooses an effort level $j$ to incentivize, and offers the optimal contract that incentivizes that effort level. The principal’s collected reward is the sum of the rewards of the jobs that the agents get to complete with a total processing time less than the budget $S$. The principal aims to maximize their profit: the expected total reward minus the expected sum of transfers made to the agents.
Adaptive versus non-adaptive policies.
Definition 2.2.
(Fully-Adaptive Policies). A fully-adaptive policy, $\varphi:2^{I}\times[0,1]\rightarrow[n]\times[m]\cup\{\perp\}$, takes as input the unassigned jobs $J\subseteq I$ and the remaining time budget $S^{\prime}\leq 1$, and chooses which agent to hire next and at which implementable effort level. This procedure is repeated until:
  • an agent finishes their job after the deadline; or
  • the policy decides not to execute another job (outputs $\perp$ as the next job).
Let $\mathrm{profit}(\varphi)$ be a random variable representing the profit obtained by the principal when using policy $\varphi$ on instance $I$. We denote by $\mathrm{ADAPT}(I)=\max_{\varphi}\operatorname{\mathbb{E}}[\mathrm{profit}(\varphi)]$ the optimum expected profit obtained by an adaptive policy for instance $I$. When referring to an algorithm $\mathcal{A}$ that is given an instance $I$ and returns a policy $\varphi_{I}$, we denote by $\mathcal{A}(I)=\mathrm{profit}(\varphi_{I})$ the random variable representing the profit obtained by the principal when using the policy outputted by algorithm $\mathcal{A}$’s policy on instance $I$.
Definition 2.3.
(Non-Adaptive Policies) A non-adaptive policy is an ordering of agents $\pi:[n]\rightarrow[n]\cup\{\perp\}\times[m]$. For some $i\in[n]$, $\pi(i)$ outputs the $i$th selected agent and the effort level the policy incentivizes the agent to exert.
Let $\Pi$ be the set of non-adaptive policies. We denote by $\mathrm{NON\_ADAPT}(I)=\max_{\pi\in\Pi}\operatorname{\mathbb{E}}[\mathrm{profit}(\pi)]$ the optimum expected profit obtained by a nonadaptive policy for instance $I$.
Definition 2.4.
(Stopping-Time Adaptive Policies) Similarly to non-adaptive policies, a stopping-time adaptive policy chooses a non-adaptive ordering of the agents and their effort levels $\pi$. However, it may adaptively choose to stop at any point. Without loss of generality, we assume the policy stops if every remaining sequence of agents yields a negative expected profit.
Let $\Pi^{\prime}$ be the set of stopping-time adaptive policies. $\mathrm{STOP\_ADAPT}(I)=\max_{\pi\in\Pi^{\prime}}\operatorname{\mathbb{E}}[\mathrm{profit}(\pi)]$ is the optimum expected profit obtained by a stopping-time policy for instance $I$.
2.2 Multi-Choice Stochastic Knapsack Problem With Costs
We define a generalization of Stochastic Knapsack, where items have multiple choices of assignment, and each is associated with a size distribution, and a cost. We show that our Knapsack Contracts problem reduces to this problem. This implies that our algorithmic results can be obtained in this setting, and the results will carry through to the contracts setting.
Definition 2.5 (Multi-Choice Stochastic Knapsack Problem With Costs).
An instance of the multi-choice stochastic knapsack problem with costs (or MSKC for short) consists of $n$ items, where each item has $m$ choices, each choice has a different cost and size distribution. That is $I=\{v_{i},\{p_{i}^{j},F_{i}^{j}\}_{j\in[m]}\}_{i\in[n]}.$ An item $i$ can be selected with a single choice $j$. In this case, the cost $p_{i}^{j}$ must be incurred up front, and the item’s size is realized from distribution $F_{i}^{j}$. There is a budget of size $S$ ($S=1$ without loss of generality). A policy selects item-choice pairs sequentially, and terminates either when it decides not to select any item, or when the total size of selected items exceeds the budget. The profit of a policy is the total value of items selected that do not exceed the budget minus the total cost of the selected items, including the cost of an item that exceeds the budget, if one exists.
Adaptive, Stopping-time Adaptive and Non-adaptive policies for the MSKC problem are defined analogously to the way they are defined for the Knapsack Contracts problem. The following lemma (Proof in Appendix F) shows that in order to find an approximately optimal policy for the Knapsack Contracts problem, it suffices to find an approximately optimal policy for the MSKC problem.
Lemma 2.1 (Reduction to Multi-Choice Stochastic Knapsack with Costs).
Given an efficient algorithm that finds a policy that $c$-approximates the optimal policy MSKC problem, one can efficiently find a policy that $c$-approximates the optimal policy for the knapsack contracts problem.
2.3 Inverse of Return on Investment Parameter
Consider an instance of the Knapsack Contracts problem. For a job and implementable effort level pair, $(i,j)$, let
$\tilde{t}_{i}^{j}=\operatorname{\mathbb{E}}_{s\sim F_{i}^{j}}[t_{i}(s)],$
where $t_{i}$ is the contract that minimizes the expected payment while incentivizing effort level $j$, as given by the LP in Definition 2.1. That is, $\tilde{t}_{i}^{j}$ is the minimal expected transfer required in order to incentivize agent $i$ to take action $j$.
The hardness results presented in the paper critically exploit problem instances characterized by pairs (job, action) with high values $v_{i}$ and expected transfers $\tilde{t}_{i}^{j}$, but a comparatively low net profit
$v_{i}\cdot\Pr_{s_{i}^{j}\sim F_{i}^{j}}[s_{i}^{j}\leq 1]-\tilde{t}_{i}^{j}.$
In such cases, a job may yield a small positive profit of $v_{i}-\tilde{t}_{i}^{j}$ when the knapsack is empty, yet incur a large negative penalty of up to $\tilde{t}_{i}^{j}$ when capacity is low. Moreover, this fluctuation can be abrupt: even an $\epsilon$ increase in load may decrease the expected profit to $\tilde{t}_{i}^{j}$. We show that this drastic fluctuation property is an inherent obstacle in our model. To do so, we define the Inverse of the Return on Investment parameter:
Definition 2.6 (Inverse of ROI parameter).
We define
$\mathrm{ROI}(i,j)\ =\ \frac{v_{i}\cdot\Pr_{s_{i}^{j}\sim F_{i}^{j}}[\,s_{i}^{j}\leq 1\,]\;-\;\tilde{t}_{i}^{j}}{\tilde{t}_{i}^{j}},$
and
$\alpha\;=\;\max_{i,j\in\mathsf{I}_{i}}\frac{1}{\mathrm{ROI}(i,j)}\,$,
where $\mathsf{I}_{i}$ is the set of implementable actions for job $i$. We sometimes refer to $\alpha$ as the Inverse of ROI (or IOR) parameter.
Intuitively, a large value of $\alpha$ signifies the presence of at least one job where for some effort level, the transfer required to incentivize it is much larger than the net profit from successfully executing the job before the deadline. We show a $\Theta(\alpha)$ adaptivity gap and an $\Omega(\alpha)$ information gap (Section 5), suggesting that $\alpha$ captures precisely the “hardness” aspect for our problem.
Analogously, we define the IOR parameter for the MSKC problem as in Definition 2.6, but using $p_{i}^{j}$ instead of $\tilde{t}_{i}^{j}$. Given a Knapsack Contracts instance with an $\alpha$ IOR parameter, applying the reduction in Lemma 2.1 yields an MSKC instance with the same $\alpha$ as the analogous IOR parameter. This implies that any approximation as a function of $\alpha$ a policy in the MSKC instance would obtain will imply the same approximation as a function of $\alpha$ in the Stochastic Knapsack instance.
3 $O(\alpha)$-Approximate Non-Adaptive Contracts
In this section, we show how to find a non-adaptive set of contracts $\pi:[n]\to[n]\times[m]\cup\{\perp\}$ that obtains an $O(\alpha)$ approximation to the optimal profit the principal can obtain by any (possibly highly adaptive) policy.
By Lemma 2.1, we focus on finding an $O(\alpha)$-approximate non-adaptive policy for the MSKC problem, which yields an $O(\alpha)$-approximate non-adaptive set of contracts for the Knapsack Contracts problem. Our main theorem is the following:
Theorem 3.1.
There exists a polynomial time algorithm for MSKC that computes for every input $I$ a non-adaptive policy $\pi_{I}:[n]\to[n]\times[m]\cup\{\perp\}$ such that
$\frac{\mathrm{ADAPT}(I)}{\operatorname{\mathbb{E}}[\mathrm{profit}(\pi_{I})]}\leq O(\alpha).$
We first define an LP that both guides our solution and gives an upper bound on the performance of every adaptive policy. For every $i\in[n],j\in[m]$ and for a knapsack of size $t$, denote by $w_{ij}=v_{i}\cdot\Pr_{{s_{i}^{j}\sim F_{i}^{j}}}[s_{i}^{j}\leq t]-p_{i}^{j}$ and $\mu_{ij}=\operatorname{\mathbb{E}}_{s_{i}^{j}\sim F_{i}^{j}}[\min\{s_{i}^{j},t\}]$. The first parameter can be viewed as a proxy to the profit from job $i$ when executed at effort level $j$ and the second parameter is the expected truncated size.
Consider the following LP, parameterized by the size of the knapsack $t$.
$\Phi(t)=\max\left\{\sum_{ij}w_{ij}\,x_{ij}\;:\;\sum_{ij}\mu_{ij}\,x_{ij}\,\leq\,t,\quad\forall\,i,\;\sum_{j}x_{ij}\,=\,1,\quad x_{ij}\,\geq\,0\right\}.$ (2)
We assume, without loss of generality, that every item $i$ includes a dummy “null” choice $j_{\bot}$ with zero size and zero value; selecting this choice represents the case where none of the actual choices for the item are chosen. In addition, we will use the solution to this LP to guide our algorithm. First, we characterize the structure of optimal solutions:
Lemma 3.2.
There exists an optimal solution to $\Phi(t)$ satisfying the following properties:
  1. 1.
    At most one item $i^{*}$ has exactly two choices $j,j^{\prime}$ such that $x_{i^{*}j},x_{i^{*}j^{\prime}}>0$.
  2. 2.
    Every other item $i\neq i^{*}$ has at most one choice $j$ with $x_{ij}=1$.
Moreover, such a solution can be computed in polynomial time.
Note that every solution returned by the Simplex algorithm satisfies the above properties; the basic feasible solution of the LP has at most $n+1$ non-negative variables, and the constraints ensure that for each item $i$, there is at least one choice $j$ with $x_{ij}>0$, hence at most one item has two non-zero choices. However, the Simplex algorithm may run at exponential time in the worst case. Thus, we describe a polynomial time algorithm to compute such a solution.
First, for each item $i$, we characterize the choices $j$ for which there exists an optimal solution with $x_{ij}=0$. We then show that there always exists an optimal solution in which at most two choices have positive values.
Definition 3.1 (Single Dominance).
For an item $i$, choice $j^{\prime}$ single dominates choice $j$ if:
$(1)$ $\quad w_{ij^{\prime}}\geq w_{ij}$
$(2)$ $\quad\mu_{ij^{\prime}}\leq\mu_{ij}$
Definition 3.2 (Double Dominance).
For an item $i$, choices $j^{\prime}$ and $j^{\prime\prime}$ double dominate choice $j$ if:
$(1)$ $\quad w_{ij^{\prime}}<w_{ij}<w_{ij^{\prime\prime}}$
$(2)$ $\quad\mu_{ij^{\prime}}\leq\mu_{ij}\leq\mu_{ij^{\prime\prime}}$
$(3)$ $\quad(\mu_{ij}-\mu_{ij^{\prime}})/(w_{ij}-w_{ij^{\prime}})\geq(\mu_{ij^{\prime\prime}}-\mu_{ij})/(w_{ij^{\prime\prime}}-w_{ij})$
Finally, we say that a choice $j$ is dominated if it is either single-dominated or double-dominated.
Lemma 3.3.
There exists an optimal solution $x$ to $\Phi(t)$ in which every dominated choice is assigned zero weight; namely, for any item $i$ and any choice $j$ that is either single‑dominated or double‑dominated, we have $x_{ij}=0$.
Proof.
Suppose that for some feasible solution $x$, we have $x_{ij}\neq 0$.
First, if $j^{\prime}$ single-dominates $j$, define a new solution by setting
$x^{*}_{ij^{\prime}}=x_{ij^{\prime}}+x_{ij},\quad x^{*}_{ij}=0~.$
This yields a feasible solution whose objective value is at least as large as that of $x$. Hence, there exists an optimal solution in which $x_{ij}=0$.
Second, if $j^{\prime}$ and $j^{\prime\prime}$ double-dominate $j$, define a new solution by setting
$x^{*}_{ij^{\prime}}=x_{ij^{\prime}}\;+\;x_{ij}\cdot\frac{w_{ij^{\prime\prime}}-w_{ij}}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}},\quad x^{*}_{ij^{\prime\prime}}=x_{ij^{\prime\prime}}\;+\;x_{ij}\cdot\frac{w_{ij}-w_{ij^{\prime}}}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}},\quad x^{*}_{ij}=0~.$
This again yields a feasible solution whose objective value is at least as large as that of $x$ (see Fig. 1 for illustration, and Appendix E for the full proof). Therefore, in either case, there exists an optimal solution in which $x_{ij}=0$.
[…]
Figure 1: Visualization of the removal of double-dominated points. The hollow circles ($j^{\prime},j^{\prime\prime}$) represent two choices for an item. The red point ($j$) is ”double-dominated” because it lies above the line segment representing the convex combination of $j^{\prime}$ and $j^{\prime\prime}$. In the LP relaxation, this dominated point can be replaced by its vertical projection (the green point) onto the segment, which yields a solution that is no worse in objective value and potentially better in terms of feasibility ($\mu$).
Lemma 3.4.
There exists an optimal solution $x$ to $\Phi(t)$ in which, for every item $i$, at most two choices satisfy $x_{ij}>0$.
Proof.
Let $x$ be an optimal solution in which no choice is dominated; such a solution exists by Lemma 3.3. Without loss of generality we may further assume that no additional choices remain and that
$w_{i,j}\;<\;w_{i,j+1}\qquad\text{for every item }i\text{ and every }j\in\{1,\dots,m-1\}.$
Lemma 3.3 then also yields
$\mu_{i,j}\;<\;\mu_{i,j+1}\qquad\text{for all }i\text{ and }j\in\{1,\dots,m-1\}.$
Consider the pairs $\{\bigl{(}w_{ij},\mu_{ij}\bigr{)}\}_{j\in[m]}$ for a fixed index $i$. Connecting consecutive points by straight-line segments (the standard piecewise-linear interpolation) yields a polyline with non-decreasing slopes. Consequently, the boundary is convex.
Suppose for some item $i$ there are strictly more than two choices that satisfy $x_{ij}>0$. Define $\tilde{\mu}=\sum_{j}x_{ij}\mu_{ij}$. Let $j^{\prime}=\arg\max_{j}\bigl{\{}\mu_{ij}\leq\tilde{\mu}\bigr{\}}$ and $j^{\prime\prime}=\arg\min_{j}\bigl{\{}\mu_{ij}\geq\tilde{\mu}\bigr{\}}$. By the definition of $\tilde{\mu}$ and by our assumption that at least two choices satisfy $x_{ij}>0$, such indices $j^{\prime}$ and $j^{\prime\prime}$ always exist. Set $\lambda=\frac{\mu_{ij^{\prime\prime}}-\tilde{\mu}}{\mu_{ij^{\prime\prime}}-\mu_{ij^{\prime}}}$. Consider the solution defined by $x^{*}_{ij^{\prime}}=\lambda$, $x^{*}_{ij^{\prime\prime}}=1-\lambda$, and $x^{*}_{ij}=0$ for all $j\notin\{j^{\prime}j^{\prime\prime}\}$. Note that this construction satisfies $\sum_{j}x^{*}_{ij}\,\mu_{ij}=\tilde{\mu}$. Therefore, it yields a feasible solution. Since the resulting point lies on the boundary of the convex set, its objective value is at least as large as that of the original solution $x$, (see Fig. 2 for illustration).
[…]
Figure 2: Visualization of the set of undominated points, $\Gamma_{i}$. The blue line represents the efficient (Pareto) frontier formed by undominated choices and their convex combinations. An arbitrary combination of choices that lies above this frontier (the red point) is inefficient. It can be strictly improved by projecting it horizontally onto the frontier (the green point), which achieves the same $\mu$ value for a greater $w$ value.
We are now ready to prove Lemma 3.2.
Proof of Lemma 3.2.
By Lemmas 3.3, 3.4, it follows that there exists an optimal solution $x$ where at most two choices have positive values.
Consider a solution where there exist indices $i_{1},i_{2}$ and $j^{\prime}_{1},j^{\prime\prime}_{1},j^{\prime}_{2},j^{\prime\prime}_{2}$ such that $x_{i_{k},j^{\prime}_{k}}>0$ and $x_{i_{k},j^{\prime\prime}_{k}}>0$ for each $k\in\{1,2\}$. Let
$d_{k}=\frac{w_{i_{k},j^{\prime}_{k}}-w_{i_{k},j^{\prime\prime}_{k}}}{\mu_{i_{k},j^{\prime}_{k}}-\mu_{i_{k},j^{\prime\prime}_{k}}}$
be the current slope of the $k$-th item, and assume without loss of generality that $d_{1}\geq d_{2}$. Define
$\tau=\frac{\mu_{i_{1},j^{\prime}_{1}}-\mu_{i_{1},j^{\prime\prime}_{1}}}{\mu_{i_{2},j^{\prime}_{2}}-\mu_{i_{2},j^{\prime\prime}_{2}}}.$
For $\epsilon>0$, set
$x^{*}_{i_{1},j^{\prime}_{1}}=x_{i_{1},j^{\prime}_{1}}+\epsilon,\quad x^{*}_{i_{1},j^{\prime\prime}_{1}}=x_{i_{1},j^{\prime\prime}_{1}}-\epsilon,$
$x^{*}_{i_{2},j^{\prime}_{2}}=x_{i_{2},j^{\prime}_{2}}-\tau\cdot\epsilon,\quad x^{*}_{i_{2},j^{\prime\prime}_{2}}=x_{i_{2},j^{\prime\prime}_{2}}+\tau\cdot\epsilon.$
By construction, for any $\epsilon>0$ we have
$\sum_{j}x_{i_{1},j}\,\mu_{i_{1},j}+\sum_{j}x_{i_{2},j}\,\mu_{i_{2},j}=\sum_{j}x^{*}_{i_{1},j}\,\mu_{i_{1},j}+\sum_{j}x^{*}_{i_{2},j}\,\mu_{i_{2},j}.$
By choosing the smallest $\epsilon$ for which $x^{*}_{i_{1},j^{\prime}_{1}}=1$, or $x^{*}_{i_{2},j^{\prime}_{2}}=0$, or $x^{*}_{i_{1},j^{\prime\prime}_{1}}=0$, or $x^{*}_{i_{2},j^{\prime\prime}_{2}}=1$, we obtain a feasible solution whose objective value is at least as large as that of the original solution $x$, and whose number of items with two positive choices is strictly smaller. By iteratively applying this procedure, we eventually obtain a solution in which at most a single item has two positive choices.
Finally, we observe that all these transformations are applicable in polynomial time. Therefore, given any solution to $\Phi(t)$, it is possible to maintain a solution satisfying the above properties in polynomial time.
[…]
Figure 3: For an optimal solution $x$, items $i_{1}$ and $i_{2}$ each have two choices with nonzero values. The solution for each item is a convex combination of the two choices, corresponding to the red dots in the graphs $\Gamma_{i_{1}}$ and $\Gamma_{i_{2}}$, which lie on the segments connecting the two choices. As the slope of the first item is larger, the modified solution indicated by the dashed arrows illustrates one possible update that maintain feasibility and increases the objective value.
For a set of item choices $\mathcal{S}\subset[n]\times[m]$, denote $size(\mathcal{S})=\sum_{(i,j)\in\mathcal{S}}s_{i}^{j}$ and $\mu(S)=\sum_{(i,j)\in\mathcal{S}}\mu_{ij}$. The following Lemmas are direct extensions of key lemmas from Dean et al. [11], originally established for the classic Stochastic Knapsack problem with a budget of 1. We adapt these lemmas to our knapsack variant with multi-choice and costs, and for a general budget $t$. In this section, we only need the case where $t=1$; however, in Section 6 we use the generalized versions. The proofs are presented in Appendix E for completeness.
Lemma 3.5.
For any set of item choices $\mathcal{S}$:
$\Pr[size(\mathcal{S})<t]\geq 1-\frac{1}{t}\mu(\mathcal{S}).$
Lemma 3.6.
For any adaptive policy for the MSKC problem with budget $t$, let $\mathcal{S}$ be the random set of item choices that the policy inserts, then $\operatorname{\mathbb{E}}[\mu(\mathcal{S})]\leq 2t$.
Lemma 3.7.
For any constant $c\geq 2$ and budget $t$,
$\mathrm{ADAPT}\leq\Phi(ct)\leq c\cdot\Phi(t)~,$
where $\mathrm{ADAPT}$ is the optimal adaptive policy for budget $t$.
We are now ready to present our algorithm, SKC (Stochastic Knapsack with Costs):
Input: Instance II with items $i\in[n]$, choices $j\in[m]$, values $v_{i}$, costs $p_{i}^{j}$, distributions $F_{i}^{j}$
Output: Non-adaptive policy $\pi$
// Step 1: LP Solution
Solve linear program $\Phi(1)$ to obtain optimal solution $x$ with the structure from Lemma 3.2;
Let $\mathcal{C}=\{(i,j)\mid x_{ij}>0\}$ be the candidate item-choice pairs;
// Step 2: Virtual Item Construction
if $\exists$ item $i^{*}$ with two choices $j,j^{\prime}$ such that $x_{i^{*},j},x_{i^{*},j^{\prime}}>0$ then
   Remove $(i^{*},j)$ and $(i^{*},j^{\prime})$ from $\mathcal{C}$;
   Create virtual item $(i^{*},j^{*})$ with:
 $w_{i^{*},j^{*}}\leftarrow w_{i^{*},j}x_{i^{*},j}+w_{i^{*},j^{\prime}}x_{i^{*},j^{\prime}}$
 $\mu_{i^{*},j^{*}}\leftarrow\mu_{i^{*},j}x_{i^{*},j}+\mu_{i^{*},j^{\prime}}x_{i^{*},j^{\prime}}$
   Add $(i^{*},j^{*})$ to $\mathcal{C}$;
 // Randomly select between original choices
   When attempting to insert $(i^{*},j^{*})$, choose $j$ with probability $x_{i^{*},j}$ and $j^{\prime}$ with probability $x_{i^{*},j^{\prime}}$;
end if
// Step 3: Item Insertion with Threshold Check
$w_{\max}\leftarrow\max_{(i,j)\in\mathcal{C}}w_{ij}$
$\mathcal{T}\leftarrow\frac{\Phi(1)}{4(1+\alpha)}$
if $w_{\max}>\mathcal{T}$ then
   Insert item $(i_{\max},j_{\max})=\arg\max_{(i,j)\in\mathcal{C}}w_{ij}$;
 return policy that executes only this item;
end if
else
   Sort items in $\mathcal{C}$ by decreasing density $\frac{w_{ij}}{\mu_{ij}}$;
 return policy that executes items in decreasing density order;
end if
Algorithm 1 SKC (Stochastic Knapsack with Costs)
An overview of the algorithm.
We begin by solving $\Phi(1)$ to obtain an optimal solution $x$ with the properties established in Lemma 3.2: each item $i$ has one choice $j$ such that $x_{ij}=1$, except for possibly one item $i^{*}$ with two choices $j,j^{\prime}$ such that $x_{i^{*}j}+x_{i^{*}j^{\prime}}=1$. Let $\mathcal{C}$ denote the set of item-choice pairs selected by the LP, and let $r=|\mathcal{C}|$. If there exists an item with two positive choices, we remove these two choices and replace them with a single virtual choice; whenever the algorithm attempts to insert this virtual item, it randomly selects between the two original choices with probabilities proportional to their values in $x$. Since each item now has a single (possibly virtual) choice, we refer to each item-choice pair $(i,j)$ by the item $i$. Note that from the properties of our solution $x$, it holds that $\sum_{k=1}^{r}{w_{k}}=\Phi(1)$ and $\sum_{k=1}^{r}{\mu_{k}}\leq 1$. We then define a threshold $\mathcal{T}(\alpha)$ according to which we decide whether to insert the item with the maximum profit $w_{\text{max}}$. If the threshold is not met, we insert the items in decreasing order of their density $\frac{w_{i}}{\mu_{i}}$, until the knapsack is full or no more items remain in the candidate set $\mathcal{C}$. Our threshold determines whether or not following the greedy ordering is worth the risk of overflow compared to the alternative of inserting the best item. The bigger $w_{\text{max}},\alpha$ are, the more likely we are to insert the best item.
Note that this algorithm can be easily de-randomized by choosing the choice among the two choices that maximizes the expected gain.
Claim 3.8.
$\mathbb{E}[\textbf{SKC}]\geq\mathcal{T}=\frac{\Phi(1)}{4(1+\alpha)}$
Proof.
First, if $w_{\max}>\mathcal{T}$, the algorithm attempts only item $(i_{\max},j_{\max})$, gaining an expected net reward of $w_{\max}$. Since $w_{\max}>\mathcal{T}$, we have $\mathbb{E}[\textbf{SKC}]=w_{\max}>\mathcal{T}$, as required.
Now, consider the case $w_{\max}\leq\mathcal{T}$. Define $\textbf{SKC}_{k}$ as the contribution of item $k$ to the algorithm’s profit. It holds that
$\operatorname{\mathbb{E}}[\textbf{SKC}]=\sum_{k=1}^{r}{\operatorname{\mathbb{E}}[\textbf{SKC}_{k}]}~.$
We want to lower bound $\operatorname{\mathbb{E}}[\textbf{SKC}_{k}]$. If item $k$ is inserted and does not overflow, we obtain a reward of $v_{k}-p_{k}\geq w_{k}$. On the other hand, if the item overflows, we pay a cost of $p_{k}$. By the definition of the IOR parameter (Definition 2.6), $\alpha\geq\frac{p_{k}}{v_{k}\Pr[s_{k}\leq 1]-p_{k}}=p_{k}/w_{k}$, which implies that an overflowing job incurs a cost of $p_{k}\leq\alpha w_{k}\leq\alpha w_{\max}$.
We get that
$\operatorname{\mathbb{E}}[\textbf{SKC}_{k}]$ $=(v_{k}-p_{k})\cdot\Pr[k\mbox{ successfully inserted}]-p_{k}\cdot\Pr[k\mbox{ overflows}]$
$\geq(v_{k}-p_{k})\left(1-\sum_{j=1}^{k}{\mu_{j}}\right)-p_{k}\cdot\Pr[k\mbox{ overflows}]$
$\geq w_{k}\left(1-\sum_{j=1}^{k}{\mu_{j}}\right)-\alpha\cdot w_{\max}\cdot\Pr[k\mbox{ overflows}],$ (3)
where the first inequality follows from Lemma 3.5 (with $t=1$), the second inequality follows from the definition of $w_{k}\leq v_{k}-p_{k}$ and the argument above that $p_{k}\leq\alpha\,w_{\max}$.
In the above, this bound is also true for the virtual item $i^{*}$: the item’s expected truncated size is $\mu_{i^{*}}$, and its expected profit conditioned on the event that it was successfully inserted is at least $w_{i^{*}}$.
Summing over all $k\in[r]$, we have that
$\operatorname{\mathbb{E}}[\textbf{SKC}]$ $\geq\sum_{k=1}^{r}{w_{k}\left(1-\sum_{j=1}^{k}{\mu_{j}}\right)}-\alpha w_{\max}\sum_{k=1}^{r}{\Pr[k\mbox{ overflows}]}$
$\geq\Phi(1)-\sum_{k=1}^{r}{w_{k}}\sum_{j=1}^{k}{\mu_{j}}-\alpha w_{\max}$ $(*)$
$=\Phi(1)-\sum_{1\leq j\leq k\leq r}{w_{k}\mu_{j}}-\alpha w_{\max}$
$=\Phi(1)-\sum_{1\leq j<k\leq r}{w_{k}\mu_{j}}-\sum_{k=1}^{r}{w_{k}\mu_{k}}-\alpha w_{\max}$
$\geq\Phi(1)-\frac{1}{2}\sum_{k,j=1}^{r}{w_{k}\mu_{j}}-\sum_{k=1}^{r}{w_{k}\mu_{k}}-\alpha w_{\max}$ $(**)$
$\geq\frac{\Phi(1)}{2}-\sum_{k=1}^{r}{w_{k}\mu_{k}}-\alpha w_{\max}$ $\sum_{k,j=1}^{r}{w_{k}\mu_{j}}\leq\Phi(1)$
$\geq\frac{\Phi(1)}{2}-w_{\max}\cdot\left(\sum_{k=1}^{r}{\mu_{k}}+\alpha\right)$
$\geq\frac{\Phi(1)}{2}-w_{\max}(1+\alpha)$ $\sum_{k=1}^{r}{\mu_{k}}\leq 1$
$\geq\frac{\Phi(1)}{4}\geq\frac{\Phi(1)}{4(1+\alpha)}=\mathcal{T}~.$ $w_{\max}\leq\mathcal{T}=\frac{\Phi(1)}{4(1+\alpha)}.$ (4)
In ()(*) we use the fact that $\sum_{k\in[r]}\Pr[k\mbox{ overflows}]\leq 1$ as at most one item overflows, and $\sum_{k=1}^{r}{w_{k}}=\Phi(1)$. In ()(**) we use the decreasing order of density: for every $j<k$ it holds that $w_{k}\mu_{j}\leq w_{j}\mu_{k}$, and thus $w_{k}\mu_{j}\leq\frac{1}{2}\left(w_{k}\mu_{j}+w_{j}\mu_{k}\right)$.
We are ready to prove the main theorem of this section:
Proof of Theorem 3.1.
From 3.8 and Lemma 3.7, we conclude:
$\frac{\mathrm{ADAPT}}{\operatorname{\mathbb{E}}[\textbf{SKC}]}\leq\frac{2\Phi(1)}{\frac{\Phi(1)}{4(1+\alpha)}}=8(1+\alpha)=O(\alpha)~.$
In our proof, we show that our algorithm finds a policy that approximates the LP value up to a factor $O(\alpha)$. In [11], a similar LP helped guide the solution as its value was a proxy (up to constant factors) of the optimal policy for the stochastic knapsack problem. In our case, this LP provides an upper bound to the optimal policy (see Lemma 3.7), but its value can be larger by a factor $\Omega(\alpha)$ than the profit of the optimal policy (see Appendix A). This shows that we cannot hope for a tighter approximation to the LP’s value.
In Section 4, we show that every non-adaptive policy cannot have a better approximation than $O(\alpha)$ to the optimal adaptive policy (Theorem 4.1).
4 Adaptivity Gaps
In this section we study the power of different policies. In Section 4.1, we start by comparing Non-Adaptive policies (Definition 2.3) to Stopping-time Adaptive policies (Definition 2.4). Theorem 4.1 shows that the profit of a principal that uses the latter can be higher by a factor of $\Omega(\alpha)$ (which is tight according to Theorem 3.1). In Section 4.2 we study the power of Fully-Adaptive policies (Definition 2.2), and show they are more powerful than policies that fix the order of jobs and only decide when to stop processing new jobs. Namely, the gap between Fully-Adaptive policies and Stopping-time Adaptive policies can be as high as $\Omega(\alpha^{1/3})$ (Theorem 4.2). We note that in the instances that give the lower bound, all jobs have a single choice, and therefore, in order to incentivize jobs, we need to pay their cost exactly. Specifically, when determining the IOR, we can substitute the expected payment in Definition 2.6 with the jobs’ costs.
4.1 Stopping-time Adaptive vs. Non-Adaptive $\Omega(\alpha)$ Gap
We show that there exists a $\Omega(\alpha)$ gap between the optimal Non-Adaptive policy and the optimal Stopping-time Adaptive policy. Recall that both models require the algorithm to choose an item ordering at the beginning; the stopping-time adaptive policy can then adaptively choose when to halt, while the non-adaptive policy must attempt all items in the chosen order.
Theorem 4.1.
For every $\alpha$ there exists an instance $I$ of the knapsack-contracts problem such that
$\frac{\mathrm{STOP\_ADAPT}(I)}{\mathrm{NON\_ADAPT}(I)}=\Omega(\alpha).$
Proof.
For a constant $\alpha$, the claim is trivially true. We prove the claim for every $\alpha>3$. We build an instance $I$ as follows. Consider some $\epsilon<\frac{1}{2}$ and any $\gamma\in(0,\tfrac{1}{2})$, where $\epsilon$ will be used to determine $\alpha$. Our instance has $n$ jobs for $n=\lceil\frac{1}{\epsilon}\rceil$. Each job has value $v=2$ and cost $c=2-\epsilon$.
The size distribution of job $i\in[n]$ is
$s_{i}\;=\;\begin{cases}a_{i},&\text{with probability }1-\epsilon,\\[4.0pt] b_{i},&\text{with probability }\epsilon,\end{cases}$
where
$a_{i}=\gamma^{\,n-i+1},\qquad b_{i}=1-\displaystyle\sum_{j=1}^{i-1}a_{j}$
By Definition 2.6, we have
$\alpha=\frac{2-\epsilon}{\epsilon}\;=\;\Theta\!\left(\frac{1}{\epsilon}\right).$
Note that for every $\alpha>3$, we can choose an $\epsilon<1/2$ in order to set the IOR of our instance to be $\alpha$.
Since for every $i$, $\sum_{j=1}^{i-1}a_{j}+b_{i}=1,$ a stopping-time adaptive policy can safely schedule the jobs sequentially until it encounters the first job $i$ whose realized size equals $b_{i}$, and then halt.
We show that since $b_{i}+a_{i^{\prime}}>1$ for every pair $i<i^{\prime}$, attempting to schedule job $i^{\prime}$ without considering the realization of earlier jobs’ size yields negative expected profit. Consequently, the optimal $\mathrm{NON\_ADAPT}$ policy executes at most a single job. Our proof combines these two facts to yield the desired bound.
Formally, consider a stopping‑time adaptive policy that executes jobs in increasing index order ($i=1,2,\ldots$) until it encounters the first job whose size realizes as $s_{i}=b_{i}$, and then stops. Let $X$ denote the number of jobs the policy executes. We have
$\operatorname{\mathbb{E}}[X]\;=\;\sum_{i=1}^{n}\Pr[X\geq i]\;=\;\sum_{i=1}^{n}(1-\epsilon)^{i-1}\;=\;\frac{1-(1-\epsilon)^{n}}{\epsilon}.$
Since $n\geq 1/\epsilon$, we get that $\operatorname{\mathbb{E}}[X]\geq\frac{1-e^{-1}}{\epsilon}$. Therefore the expected profit is at least
$\mathrm{STOP\_ADAPT}(I)\;=\;\operatorname{\mathbb{E}}[X]\,(v-c)\;\geq\;\frac{1-e^{-1}}{\epsilon}\,\cdot\,\epsilon\;=\;1-e^{-1}.$
Consider now a non‑adaptive policy. This policy fixes the number of jobs to execute, $k$, and an ordering $\pi=i_{1},i_{2},\ldots,i_{k}$. For each $\ell\in[k]$, let $\mathrm{profit}(\ell)$ denote the profit from the $\ell$‑th job in the sequence. Clearly, $\operatorname{\mathbb{E}}[\mathrm{profit}(1)]=\epsilon$ for every non‑empty sequence of jobs. We will show that for every sequence with $k>1$ we have $\operatorname{\mathbb{E}}[\mathrm{profit}(k)]<0$, which implies that it is always better to execute at most a single job, since we can iteratively remove jobs and only improve the expected profit of the policy.
Given a sequence $\pi=i_{1},i_{2},\ldots,i_{k}$ of jobs, consider first the case in which $i_{\ell}>i_{k}$ for some earlier job in the sequence ($\ell<k$). If $i_{k}$ is reached and its size realises as $s_{i_{k}}=b_{i_{k}}$, the job terminates after the deadline, because $s_{i_{\ell}}+s_{i_{k}}\geq a_{i_{\ell}}+b_{i_{k}}>1$. Hence
[…]
$\geq\Pr[i_{k}\text{ is reached}]\Bigl{(}\Pr[s_{i_{k}}=b_{i_{k}}](-c)+\Pr[s_{i_{k}}=a_{i_{k}}](v-c)\Bigr{)}$
$=\Pr[i_{k}\text{ is reached}]\bigl{(}\epsilon(-c)+(1-\epsilon)(v-c)\bigr{)}$
$=\Pr[i_{k}\text{ is reached}]\bigl{(}-2\epsilon+\epsilon^{2}+\epsilon-\epsilon^{2}\bigr{)}$
$=-\,\epsilon\,\Pr[i_{k}\text{ is reached}]\;<\;0.$
Now consider the case where $i_{\ell}<i_{k}$ for all $\ell<k$. Let $j=\arg\max_{\ell<k}i_{\ell}$, the largest‑index job among the first k1k-1 scheduled. Note that, $i_{j}<i_{k}$ by our assumption. We have that
$\operatorname{\mathbb{E}}[\mathrm{profit}(k)]$ $=$ $\operatorname{\mathbb{E}}[\mathrm{profit}(k)|i_{k}\mbox{ is reached}]\cdot\Pr[i_{k}\mbox{ is reached}]$
$=$ $\operatorname{\mathbb{E}}[\mathrm{profit}(k)|s_{i_{\ell}}=a_{i_{\ell}}\quad\forall\ell<k]\cdot\Pr[s_{i_{\ell}}=a_{i_{\ell}}\quad\forall\ell<k]$
$+\operatorname{\mathbb{E}}[\mathrm{profit}(k)|i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]\cdot\Pr[i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]$
$=$ $(v-c)\cdot(1-\epsilon)^{k-1}+(-c)\cdot\Pr[i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]$
$=$ $\epsilon\cdot(1-\epsilon)^{k-1}+(-c)\cdot\Pr[i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]$
$\leq$ $\epsilon\cdot(1-\epsilon)^{k-1}+(-c)\cdot\Pr[s_{i_{j}}=b_{i_{j}}\ \wedge\ (s_{i_{\ell}}=a_{i_{\ell}}\quad\forall\ell<k,\ell\neq j)]$
$=$ $\epsilon\cdot(1-\epsilon)^{k-1}+(-2+\epsilon)\cdot(1-\epsilon)^{k-2}\cdot\epsilon$
$=$ $\epsilon\cdot(1-\epsilon)^{k-2}(1-\epsilon-2+\epsilon)$
$=$ $-\epsilon\cdot(1-\epsilon)^{k-2}<0.$
In the above, the third equality follows since $i_{k}$ always overflows in case some $s_{i_{\ell}}=b_{i_{\ell}}$ for some $i_{\ell}<i_{k}$, as $b_{i_{\ell}}+a_{i_{k}}>1$. The first inequality follows from the fact that c-c is negative, and $\Pr[i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]\geq\Pr[s_{i_{j}}=b_{i_{j}}\ \wedge\ (s_{i_{\ell}}=a_{i_{\ell}}\quad\forall\ell<k,\ell\neq j)]$ since for the event $s_{i_{j}}=b_{i_{j}}$ and for all other $\ell<k$, $s_{i_{\ell}}=a_{i_{\ell}},$ we have that $i_{k}$ is reached. Since
$\sum_{\ell<k,\ell\neq j}a_{i_{\ell}}\leq\sum_{\ell=1}^{i_{j}-1}a_{\ell}=1-b_{i_{j}},$
and there will be no overflowing job before $i_{k}$ is executed.
We conclude that the optimal non-adaptive policy only executes a single job, and obtains a profit of $\epsilon$, while the optimal stopping-time adaptive policy obtains a profit of 1, as argued above. Therefore,
$\frac{\mathrm{STOP\_ADAPT}(I)}{\mathrm{NON\_ADAPT}(I)}\geq\frac{1-e^{-1}}{\epsilon}=\Omega(\alpha).$
4.2 Fully-adaptive vs. Stopping-time Adaptive $\Omega(\alpha^{1/3})$ Gap
We further show that there is an $\Omega(\alpha^{1/3})$ gap between the optimal stopping–time adaptive policy and the optimal fully adaptive policy. Recall that the stopping-time adaptive model requires the algorithm to fix an ordering of the items at the outset, and it halts if all future sequences of jobs yield a negative expected profit.
Theorem 4.2.
For every $\alpha$ there exists an instance $I$ of the knapsack-contracts problem such that
$\frac{\mathrm{ADAPT}(I)}{\mathrm{STOP\_ADAPT}(I)}=\Omega\bigl{(}{\alpha}^{1/3}\bigr{)}.$
In order to prove the above theorem, we define an instance $I$ consisting of an infinite collection of jobs, each with with value $v=15/\epsilon$ and cost $c=15/\epsilon-\epsilon^{2}$. As in Section 4.1, $\epsilon$ is used to determine $\alpha$. Again, as the claim is trivial for bounded $\alpha$, we prove the claim for large values of $\alpha$.
The job set is partitioned into $1/\epsilon$ types of jobs $L_{1},\ldots,L_{1/\epsilon}$, each containing an infinite number of jobs. A job of type $L_{i}$ has the size distribution
Probability Size
$1-\epsilon-\epsilon^{2}$ 0
$\epsilon$ $a_{i}$
$\epsilon^{2}$ $b_{i}$
where
$a_{i}=\gamma^{\,n-i+1},\qquad b_{i}=1-\displaystyle\sum_{j=1}^{i-1}a_{j}$
for $\gamma<\tfrac{1}{2}$.
By the definition of the IOR parameter (Definition 2.6),
$\alpha=\frac{15/\epsilon-\epsilon^{2}}{\epsilon^{2}}=15/\epsilon^{3}-1=O\left(\frac{1}{\epsilon^{3}}\right)$
For every large enough $\alpha$, we can choose an appropriate $\epsilon$ that sets the IOR to be $\alpha$.
We first analyze the expected reward from a fully adaptive policy.
Lemma 4.3.
For the instance given above,
$\mathrm{ADAPT}(I)\;\geq\;\frac{1}{2e}.$
Proof.
Consider an adaptive policy that selects jobs by type $L_{1}$, then type $L_{2}$, and so on until type $L_{1/\epsilon}$, according to the following rule:
  • When selecting jobs of type $L_{i}$, the policy continues to select such jobs as long as $s_{i}=0$.
  • Once $s_{i}=a_{i}$, the policy moves on to select jobs of type $L_{i+1}$ if $i<1/\epsilon$.
  • If $s_{i}=b_{i}$, the policy terminates.
Notice that the execution order ensures each executed job finishes within the time budget, so the policy obtains a reward of exactly $v-c=\epsilon^{2}$ for every executed job. Let $X$ be a random variable for the number of jobs the adaptive policy completes. Since the profit from each job is $\epsilon^{2}$, the policy gets a reward of $\operatorname{\mathbb{E}}[X]\cdot\epsilon^{2}$. Let $X_{i}$ be the number of jobs of type $L_{i}$ completed by the policy. Notice that
[…]
By the description of the policy,
$\operatorname{\mathbb{E}}\bigl{[}X_{i}\mid\varphi\text{ executes jobs of type }L_{i}\bigr{]}=\frac{1}{\epsilon+\epsilon^{2}}.$
We also have
[…]
The second inequality follows because in order for the policy to reach jobs of type $L_{1/\epsilon}$, each time a job of type $L_{i}$ (for $i<1/\epsilon$) has its size realized at a value larger than 0, the policy moves on to jobs of type $L_{i+1}$ only if the job’s size is realized at $a_{i}$, which happens with probability $\frac{\epsilon}{\epsilon+\epsilon^{2}}$. Since this must occur at most $1/\epsilon$ times, the probability is at least
$\left(\frac{\epsilon}{\epsilon+\epsilon^{2}}\right)^{1/\epsilon}.$
Therefore,
$\operatorname{\mathbb{E}}[X_{i}]\;\geq\;\frac{1}{e(\epsilon+\epsilon^{2})}\;>\;\frac{1}{2e\epsilon},$
which implies
$\operatorname{\mathbb{E}}[X]=\sum_{i=1}^{1/\epsilon}\operatorname{\mathbb{E}}[X_{i}]\;>\;\frac{1}{2e\epsilon}\cdot\frac{1}{\epsilon}=\frac{1}{2e\epsilon^{2}}.$
This yields
$\mathrm{ADAPT}(I)\;>\;\epsilon^{2}\cdot\frac{1}{2e\epsilon^{2}}\;=\;\frac{1}{2e},$
as desired. ∎
We next bound the expected reward from a stopping-time adaptive policy.
Lemma 4.4.
For the instance given above,
$\mathrm{STOP\_ADAPT}(I)\;\leq\;4\epsilon.$
Proof.
Consider some sequence $\pi$ of jobs. Let $Y$ denote the following event: the policy is to process a job of type $L_{i}$, when some prior job of type $L_{i}$ was processed with a size realized to be larger than 0. We claim that once event $Y$ occurs, then for every sequence of future jobs, the expected profit is negative. Consider first the expected profit of only processing the next job, of type $L_{i}$. This is at most
$(1-\epsilon^{2})(v-c)+\epsilon^{2}(-c)=(1-\epsilon^{2})\cdot\epsilon^{2}-\epsilon^{2}(15/\epsilon-\epsilon^{2})=-15\epsilon+\epsilon^{2}<0,$
since if the next job’s size is realized to be $b_{i}$, then the job finishes after the deadline for sure.
As the profit of any job is at most $v-c=\epsilon^{2},$ this implies that for any sequence to have a positive profit, it has to be of length larger than $(15\epsilon-\epsilon^{2})/\epsilon^{2}$. Consider a partition of the sequence into blocks of size
$k=\frac{1}{\epsilon+\epsilon^{2}}+\frac{1}{\epsilon},$
If the sequence ends at a block of size smaller than $k$, we merge the last two blocks together to form a single block, meaning that the size of each block is between $k$ and $2k$. Denote by $r$ the number of blocks in the sequence. Consider some block $j\in[r]$, and let $n_{j}(i)$ denote the number of jobs of type $L_{i}$ it contains. From our observation, if $n_{j}(i)>1$, then the expected profit of processing the second job of type $L_{i}$ in the block is at most $15\epsilon+\epsilon^{2} \ZoharcommentThis$ is only true given $Y$, right? I think the proof might be more clear if it is partitioned (e.g. having a couple of clearly stated claims in the proof), in which case the expected profit of the entire block is at most
$15\epsilon+\epsilon^{2}+\epsilon^{2}(2k-1)<-15\epsilon+2\epsilon^{2}+4\epsilon=-11\epsilon+2\epsilon^{2}<0.$
Therefore, to have a positive expected profit, for every $i$, if $b_{j}(i)>1$, then the first $b_{j}(i)-1$ jobs of type $L_{i}$ must have size 0, and in total, at least
$\sum_{i=1}^{1/\epsilon}{\max\{n_{j}(i)-1,0\}}\geq k-\frac{1}{\epsilon}=\frac{1}{\epsilon+\epsilon^{2}}$
jobs must have size 0. This occurs with probability at most
$(1-(\epsilon+\epsilon^{2}))^{\frac{1}{\epsilon+\epsilon^{2}}}<\frac{1}{2}$~
in which case the expected profit of the block is at most $2k\epsilon^{2}\leq 4\epsilon$.
Else, with probability at least $1/2$, the expected profit of the block is at most $11\epsilon+2\epsilon^{2}$. Thus, the expected profit of a block bounded given by
$\frac{1}{2}(-11\epsilon+2\epsilon^{2})+4\epsilon=-\frac{3}{2}\epsilon+\epsilon^{2}<0.$
Thus, we get that once event $Y$ occurs, since every sequence of jobs has a negative profit for the policy, it will halt. Let $X$ be a random variable of the number of jobs processed before event $Y$ occurs. Since every job has profit at most $\epsilon^{2}$, we have that $\mathrm{STOP\_ADAPT}(I)\leq\operatorname{\mathbb{E}}[X]\epsilon^{2}$. We again break $\pi$ into blocks of size $k$. As analyzed above, for every block, we have that event $Y$ occurs with probability at least $1/2$; therefore, the expected number of blocks processed is bounded by 2, and the expected number of jobs processed is bounded by $2k$. Therefore,
$\mathrm{STOP\_ADAPT}(I)\leq\operatorname{\mathbb{E}}[X]\epsilon^{2}\leq 2k\epsilon^{2}\leq 4\epsilon,$
as desired. ∎
Combining Lemmas 4.3 and 4.4 yields Theorem 4.2.
5 $\Omega(\alpha)$ Informational Gap
In this section we establish a lower bound that exposes a fundamental information gap in the problem of Stochastic Knapsack with Costs. Previous constant-factor approximation algorithms for Stochastic Knapsack (without costs) have used limited distributional information, specifically the expected (or truncated) sizes and the probability of fitting within the current capacity. In Proposition 5.1 we demonstrate that this information is insufficient. We show that any algorithm, even one that finds an adaptive policy, which relies solely on such limited data, cannot achieve a constant-factor approximation in the presence of costs. More generally, we prove in Theorem 5.2 that even providing the first $k$ moments of each distribution for any constant $k$ is insufficient. This showcases the inherent difficulty of the added costs, implying that future constant-approximation algorithms will need to utilize almost the full distributional structure in novel ways.
5.1 Warm-Up: First Moment
Proposition 5.1.
Let $\mathcal{A}$ be any algorithm whose knowledge of the item size distributions is restricted to the expected values $\operatorname{\mathbb{E}}[s_{i}]$, the truncated expected values $\mu_{i}=\operatorname{\mathbb{E}}[\min\{s_{i},1\}]$, and oracle access to the function $q_{i}(B)=\Pr[s_{i}\leq B]$ for each $i\in[n]$, where $B$ is the current remaining knapsack capacity. For every $\alpha$ there exists an input instance $I$ on which the algorithm’s performance is bounded by:
$\frac{\mathrm{ADAPT}(I)}{\operatorname{\mathbb{E}}[\mathcal{A}(I)]}\geq\Omega(\alpha).$
The proof relies on constructing two types of items, one “good” and one “bad”. The optimal strategy would be to insert only the “good” items, and any algorithm that starts by inserting a “bad” item will result in a solution with a value that is only a $\frac{1}{\alpha}$-fraction of OPT. We design the size distributions to be indistinguishable to algorithm $\mathcal{A}$, forcing $\mathcal{A}$ to arbitrarily pick an item to insert into the knapsack. By creating an instance where the number of “bad” items far exceeds the number of “good” items, algorithm $\mathcal{A}$ will almost surely pick a “bad” item first, resulting in our lower bound.
Proof of Proposition 5.1.
Fix $\epsilon\in(0,1)$. The two item types are defined as follows:
Property Type 1 Item (“good”) Type 2 Item (“bad”)
Reward $v_{i}$ 22 22
Cost $p_{i}$ $2-\epsilon$ $2-\epsilon$
Size $s_{i}$ dist. $s_{1}=\begin{cases}0,&\text{w.p. }1-\epsilon,\\ 1,&\text{w.p. }\epsilon;\end{cases}$ $s_{2}=\begin{cases}\epsilon/2,&\text{w.p. }1-\frac{\epsilon}{2-\epsilon},\\ 1,&\text{w.p. }\frac{\epsilon}{2-\epsilon}.\end{cases}$
Note that in this instance the expected values and the truncated expected values are identical: $\operatorname{\mathbb{E}}[s_{1}]=\operatorname{\mathbb{E}}[s_{2}]=\mu_{1}=\mu_{2}=\epsilon$ and $q_{1}(1)=q_{2}(1)=1$, hence, algorithm $\mathcal{A}$ cannot distinguish between the two items. Inserting one item of type 2 results in a net profit of $\epsilon$, and reduces the remaining capacity $B$ by at least $\epsilon/2$. Inserting any further items (of either types) results in a negative expected profit:
$\operatorname{\mathbb{E}}[\text{inserting item of type $i$}\;|\;B\leq 1-\epsilon/2]$ $\leq\Pr[s_{i}\neq 1]\cdot v_{i}-p_{i}$
$\leq(1-\frac{\epsilon}{2-\epsilon})\cdot 2-(2-\epsilon)$
$=\epsilon(1-\frac{2}{2-\epsilon})\;<\;0.$
Now, consider an input instance $I$ with $n>\frac{1}{\epsilon}$ items of type 1 and $n^{2}-n$ items of type 2. The optimal policy will attempt to insert items of type 1, and stop if an item’s size instantiates $s_{i}=1$, or if no more items of type 1 remain. The expected number of items successfully inserted by OPT is $\frac{1}{\epsilon}$, each item contributing $\epsilon$ net profit, resulting in
$\mathrm{ADAPT}(I)\geq\frac{1}{\epsilon}\cdot\epsilon=1.$
Since $\mathcal{A}$’s only strategy is to randomly select an item to insert, with probability $1-\frac{1}{n}$ it begins by inserting a “bad” item. By taking $n\to\infty$, the expected profit of $\mathcal{A}$ approaches $\epsilon$:
$\lim_{n\to\infty}\operatorname{\mathbb{E}}[\mathcal{A}(I)]=\epsilon.$
Thus for a large enough $n$, the performance ratio of the algorithm can be bounded:
$\frac{\mathrm{ADAPT}(I)}{\operatorname{\mathbb{E}}[\mathcal{A}(I)]}\geq\frac{1}{2\epsilon}\geq\frac{1}{4}\alpha.$
5.2 $k$ Moments
We generalize the previously shown distributional informational gap to show that, even an algorithm with knowledge of the first $k$ (or truncated) moments of the distribution, has an $\Omega(\alpha)$ gap compared to an algorithm with knowledge of the exact distribution.
Theorem 5.2.
Fix $k\geq 1$. Let $\mathcal{A}$ be any algorithm whose knowledge of the item size distributions is restricted to the moment $\operatorname{\mathbb{E}}[s_{i}^{j}]$, and truncated moment $\mu_{i}^{j}=\operatorname{\mathbb{E}}[(\min\{s_{i},1\})^{j}]$ for every item $i\in[n]$ and moment $j\in[k]$, and oracle access to the function $q_{i}(B)=\Pr[s_{i}\leq B]$, where $B$ is the current remaining knapsack capacity. For every $\alpha$ there exists an input instance $I$ on which the algorithm’s performance is bounded by:
$\frac{\mathrm{ADAPT}(I)}{\operatorname{\mathbb{E}}[\mathcal{A}(I)]}\geq\Omega(\alpha).$
Similarly to the previous proof, we construct two types of items, one “good” and one “bad”, which are indistinguishable to $\mathcal{A}$ yet are critically different for the optimal algorithm.
Proof of Theorem 5.2.
Fix $\epsilon\in(0,1)$. Fix distinct sizes
$t_{j}:=1-2^{-j},\quad j=1,2,\dots,k,$
$t_{k+1}:=1.$
The two item types are defined as follows. For sufficiently small $\delta>0$ and for $p_{j}\in[0,1]$ to be determined later:
Property Type 1 Item (“good”) Type 2 Item (“bad”)
Reward $v_{i}$ 22 22
Cost $c_{i}$ $2-\epsilon$ $2-\epsilon$
Size $s_{i}$ dist. $s_{1}=\begin{cases}0,&\text{w.p. }1-2\epsilon\\ t_{j},&\text{w.p. }\tfrac{\epsilon}{k+1},\;j=1,\dots,k+1\\ 1,&\text{w.p. }\epsilon\end{cases}$ $s_{2}=\begin{cases}\delta,&\text{w.p. }1-2\epsilon\\ t_{j},&\text{w.p. }p_{j},\;j=1,\dots,k+1\\ 1,&\text{w.p. }\epsilon\end{cases}$
Note that since $\Pr[s_{i}\leq 1]=1$, the truncated moments are identical to the original moments. We now focus on finding the correct parameters such that the two distributions have the same moments; We seek $p_{j}$ with $\sum_{j}p_{j}=\epsilon$, satisfying for each $r=1,\dots,k$:
$\operatorname{\mathbb{E}}[s_{1}^{r}]=\operatorname{\mathbb{E}}[s_{2}^{r}].$
That is, for $r=1,\dots,k$:
$\sum_{j=1}^{k+1}\frac{\epsilon}{k+1}\,t_{j}^{r}+\epsilon$ $=(1-2\epsilon)\,\delta^{r}+\sum_{j=1}^{k+1}p_{j}\,t_{j}^{r}+\epsilon$
$\iff\sum_{j=1}^{k+1}t_{j}^{r}\,\left(\frac{\epsilon}{k+1}-p_{j}\right)$ $=(1-2\epsilon)\,\delta^{r}.$
Denote $x_{j}:=\frac{\epsilon}{k+1}-p_{j}$. Then $\sum_{j}x_{j}=0$ (since $\sum p_{j}=\epsilon$ and $\sum\frac{\epsilon}{k+1}=\epsilon$). We get the following system of equations:
$\sum_{j=1}^{k+1}x_{j}$ $=0;$ (5)
$\sum_{j=1}^{k+1}t_{j}^{r}\,x_{j}$ $=(1-2\epsilon)\,\delta^{r},\quad\text{for }r=1,\dots,k.$ (6)
We now rewrite (5) and (6) in matrix form $Ax=b$, where
$A=\begin{pmatrix}1&1&\cdots&1\\ t_{1}&t_{2}&\cdots&t_{k+1}\\ t_{1}^{2}&t_{2}^{2}&\cdots&t_{k+1}^{2}\\ \vdots&\vdots&&\vdots\\ t_{1}^{k}&t_{2}^{k}&\cdots&t_{k+1}^{k}\end{pmatrix},x=\begin{pmatrix}x_{1}\\ \vdots\\ x_{k+1}\end{pmatrix},b=\begin{pmatrix}0\\ (1-2\epsilon)\delta\\ \vdots\\ (1-2\epsilon)\delta^{k}\end{pmatrix}$.
Note that $A$ is a $(k+1)\times(k+1)$ matrix where the first row corresponds to the zero-sum constraint (5), and the subsequent $k$ rows correspond to the moment-matching conditions (6). This is a transposed Vandermonde matrix [17]. Since the $t_{j}$ are distinct, it yields a unique solution $x=A^{-1}b$. Hence
$p_{j}=\frac{\epsilon}{k+1}-x_{j}.$
It is left to prove that $p_{j}\geq 0$ to conclude that $s_{2}$ is indeed a valid distribution. Note that since $A$ and $A^{-1}$ are functions of $k$ (and the fixed $t_{j}$ values), then after fixing $k$, $x_{j}$ is a function of $\epsilon$ and $\delta$:
$x_{j}(\epsilon,\delta)=\sum_{r=1}^{k}(A^{-1})_{j,r+1}[(1-2\epsilon)\delta^{r}].$
For a fixed $\epsilon$, we have that $|x_{j}(\delta)|=O(\delta)$. Thus, for $p_{j}=\frac{\epsilon}{k+1}-x_{j}$ to be non-negative, we can choose $\delta$ sufficiently small such that $\frac{\epsilon}{k+1}$ dominates the $x_{j}$ terms, ensuring $p_{j}\geq 0$. Hence, the two distributions are indistinguishable to $\mathcal{A}$.
Inserting a single item of type 2 results in a net profit of $\epsilon$, and reduces the remaining capacity $B$ by at least $\delta$. Inserting any further items (of either types) results in a negative expected profit:
$\operatorname{\mathbb{E}}[\text{inserting item of type $i$}\;|\;B\leq 1-\delta]$ $\leq\Pr[s_{i}\neq 1]\cdot v_{i}-c_{i}$
$=(1-\epsilon)\cdot 2-(2-\epsilon)$
$=-\epsilon<0.$
We conclude the proof by considering an input instance $I$ with $n>\frac{1}{\epsilon}$ items of type 1 and $n^{2}-n$ items of type 2. The optimal policy does as well as the following policy: it will attempt to insert items of type 1, and stop if an item’s size instantiates $s_{i}\neq 0$, or if no more items of type 1 remain. The expected number of items successfully inserted by OPT is $\frac{1}{2\epsilon}$, each item contributing $\epsilon$ net profit, resulting in
$\mathrm{ADAPT}(I)\geq\frac{1}{2\epsilon}\cdot\epsilon=\frac{1}{2}.$
Since $\mathcal{A}$’s only strategy is to randomly select an item to insert, with probability $1-\frac{1}{n}$ it begins by inserting a “bad” item. By taking $n\to\infty$, the expected profit of $\mathcal{A}$ approaches $\epsilon$:
$\lim_{n\to\infty}\operatorname{\mathbb{E}}[\mathcal{A}(I)]=\epsilon.$
Thus for a large enough $n$, the performance ratio of the algorithm can be bounded:
$\frac{\mathrm{ADAPT}(I)}{\operatorname{\mathbb{E}}[\mathcal{A}(I)]}\geq\frac{\frac{1}{2}}{2\epsilon}\geq\frac{1}{8}\alpha.$
6 $O(1/\delta)$-approximation for Bounded-Size Jobs
In this section, we assume that all jobs are size-bounded with high probability; specifically, we assume that for some $\delta\in(0,1)$, for all $i,j$,
$\Pr_{s_{i}^{j}\sim F_{i}^{j}}[s_{i}^{j}>1-\delta]\leq\frac{1}{2n\alpha}.$
By Lemma 2.1, we focus on providing an approximately optimal policy for the MSKC problem.
As a first step, we analyze a modified problem, where there is no bound on the job sizes, and if a job exceeds the deadline, the principal collects the job’s reward, and then the process terminates. For this case, we give a policy that approximates the profit of the optimal policy that does not receive the reward of the job that overflows. We then show how to transform this policy into a policy that handles bounded-size jobs and does not collect the reward of overflowing jobs.
Our algorithm for the case where jobs can overflow, SKC-OF, runs SKC with budget $\delta$ and a different threshold $\mathcal{T}=\Phi(\delta)/4$. Note that, unlike the SKC algorithm, this threshold is independent of $\alpha$. We denote by $\textbf{SKC-OF}(t)$, the algorithm that runs SKC with a budget $t$, and a threshold $\mathcal{T}=\Phi(t)/4$. Specifically, $\textbf{SKC-OF}=\textbf{SKC-OF}(\delta)$. We have the following.
Lemma 6.1.
For any constant $c\geq 1$,
$\operatorname{\mathbb{E}}[\textbf{SKC-OF}(t)]\geq\frac{1}{c}\Phi(ct)/4.$
Proof.
In case $w_{\max}\geq\Phi(t)/4$, we only take $w_{\max}$ and obtain a profit of
$v_{i_{\max}}-p_{i_{\max}}^{j_{\max}}\geq w_{\max}\geq\Phi(t)/4\geq\frac{1}{c}\Phi(ct)/4,$
where the last inequality follows Lemma 3.7. Therefore, assume $w_{\max}<\Phi(t)/4$.
Recall that $\textbf{SKC-OF}_{k}$ is the profit of item $k$ in the SKC-OF algorithm, as defined in Section 3. In our case, the profit from item $k$ will be as large as in the MSKC instance, and will not incur a cost in case of an overflow. Therefore, by adapting Eq. (3), our bound on $\textbf{SKC-OF}_{k}$ will be
$\operatorname{\mathbb{E}}[\textbf{SKC-OF}_{k}]\geq w_{k}\left(1-\frac{1}{t}\sum_{j=1}^{k}{\mu_{j}}\right).$
We adapt the bound on SKC-OF from Eq. (4) accordingly to be
$\operatorname{\mathbb{E}}[\textbf{SKC-OF}]\geq\Phi(t)/2-w_{\max}\geq\Phi(t)/4\geq\frac{1}{c}\Phi(ct)/8,$ (7)
where the second inequality follows our assumption that $w_{\max}\leq\Phi(t)/4$ and the last inequality follows Lemma 3.7. ∎
We now turn to the bounded job size case. Our algorithm for this case, SKC-BOUND, will run SKC-OF, and pretend that the budget is $\delta$, i.e. it will terminate if the total size of jobs exceeds $\delta$.
Theorem 6.2.
$\frac{\mathrm{ADAPT}}{\operatorname{\mathbb{E}}[\textbf{SKC-BOUND}]}\leq O\left(\frac{1}{\delta}\right)$
Proof.
If all items are instantiated with sizes $s_{i}^{j}<1-\delta$, then since the policy terminates when we exceed the budget of $\delta$, the policy never overflows, achieving the expected value of SKC-OF. However, if $s_{i}^{j}>1-\delta$ for some $i,j$, the policy terminates and an overflow may occur, resulting in a loss of at most $\alpha w_{\max}$. We have that
$\operatorname{\mathbb{E}}[\textbf{SKC-BOUND}]$ $\geq\operatorname{\mathbb{E}}[\textbf{SKC-OF}]-\alpha w_{\max}\cdot\Pr[\exists\text{ item $i$ with }s_{i}^{j}>1-\delta]$
$\geq\Phi(\delta)/4-\alpha w_{\max}\cdot\sum_{i=1}^{n}\Pr[s_{i}^{j}>1-\delta]$
$\geq\Phi(\delta)/4-\alpha w_{\max}\cdot n\cdot\frac{1}{2n\alpha}$
$\geq\Phi(\delta)/4-\Phi(\delta)/8$
$=\Phi(\delta)/8.$
The theorem follows from Lemma 3.7,
$\frac{\mathrm{ADAPT}}{\operatorname{\mathbb{E}}[\textbf{SKC-BOUND}]}\leq\frac{\Phi(2)}{\Phi(\delta/8)}\leq\frac{\frac{16}{\delta}\Phi(\delta/8)}{\Phi(\delta/8)}=O\left(\frac{1}{\delta}\right)~.$
We note that SKC-BOUND determines a non-adaptive order of jobs-effort levels (as it applies SKC), but it applies an adaptive stopping rule, as it stops after the total realized sizes first exceed $\delta$. In Appendix B, we show that even for bounded-sized jobs, any non-adaptive policy is bounded away by a factor $\Omega(\sqrt{\alpha})$ from the optimal policy (Theorem B.1).
In Appendix C, we show that if indeed the principal gets the reward of any job that is initiated before the deadline (even if it finishes after the deadline), then we have a constant adaptivity gap Lemma C.1.
7 Conclusion and Future Directions
In this work, we initiate the study of contract design where rewards are fixed, but the principal incentivizes the agents to invest effort in finishing their jobs on time. We identify an economically meaningful Inverse of the Return-on-Investment parameter, $\alpha$, which captures many aspects of the problem’s complexity. Dean et al. [11] show a constant adaptivity gap, while we show that in our case, the adaptivity gap is $\Theta(\alpha)$. As the optimal adaptive contract is intractable, a remaining open problem is whether one can efficiently compute an adaptive contract whose profit is a good approximation to the optimal adaptive contract.
In Section 4.2, we consider a policy with limited adaptivity: the order is non-adaptive, but the policy stops processing jobs once it expects a negative profit. We show that for such policies, their profit can be a factor $\Omega(\alpha^{1/3})$ away from the optimal profit. As our algorithm in Section 3 guarantees an $O(\alpha)$-approximation, this leaves a gap between the lower and upper bound for such policies. It is also interesting to consider other limited-adaptivity policies, e.g., a policy that predetermines the agents’ order, but not their effort levels. Another direction is quantifying the profit loss from using a family of restricted contracts, such as contracts that do not pay agents who finish after the deadline.
Finally, many additional stochastic optimization problems, such as stochastic routing and others, have natural strategic variants were the uncertainty is controlled by an agent’s actions. E.g., in routing, service times are likely to result from a strategic choice of an economic entity. Marrying the classic algorithms developed for such problems with contractual incentives is an intriguing avenue for future research.
Appendix A LP Overconfidence: An $\Omega(\alpha)$ Overestimation of the True Profit
In this section, we provide an instance of the MSKC problem where the value of the LP introduced in Eq. (2) is larger by a factor $\Omega(\alpha)$ from the obtainable value of any policy. This is in stark contrast to the case without costs, where Dean et al. [11] show a non-adaptive policy who’s reward is a constant factor away from the value of the analogous LP. As our non-adaptive policy in Section 3 gives $O(\alpha)$-approximation to the value of the LP in Eq. (2), this approximation is the best possible with respect to the value of the LP.
In our instance, we have an infinite number of identical jobs with value $1$ and cost $1-\epsilon$. The size-distribution of the jobs is
$s=\begin{cases}\epsilon^{n},&\text{w.p. }1-\epsilon,\\ 1,&\text{w.p. }\epsilon;\end{cases}.$
By Definition 2.6, the value of the IOR parameter is $\alpha=\frac{1-\epsilon}{\epsilon}=O(1/\epsilon)$. For every job $i$, we have $\mu_{i}=(1-\epsilon)\cdot\epsilon^{n}+\epsilon\cdot 1=\epsilon^{n}+\epsilon-\epsilon^{n+1}\leq 2\epsilon$ and $w_{i}=\epsilon$. Setting $x_{i}=$ for $1/2\epsilon$ is a feasible assignment to the LP, which implies that $\Phi(1)\geq\frac{1}{2\epsilon}\cdot\epsilon=1/2.$ We next bound the reward of any policy for this instance.
For any policy, after executing the first job, we have that the any job fits with probability at most $1-\epsilon$; thus, we have that the benefit from trying to execute any job after the first one is at most $(1-\epsilon)\epsilon-\epsilon\cdot(1-\epsilon)=0.$ Therefore, the policy that only executes one job is optimal, and it obtains a reward of $\epsilon\leq\frac{\epsilon}{2}\Phi(1)=\Phi(1)/\Omega(\alpha).$
Appendix B The Need for Adaptivity for Bounded-Size Jobs
Theorem B.1.
There exists an instance $I$ of the Knapsack Contracts problem such that
$\frac{\mathrm{STOP\_ADAPT}(I)}{\mathrm{NON\_ADAPT}(I)}=\Omega\bigl{(}\sqrt{\alpha}\bigr{)}.$
Proof.
Let item:
Property Item
Reward $v$ 22
Cost $c$ $2-\epsilon^{2}$
Size $s$ dist. $s_{1}=\begin{cases}\epsilon^{n},&\text{w.p. }1-\epsilon,\\ \frac{1}{2},&\text{w.p. }\epsilon;\end{cases}$
We consider the case where $S^{\prime}<\frac{1}{2}$, and compute the expected profit from each job.
$\operatorname{\mathbb{E}}[w|S^{\prime}<\frac{1}{2}]=v\cdot\Pr[s\leq S^{\prime}]-c=2\cdot(1-\epsilon)-(2-\epsilon^{2})=2-2\epsilon-2+\epsilon^{2}=-2\epsilon+\epsilon^{2}<0$
In expectation, we can include approximately $1/\epsilon$ jobs before reaching a state where $S^{\prime}<\frac{1}{2}$ so, the expected number of jobs executed by the optimal stopping-time adaptive policy is $1/\epsilon$, and therefore, it obtains an expected profit of $\frac{1}{\epsilon}\cdot\epsilon^{2}=\epsilon.$
Consider now a non-adaptive policy. This policy decides on the number of jobs to execute, and ends whether it overflows or completes the number of jobs it sets to execute. Consider first a policy that only executes one job. Its reward is exactly $v-c=\epsilon^{2}$. Let $\mathrm{profit}(i)$ be the expected profit from the $i$th job in the sequence. Since the first job always finishes before the deadline, it is clear that $\mathrm{profit}(1,2)=c-v=\epsilon^{2}$. Consider now $\mathrm{profit}(i)$ for $i>2$. We have that
$\mathrm{profit}(i)$ $=$ $\operatorname{\mathbb{E}}[\mathrm{profit}(i)|s_{j}=\epsilon^{n}\ \forall j<i]\cdot\Pr[s_{j}=\epsilon^{n}\quad\forall j<i]$
$+\operatorname{\mathbb{E}}[\mathrm{profit}(i)|\exists!j<i:\ s_{j}>\epsilon^{n}]\cdot\Pr[\exists!j<i:\ s_{j}>\epsilon^{n}]$
$=$ $(v-c)\cdot(1-\epsilon)^{i-1}+((v-c)\cdot(1-\epsilon)-c\cdot\epsilon)\cdot\Pr[\exists!j<i:\ s_{j}>0]$
$=$ $\epsilon^{2}\cdot(1-\epsilon)^{i-1}+((1-\epsilon)v-c)\cdot\Pr[\exists!j<i:\ s_{j}>0]$
$=$ $\epsilon^{2}\cdot(1-\epsilon)^{i-1}+(\epsilon^{2}-2\epsilon)\cdot\Pr[\exists!j<i:\ s_{j}>0]$
$\leq$ $\epsilon^{2}\cdot(1-\epsilon)^{i-1}+(\epsilon^{2}-2\epsilon)\cdot\Pr[s_{i-1}>0\ \wedge\ s_{j}=0\quad\forall j<i-1]$
$=$ $\epsilon^{2}\cdot(1-\epsilon)^{i-1}+(\epsilon^{2}-2\epsilon)\cdot(1-\epsilon)^{i-2}\cdot\epsilon$
$=$ $\epsilon\cdot(1-\epsilon)^{i-2}(\epsilon\cdot(1-\epsilon)+\epsilon^{2}-2\epsilon)$
$\leq$ $\epsilon\cdot(1-\epsilon)^{i-2}(\epsilon^{2}-\epsilon)$
$<$ $0.$
In the above, the first inequality follows from the fact that $\epsilon^{2}-2\epsilon$ is negative, and that the probability that exactly one of the first i1i-1 jobs has size greater than $\epsilon^{n}$ is larger than the probability that only job i1i-1 has size greater than $\epsilon^{n}$. By Definition 2.6, the value of the IOR parameter is $\alpha=\frac{2-\epsilon^{2}}{\epsilon^{2}}=O(1/\epsilon^{2})$.
Therefore, we have that a non-adaptive policy only loses from having a sequence larger than 2, and the optimal non-adaptive policy has a profit of $2\epsilon^{2}$. We conclude that for this instance, we have $\frac{\mathrm{STOP\_ADAPT}(I)}{\mathrm{NON\_ADAPT}(I)}=\frac{1}{2\epsilon}=O(\sqrt{\alpha}).$
Appendix C Constant Approximation for MSKC with Overflows
We first provide an upper-bound for expected profit of the optimal adaptive policy that can overflow as well. Recall the definition of $\Phi$ given in Eq (2). Let $\mathrm{ADAPT}_{OF}$ be the optimal adaptive policy that can overflow.
Lemma C.1.
For every instance of the knapsack contract problem $I$, $\mathrm{ADAPT}_{OF}(1)\leq 2\Phi(1)+w_{max}$.
Proof.
let $\mathcal{S}$ be the random set of item choices that the policy inserts, and $w_{ij}=v_{i}\cdot\Pr_{{s_{i}^{j}\sim F_{i}^{j}}}[s_{i}^{j}\leq 1]-p_{i}^{j}$ (like in Section 3), $w_{\max}=\max_{(i,j)\in\mathcal{C}}w_{ij}$ (Algorithm 1). $\Phi(t)$ is defined in Eq. (2), .
$\mathrm{ADAPT}_{OF}(1)$ $=$ $\sum_{ij}w_{ij}\cdot\Pr[(i,j)\in\mathcal{S}]$ (8)
$=$ $\sum_{ij}w_{ij}\cdot\Pr[(i,j)\in\mathcal{S}\wedge s_{i}\leq 1]+\sum_{ij}w_{ij}\cdot\Pr[(i,j)\in\mathcal{S}\wedge s_{i}>1]$
$\leq$ $\sum_{ij}w_{ij}\cdot\Pr[s_{i}\leq 1]\cdot\Pr[(i,j)\in\mathcal{S}]+w_{max}\sum_{ij}\Pr[(i,j)\in\mathcal{S}\wedge s_{i}>1]$
$\leq$ $\sum_{ij\in\mathcal{S}}w_{ij}\cdot x_{ij}+w_{max}$
$\leq$ $\Phi(2)+w_{max}$
$\leq$ $2\Phi(1)+w_{max}.$
Here the first inequity follows since the probability $s_{i}>1$ is independent from the fact $(i,j)\in\mathcal{S}$. The last inequality follows from the fact that, at each run of the adaptive policy, at most one element has a size larger than 1, and therefore the sum of probabilities of the right summand is at most 1. By Lemma 3.6, we know that $\operatorname{\mathbb{E}}[\mu(\mathcal{S})]\leq 2$, therefore, we get $\sum_{ij\in A}w_{ij}\cdot x_{ij}\leq\Phi(2)$. ∎
Theorem C.2.
$\textbf{SKC}_{OF}$ is a 9-approximation algorithm to $\mathrm{ADAPT}_{OF}$.
Proof.
By Equations (7),(8):
$\frac{ADAPT_{OF}}{\textbf{SKC-OF}}$ $\leq$ $\frac{2\Phi(1)+w_{max}}{\Phi(1)/4}.$
If $w_{max}>\Phi(1)/4$: so $\textbf{SKC-OF}=w_{max}$ and:
$\frac{ADAPT_{OF}}{\textbf{SKC-OF}}$ $\leq$ $\frac{8w_{max}+w_{max}}{w_{max}}=9.$
Else:
$\frac{ADAPT_{OF}}{\textbf{SKC-OF}}$ $\leq$ $\frac{2\Phi(1)+\Phi(1)/4}{\Phi(1)/4}=9.$
Appendix D Ordered Adaptive Contracts
In many real-world scenarios, the principal lacks the flexibility to reorder agents at will. For example, this can be due to procedural requirements, like a fixed sequence of suppliers in a supply chain, or pre-determined project plan with sequential dependencies. In these settings, the principal’s adaptive power is limited to deciding, after observing the outcome of each task, whether or not to continue with the next agent, and if so – which effort level to incentivize. We define the Ordered Adaptive Contracts model:
Definition D.1 (Ordered Adaptive Contracts Model).
In the Ordered Adaptive Contracts model, the principal is given a fixed ordering $\sigma=(1,2,\ldots,n)$ of the $n$ agents. For each agent $i$ in order, the principal offers a contract $t_{i}$, and the agent chooses an effort level $j$ to maximize their utility. The agents execute their jobs sequentially in the given order. The principal observes the completion time of each job and may adaptively decide whether to proceed to the next agent or stop, but cannot change the order. The objective is to maximize the expected net profit (total value from completed jobs before the budget is exhausted minus total payments to agents), given the fixed order $\sigma$.
Since the order is fixed, we are interested in comparing our policy’s performance to the performance of the optimal policy operating on the same order of agents. By using the same reduction to the multi-agent contract problem as shown in Lemma 2.1, we focus on solving the matching stochastic knapsack with costs problem. In a similar manner to the algorithm that finds an ordered adaptive policy presented in [11] and to the PTAS presented in [5] for the regular Stochastic Knapsack variant (single-choice, no-cost), we use a resource augmentation approach: the policy is allowed a knapsack of size $(1+\epsilon)$, while the optimal adaptive policy has a knapsack of size $1$. Our algorithm runs in time $O(\frac{n^{3}m}{\epsilon^{2}})$, offering a tradeoff between the augmented size of the knapsack and the running time of the algorithm.
We use a dynamic programming algorithm with discretized sizes. We denote by $\mathcal{D}(i,k)$ the optimal expected value that can be obtained from inserting items i,,ni,...,n, with $\delta k$ capacity left in the knapsack, for some small $\delta>0$ to be determined later. We are interested in computing $\mathcal{D}(1,\lfloor\frac{1}{\delta}\rfloor)$. If all of the real sizes are indeed multiples of $\delta k$, then one can obtain an optimal policy by computing:
$\mathcal{D}(i,k)=\max\left\{\mathcal{D}(i+1,k),\max_{j\in[m]}\left\{v_{i}\cdot\Pr[s_{i}^{j}\leq\delta k]-p_{i}^{j}+\sum_{t=0}^{k}{\mathcal{D}(i+1,k-t)\cdot\Pr[s_{i}^{j}=t\delta]}\right\}\right\}.$
With the base cases of $\mathcal{D}(n+1,k)=0$ for all $k$, and $\mathcal{D}(i,k)=0$ for $k<0$.
However, when the distributions are arbitrary, a slight deviation from the discretization can result in an arbitrarily bad solution. To overcome this, we must utilize the fact our policy is allowed a knapsack of size $1+\epsilon$. We do so by rounding down the true distribution at each step to the nearest multiple of $\delta$. By picking a small enough $\delta$, we ensure that the overall rounding error is not too large.
Theorem D.1.
There exists an $O(\frac{n^{3}m}{\epsilon^{2}})$ time algorithm that finds a policy for the ordered adaptive knapsack with costs problem, where the policy has a knapsack of size $(1+\epsilon)$, and the optimal adaptive policy has a knapsack of size $1$. The policy achieves an expected profit of at least the profit of the optimal policy with a knapsack of size $1$.
Proof.
Let OAC be our algorithm and $\mathrm{ADAPT}$ be the optimal adaptive policy. We set the discretization parameter $\delta=\epsilon/n$.
First, consider a modified problem where all item sizes $s_{i}^{j}$ are rounded down to the nearest multiple of $\delta$: $\hat{s}_{i}^{j}=\delta\lfloor s_{i}^{j}/\delta\rfloor$. Let $\widehat{\mathrm{ADAPT}}$ be the optimal expected profit for this problem with rounded-down sizes. Since for any policy, the probability of successfully fitting a set of items can only increase when item sizes are reduced, the optimal value for the rounded-down problem is an upper bound on the optimal value for the original problem. Thus, $\widehat{\mathrm{ADAPT}}\geq\mathrm{ADAPT}.$
Our algorithm solves the dynamic program $\mathcal{D}$ described above for the rounded-down sizes. This DP computes the exact optimal policy for the rounded-down problem. Denote by $\varphi_{\textbf{OAC}}$ the policy derived from this DP. When given a knapsack of capacity $1$, this policy achieves the optimal value for the rounded-down problem. Let $\widehat{\mathrm{profit}}(\varphi_{\textbf{OAC}},\text{capacity}=1)$ be a random variable of the profit of the policy with the original capacity and rounded-down job sizes. We have that $\operatorname{\mathbb{E}}[\widehat{\mathrm{profit}}(\varphi_{\textbf{OAC}},\text{capacity}=1)]=\widehat{\mathrm{ADAPT}}$. Combining these, we have $\operatorname{\mathbb{E}}[\widehat{\mathrm{profit}}(\varphi_{\textbf{OAC}},\text{capacity}=1)]\geq\mathrm{ADAPT}.$
Now, we analyze the performance of our algorithm, which executes the policy $\varphi_{\textbf{OAC}}$ on the original item sizes $s_{i}^{j}$ with an augmented capacity of $1+\epsilon$. Consider any execution path of $\varphi_{\textbf{OAC}}$ where a sequence of items and choices $\mathcal{S}\subset[n]\times[m]$ is attempted and successfully completed. According to the policy $\varphi_{\textbf{OAC}}$ (which was computed for capacity 1 and sizes $\hat{s}$), the sum of the rounded-down sizes must be no more than 1:
$\sum_{(i,j)\in\mathcal{S}}\hat{s}_{i}^{j}\leq 1.$
The true total size of these items is $\sum_{(i,j)\in\mathcal{S}}s_{i}^{j}$. We know that for each item, $s_{i}^{j}=\hat{s}_{i}^{j}+\eta_{i}^{j}$, where $0\leq\eta_{i}^{j}<\delta$ is the rounding error. The total size in the original problem is:
$\sum_{(i,j)\in\mathcal{S}}s_{i}^{j}=\sum_{(i,j)\in\mathcal{S}}\hat{s}_{i}^{j}+\sum_{(i,j)\in\mathcal{S}}\eta_{i}^{j}.$
Since at most $n$ items can be placed in the knapsack, the total accumulated rounding error is bounded:
$\sum_{(i,j)\in\mathcal{S}}\eta_{i}^{j}\leq n\cdot\delta=n\cdot(\epsilon/n)=\epsilon.$
Therefore, the true total size of the items is:
$\sum_{(i,j)\in\mathcal{S}}s_{i}^{j}\leq\sum_{(i,j)\in\mathcal{S}}\hat{s}_{i}^{j}+\epsilon\leq 1+\epsilon.$
This shows that any sequence of items that fits within capacity 1 in the rounded-down instance is guaranteed to fit within the augmented capacity of $1+\epsilon$ in the original instance. Since the policy $\varphi_{\textbf{OAC}}$ makes decisions based on the sequence of outcomes (which are independent of the rounding), every successful path under capacity 1 with sizes $\hat{s}$ is also a successful path under capacity $1+\epsilon$ with sizes $s$. Moreover, it crucially follows that an item can cause an overflow under capacity $1+\epsilon$ with sizes $s$ only if it causes an overflow under capacity 1 with sizes $\hat{s}$. Therefore, the expected profit can only be higher:
$\operatorname{\mathbb{E}}[\mathrm{profit}(\varphi_{\textbf{OAC}},\text{capacity}=1+\epsilon)]\geq\operatorname{\mathbb{E}}[\widehat{\mathrm{profit}}(\varphi_{\textbf{OAC}},\text{capacity}=1)].$
Chaining the inequalities, we get the final result:
$\textbf{OAC}=\operatorname{\mathbb{E}}[\mathrm{profit}(\varphi_{\textbf{OAC}},\text{capacity}=1+\epsilon)]\geq\operatorname{\mathbb{E}}[\widehat{\mathrm{profit}}(\varphi_{\textbf{OAC}},\text{capacity}=1)]\geq\mathrm{ADAPT}.$
The dynamic program $\mathcal{D}$ has $O(n)\cdot O(n/\epsilon)$ states, and computing each state takes up to $O(m\cdot n/\epsilon)$ time. This results in a running time of $O(\frac{n^{3}m}{\epsilon^{2}})$.
Appendix E Omitted Proofs from Section 3
Full proof of Lemma 3.3.
For two solutions $x,x^{*}$, denote by $\Delta_{w}$ and $\Delta_{\mu}$ the difference in the total values and the total expected truncated sizes, respectively, between $x^{*}$ and $x$. We want to show that our constructed solution $x^{*}$ has at least as much value ($\Delta_{w}\geq 0$) and at most as much expected truncated size ($\Delta_{\mu}\leq 0$) as the original solution $x$, to conclude that $x^{*}$ is a feasible optimal solution.
For the single dominance case, we define
$x^{*}_{ij^{\prime}}=x_{ij^{\prime}}+x_{ij},\quad x^{*}_{ij}=0~.$
We have that
$\Delta_{w}$ $=(x_{ij^{\prime}}^{*}-x_{ij^{\prime}})w_{ij^{\prime}}+(x_{ij}^{*}-x_{ij})w_{ij}$
$=x_{ij}w_{ij^{\prime}}-x_{ij}w_{ij}$
$=x_{ij}(w_{ij^{\prime}}-w_{ij})$
$\geq 0$ $\text{(since $w_{ij^{\prime}}\geq w_{ij}$)}~,$
$\Delta_{\mu}$ $=(x_{ij^{\prime}}^{*}-x_{ij^{\prime}})\mu_{ij^{\prime}}+(x_{ij}^{*}-x_{ij})\mu_{ij}$
$=x_{ij}\mu_{ij^{\prime}}-x_{ij}\mu_{ij}$
$=x_{ij}(\mu_{ij^{\prime}}-\mu_{ij})$
$\leq 0$ $\text{(since $\mu_{ij^{\prime}}\leq\mu_{ij}$)}~.$
For the double dominance case, we define
$x^{*}_{ij^{\prime}}=x_{ij^{\prime}}\;+\;x_{ij}\cdot\frac{w_{ij^{\prime\prime}}-w_{ij}}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}},\quad x^{*}_{ij^{\prime\prime}}=x_{ij^{\prime\prime}}\;+\;x_{ij}\cdot\frac{w_{ij}-w_{ij^{\prime}}}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}},\quad x^{*}_{ij}=0~.$
We have that
$\Delta_{w}$ $=(x_{ij^{\prime}}^{*}-x_{ij^{\prime}})w_{ij^{\prime}}+(x_{ij^{\prime\prime}}^{*}-x_{ij^{\prime\prime}})w_{ij^{\prime\prime}}+(x_{ij}^{*}-x_{ij})w_{ij}$
$=x_{ij}w_{ij^{\prime}}\cdot\frac{w_{ij^{\prime\prime}}-w_{ij}}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}}+x_{ij}w_{ij^{\prime\prime}}\cdot\frac{w_{ij}-w_{ij^{\prime}}}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}}-x_{ij}w_{ij}$
$=x_{ij}\left(\frac{w_{ij^{\prime}}(w_{ij^{\prime\prime}}-w_{ij})+w_{ij^{\prime\prime}}(w_{ij}-w_{ij^{\prime}})-w_{ij}(w_{ij^{\prime\prime}}-w_{ij^{\prime}})}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}}\right)$
$=0~,$
$\Delta_{\mu}$ $=(x_{ij^{\prime}}^{*}-x_{ij^{\prime}})\mu_{ij^{\prime}}+(x_{ij^{\prime\prime}}^{*}-x_{ij^{\prime\prime}})\mu_{ij^{\prime\prime}}+(x_{ij}^{*}-x_{ij})\mu_{ij}$
$=x_{ij}\mu_{ij^{\prime}}\cdot\frac{w_{ij^{\prime\prime}}-w_{ij}}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}}+x_{ij}\mu_{ij^{\prime\prime}}\cdot\frac{w_{ij}-w_{ij^{\prime}}}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}}-x_{ij}\mu_{ij}$
$=x_{ij}\left(\frac{\mu_{ij^{\prime}}(w_{ij^{\prime\prime}}-w_{ij})+\mu_{ij^{\prime\prime}}(w_{ij}-w_{ij^{\prime}})-\mu_{ij}(w_{ij^{\prime\prime}}-w_{ij^{\prime}})}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}}\right)~.$
From double dominance we have that
$\frac{\mu_{ij}-\mu_{ij^{\prime}}}{w_{ij}-w_{ij^{\prime}}}\geq\frac{\mu_{ij^{\prime\prime}}-\mu_{ij}}{w_{ij^{\prime\prime}}-w_{ij}}$
Since $w_{ij}-w_{ij^{\prime}}>0$ and $w_{ij^{\prime\prime}}-w_{ij}>0$, we can cross-multiply:
$(\mu_{ij}-\mu_{ij^{\prime}})(w_{ij^{\prime\prime}}-w_{ij})\geq(\mu_{ij^{\prime\prime}}-\mu_{ij})(w_{ij}-w_{ij^{\prime}})$
Expanding both sides gives:
$\mu_{ij}w_{ij^{\prime\prime}}-\mu_{ij}w_{ij}-\mu_{ij^{\prime}}w_{ij^{\prime\prime}}+\mu_{ij^{\prime}}w_{ij}\geq\mu_{ij^{\prime\prime}}w_{ij}-\mu_{ij^{\prime\prime}}w_{ij^{\prime}}-\mu_{ij}w_{ij}+\mu_{ij}w_{ij^{\prime}}$
We can cancel the $\mu_{ij}w_{ij}$ term from both sides. Rearranging the terms to match the numerator of our expression for $\Delta_{\mu}$ yields:
$\mu_{ij^{\prime}}(w_{ij^{\prime\prime}}-w_{ij})+\mu_{ij^{\prime\prime}}(w_{ij}-w_{ij^{\prime}})-\mu_{ij}(w_{ij^{\prime\prime}}-w_{ij^{\prime}})\leq 0~.$
This confirms the numerator is non-positive. Since $w_{ij^{\prime\prime}}>w_{ij^{\prime}}$ we conclude that $\Delta_{\mu}\leq 0$.
Proof of Lemma 3.5.
$\Pr[size(\mathcal{S})\geq t]$ $=\Pr[\min\{size(\mathcal{S}),t\}\geq t]$
$\leq\frac{1}{t}\operatorname{\mathbb{E}}[\min\{size(S),t\}]$
$=\frac{1}{t}\operatorname{\mathbb{E}}[{\min\{\sum_{(i,j)\in\mathcal{S}}{s_{i}^{j}},t\}}]$
$\leq\frac{1}{t}\operatorname{\mathbb{E}}[\sum_{(i,j)\in\mathcal{S}}{\min\{s_{i}^{j},t\}}]$
$=\frac{1}{t}\sum_{(i,j)\in\mathcal{S}}{\operatorname{\mathbb{E}}[\min\{s_{i}^{j},t\}]}$
$=\frac{1}{t}\sum_{(i,j)\in\mathcal{S}}{\mu_{ij}}=\frac{1}{t}\mu(S)$
Proof of Lemma 3.6.
Let $\mathcal{S^{-}}$ be the (random) set of item choices that the policy inserts except the final item choice. By the stopping condition
$size(\mathcal{S}^{-})<t~.$
It follows that for every $(i,j)\in\mathcal{S}^{-}$, $s_{i}^{j}<t$ and therefore $\mu_{ij}=\operatorname{\mathbb{E}}[\min\{s_{i}^{j},t\}]=\operatorname{\mathbb{E}}[s_{i}^{j}]$. Denote by $(i^{*},j^{*})$ the final item choice made by the policy. Then,
$\operatorname{\mathbb{E}}[\mu(\mathcal{S})]$ $=\operatorname{\mathbb{E}}[\sum_{(i,j)\in\mathcal{S}}{\mu_{ij}}]=\operatorname{\mathbb{E}}[\sum_{(i,j)\in\mathcal{S}^{-}}{\mu_{ij}}]+\operatorname{\mathbb{E}}[\mu_{i^{*}j^{*}}]$
$=\operatorname{\mathbb{E}}[\sum_{(i,j)\in\mathcal{S}^{-}}{s_{i}^{j}}]+\operatorname{\mathbb{E}}[s_{i^{*}}^{j^{*}}]=\operatorname{\mathbb{E}}[size(\mathcal{S}^{-})]+\operatorname{\mathbb{E}}[s_{i^{*}}^{j^{*}}]$
$\leq t+t=2t~.$
Proof of Lemma 3.7.
For any adaptive policy, denote by $\mathcal{S}$ the random set of pairs $(i,j)$ that the policy inserts. Set $x_{ij}=\Pr[(i,j)\in\mathcal{S}]$. From the previous lemma, the expected mass that $\mathrm{ADAPT}$ attempts to insert is $\operatorname{\mathbb{E}}(\mu(\mathcal{S}))=\sum_{ij}\mu_{ij}x_{ij}\leq 2t$, therefore, it is a feasible solution to $\Phi(2t)$ with $\sum_{ij}w_{ij}x_{ij}\leq\Phi(2t)$. Since $w_{ij}x_{ij}$ is an upper bound on the expected profit of inserting the item choice $(i,j)$, we have that
$\mathrm{ADAPT}\leq\sum_{ij}{w_{ij}x_{ij}}\leq\Phi(2t)\leq\Phi(ct)~,$
for any $c\geq 2$. Now, given an optimal solution $x$ to $\Phi(ct)$, we can construct a feasible solution $x^{\prime}$ to $\Phi(t)$ by setting $x^{\prime}_{ij}=\frac{x_{ij}}{c}$ for all $i,j$, and then increasing the value of $x^{\prime}_{ij_{\bot}}$ to ensure that $\sum_{j}x^{\prime}_{i,j}=1$ for all $i$. This results in a feasible solution to $\Phi(t)$ with $\sum_{ij}w_{ij}x^{\prime}_{ij}=\frac{1}{c}\sum_{ij}w_{ij}x_{ij}$. We conclude that
$\mathrm{ADAPT}\leq\Phi(ct)\leq c\cdot\Phi(t)$
Appendix F Proof of Lemma 2.1
Proof.
Given a knapsack contracts instance, for each agent $i\in[n]$, we create a corresponding item $i$ in the MSKC instance. For each implementable effort level $j$ of agent $i$, we create a corresponding choice $j$ for item $i$. The parameters are mapped as follows:
  • The value of item $i$ is set to the principal’s reward $v_{i}$.
  • The size of choice $(i,j)$ is a random variable with the distribution $F_{i}^{j}$.
  • The cost of choice $(i,j)$ is set to $p_{i}^{j}=\operatorname{\mathbb{E}}_{s\sim F_{i}^{j}}[t_{i}(s)]$, where $t_{i}$ is the contract that minimizes the expected payment, as given by the LP in Definition 2.1.
Now, consider any policy $\varphi$ for the contracts problem, and denote by $\varphi^{\prime}$ the corresponding policy for the knapsack problem. $\varphi$ specifies the choice of (agent, effort level) pairs for the contracts problem, and $\varphi^{\prime}$ specifies the corresponding (item, choice) pairs for the knapsack problem.
It is left to show that the expected profits are identical. Denote by $S$ the set of successfully completed jobs in the contracts and knapsack problems, and by $A$ the set of agents/items attempted in the contracts and knapsack problems. Notice that $S$ and $A$ have exactly the same distribution under $\varphi$ and $\varphi^{\prime}$, as both policies make the same choices under the same randomization.
The profit for policy $\varphi$ is a random variable that depends on the (possibly random) choices of the policy and the random sizes of the attempted jobs. Its expectation is given by
$\operatorname{\mathbb{E}}[\mathrm{profit}(\varphi)]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in S}{v_{i}}-\sum_{(i,j)\in A}{t_{i}(s_{i})}\right]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in S}{v_{i}}\right]-\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in A}{t_{i}(s_{i})}\right].$ (9)
Expanding the transfer term, we get:
$\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in A}{t_{i}(s_{i})}\right]$ $=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in A}{t_{i}(s_{i})}\right]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)}{\mathbb{1}_{(i,j)\in A}\cdot t_{i}(s_{i})}\right]=\sum_{(i,j)}\operatorname{\mathbb{E}}\left[{\mathbb{1}_{(i,j)\in A}\cdot t_{i}(s_{i})}\right]$
$\stackrel{{\scriptstyle(\star)}}{{=}}\sum_{(i,j)}\operatorname{\mathbb{E}}\left[\mathbb{1}_{(i,j)\in A}\right]\cdot\operatorname{\mathbb{E}}_{s_{i}\sim F_{i}^{j}}\left[t_{i}(s_{i})\right]\stackrel{{\scriptstyle(\star\star)}}{{=}}\sum_{(i,j)}\operatorname{\mathbb{E}}\left[\mathbb{1}_{(i,j)\in A}\right]\cdot p_{i}^{j}$
$=\operatorname{\mathbb{E}}\left[\sum_{(i,j)}\mathbb{1}_{(i,j)\in A}\cdot p_{i}^{j}\right]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in A}p_{i}^{j}\right],$
where equality $(\star)$ is due to the independence between $\mathbb{1}_{(i,j)\in A}$ and $s_{i}^{j}$, and equality $(\star\star)$ is due to the definition of $p_{i}^{j}$. Plugging this back in Eq. 9 yields:
$\operatorname{\mathbb{E}}[\mathrm{profit}(\varphi)]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in S}{v_{i}}-\sum_{(i,j)\in A}p_{i}^{j}\right]=\operatorname{\mathbb{E}}[\mathrm{profit}(\varphi^{\prime})]~,$
which implies the desired.
Thus, given a $\varphi^{\prime}$ that $c$-approximate the optimal policy in the constructed instance of the MSKC problem, the corresponding $\varphi$ $c$-approximates the optimal policy in the original knapsack contracts instance. ∎