### v2

Abstract
We revisit the Stochastic Knapsack problem, where a policy-maker chooses an execution order for jobs with fixed values and stochastic running-times, aiming to maximize the value completed by a deadline. Dean et al. (FOCS’04) show that simple non-adaptive policies can approximate the (highly adaptive) optimum, initiating the study of adaptivity gaps.
We introduce an economically motivated generalization in which each job also carries an execution cost. This uncovers new applications, most notably a new and natural variant of contract design: jobs are processed by agents who choose among effort levels that induce different processing-time distributions, and the principal must decide which jobs to run and what payments induce the intended effort.
With costs, the objective becomes mixed-sign: value from completed jobs must be balanced against costs of execution, and running a job that misses the deadline can create negative utility. This changes the algorithmic picture, and the adaptivity gap is no longer constant. We give an economic explanation: the performance of non-adaptive policies is governed by jobs’ return on investment (ROI) — utility over cost — which can be arbitrarily small.
We introduce a hierarchy of increasingly adaptive policies, trading off simplicity and adaptivity. We prove near-tight guarantees across the hierarchy, showing that with costs the adaptivity gap is $\Theta(\alpha)$, where $1/\alpha$ is the ROI. Higher in the hierarchy, we identify an efficiently computable policy with limited adaptivity that is approximately-optimal. Analogous to the centrality of ROI in economics, we believe our ROI-based, simple-vs-optimal approach to adaptivity may be useful for additional stochastic optimization and online problems with mixed-sign objectives.
1 Introduction
In many economic decision-making settings, actions have stochastic outcomes and are associated not only with value but also with inherent costs. While most of the existing literature focuses on actions that generate value alone, in this paper we study how incorporating costs can significantly change the algorithmic design of effective solution methods. Understanding how costs affect decision-making scenarios is a central challenge in designing practical, near-optimal policies in many practical settings.
Stochastic Knapsack and Adaptivity Gaps.
In this work, we consider one of the classic stochastic decision problems studied in the seminal paper of Dean et al. [10]: the Stochastic Knapsack problem. In this problem, a policy-maker designs a policy that sequentially executes jobs, each associated with a known value and an uncertain processing time. The policy-maker faces a fixed deadline and obtains value only from jobs that complete before this deadline; any job finishing after the deadline yields zero value. While processing times are uncertain, the policy-maker is assumed to know their distributions, and seeks a policy that maximizes the expected total value obtained before the deadline.
A key insight of Dean et al. [10] is the trade-off between policy simplicity and optimality, formalized through the notion of adaptivity gap. The optimal policy may adaptively choose the next job to process based on the realized processing times of previously executed jobs. Such adaptivity can be extremely complex: the optimal policy can be represented as a decision tree, whose size might be exponential in the number of jobs.
Motivated by this potential complexity, Dean et al. [10] study simple non-adaptive policies, in which jobs are ordered in advance and executed sequentially until the deadline, collecting rewards from those that complete in time. Despite their simplicity, they show that such policies achieve a constant-factor approximation of the optimal adaptive policy, implying that the adaptivity gap is bounded by a constant. While the results of Dean et al. [10] demonstrate the surprising power of simple non-adaptive policies, they rely on the assumption that jobs can be executed without incurring costs.
Costly Job Execution and Multiple Execution Modes.
In many fundamental applications, assuming job execution is cost-free is unrealistic. We therefore study a setting with costly job execution, motivated by applications such as cloud computing or contracting with workers, where executing a job requires explicit compensation and thus incurs a cost.
Moreover, each job may be executed at multiple levels of effort, where higher effort corresponds to higher cost, and different efforts induce different distributions over processing times.
For example, the choice of the machine used to process a job, or of the specialist hired to perform a task, affects both the execution cost and the processing-time distribution.
Although these extensions may appear modest, they fundamentally change the role of adaptivity. Notably, we show that when execution costs are present, the adaptivity gap of non-adaptive policies can be unbounded. Nevertheless, we identify a natural and economically meaningful parameter of the setting—capturing a notion of return on investment (ROI)—that governs the adaptivity gap in the presence of costs. We then show that a limited amount of carefully-applied adaptivity suffices to obtain policies whose performance is within a constant factor of the optimal adaptive policy.
Return on Investment.
We notice that “hard instances” of our problem have the following structural property: For some of the jobs, their cost almost eclipses their value. Thus, when executing such jobs, they provide a small return on investment (ROI). More formally, the ratio between the obtained utility (value minus cost) from running the job and the incurred cost is small.
Such jobs are intuitively problematic in the knapsack context — their execution adds to the policy-maker’s cost, but their value might be lost if they finish after the deadline, in which case the policy-maker’s loss can be substantial. We essentially show that these are the only problematic jobs.
Hierarchy of Adaptivity.
Dean et al. [10] studied the Stochastic Knapsack problem by considering two extremes:
  • Non-adaptive policies. These policies choose an ordered set of jobs. The set is then processed in order until either all jobs are completed or the deadline is met, whichever occurs first.
  • Fully-adaptive policies. This is the most general policy class: there is no predetermined job order. The policy chooses which job to run next based on the observed runtime history.
When costs are introduced, simple non-adaptive policies no longer guarantee a constant approximation to the value of the optimal fully-adaptive policies. We explore the landscape through the simplicity-optimality lens, defining and analyzing the following intermediate classes of policies:
  • Stopping-adaptive policies. Policies in this class choose an ordered set of jobs, but stop executing once, given the time left until the deadline, any future sequence of jobs results in a negative expected utility.
  • Skipping-adaptive policies. These policies also commit to an ordered list, but unlike stopping-adaptive policies, they may adaptively skip some of the listed jobs.
See Section 2.2 for formal definitions and how the decision tree policy representation size decreases along the hierarchy. In the appendix we also consider non-stopping policies, which are non-adaptive policies that never halt and execute all jobs in the ordered set even after the deadline has passed.
When there are multiple execution modes in addition to costs, the above classes of policies extend naturally by ordering pairs of (job, execution mode) rather than just ordering jobs.
Motivating Example.
To illustrate the differences among these policy types, consider the following motivating example, in which a contractor hires subcontractors to perform jobs under a time constraint. The contractor’s level of commitment and flexibility varies across policy classes.
Under a fully-adaptive policy, the contractor never commits to an execution order in advance. After each observed runtime, the contractor may choose any remaining job and execution mode next, redesigning the plan on the fly as a function of the realized history and the remaining time. Such a high degree of adaptivity, however, can significantly complicate the system, both for the contractor, who must continuously re-optimize decisions, and for the subcontractors, who face uncertainty regarding their scheduling times and effort level.
We therefore turn to simpler policy classes that impose an underlying fixed execution order on the jobs, specified by a priority list of job–execution-level pairs determined in advance. Under non-adaptive policies, once this priority list is fixed, the contractor is obliged to continue attempting jobs in order whenever any time remains. In this setting, a job that is highly unlikely to finish on time will still get executed,
incurring its execution cost without obtaining its value. As we show, this behavior can be highly inefficient, particularly when the return on investment of the final job is low.
To mitigate this issue, stopping-adaptive policies introduce a limited form of adaptivity by allowing the contractor to stop execution once future jobs cannot yield positive utility. Unfortunately, as we demonstrate, this additional adaptivity is insufficient to guarantee a constant-factor approximation.
Finally, we consider a class of skipping-adaptive policies that augments a fixed priority order with job-specific time thresholds. A job is executed upon reaching its position in the order only if the remaining time before the deadline is at least its threshold; otherwise, it is skipped. We show this rather simple policy class is expressive enough to guarantee a constant-factor approximation.
Application to Contract Design — Knapsack Contracts.
Our initial motivation in studying the stochastic knapsack with costs problem was to study a variant of the principal-agent model, which we term Knapsack Contracts. In this model, the value of each executed task is fixed, but the duration of the task may depend on the effort level of the (privately taken) action of an agent. Moreover, the principal has a deadline by which they need to complete all tasks. Using contracts to incentivize timely execution of tasks is prevalent in many time-sensitive tasks such as food delivery and contractors outsourcing tasks.
In Appendix B, we formally present the Knapsack Contracts model. We show that the Knapsack Contracts problem is reducible to the multi-choice stochastic knapsack with costs problem. Thus, providing approximation guarantees to the latter automatically translates to approximation guarantees for the former.
Proposition 1.1.
[Reduction to Multi-choice Stochastic Knapsack with Costs] Given an efficient algorithm that finds a policy that $c$-approximates the optimal policy for the MSKC problem, one can efficiently find a policy that $c$-approximates the optimal policy for the Knapsack Contracts problem.
1.1 Our Contributions and Techniques
In Section 2.1, we formally define the problem through which we explore the simplicity-optimality landscape when introducing costs, the Multi-choice Stochastic Knapsack with Costs problem (or MSKC). In this problem, each job is associated with multiple levels of execution, where each level has a different execution cost and induces a different size distribution. The decision-maker computes a policy to choose the (job, level) execution order, obtaining a utility of the value of jobs completed by the deadline minus the cost of all executed jobs. In Section 2.2 we formally define the different policies in our adaptiveness hierarchy.
As mentioned above, hard instances are ones where for some jobs, their cost almost “covers” their value. Thus, when executing such jobs, the policy is left with a small return on investment (ROI). In order to parameterize this, we inspect the Inverse of the ROI (or IOR) of a job (as the ROI diminishes, the IOR grows). We introduce an economically-meaningful parameter $\alpha$, which represents the worst IOR among all jobs—see Definition 2.1. We show this parameter is closely tied to the performance of non-adaptive policies in our problem. An overview of our main findings is given in Figure 1.
[…]
Figure 1: Hierarchy of policy adaptivity levels showing the approximation gaps between different classes.
1.1.1 $\Theta(\alpha)$ Adaptivity Gap.
In Section 3, we study the power of non-adaptive policies. Our results are the following.
Upper Bound.
We give an efficient algorithm for computing an ordered set of (job, execution level) pairs that give an $O(\alpha)$-approximation to the utility of the (intractable and highly adaptive) optimal policy (theorem 3.2).
Our algorithm formulates an LP for the MSKC problem where each pair $(i,j)$ has weight equal to the utility from executing job $i$ at execution level $j$, and size equal to the expected size of job $i$ at execution level $j$. In Dean et al. [10], a similar LP upper-bounds the utility of the optimal adaptive policy and can be approximated via the best non-adaptive policy. In our setting, the LP remains an upper bound on the optimal adaptive utility, but it may be a factor $\Omega(\alpha)$ away (see Section D.3). We show that some non-adaptive ordered set of job-level pairs achieves an $O(\alpha)$-approximation to the LP value.
We construct this set as follows. First, we efficiently obtain a well-structured LP solution in which every job is assigned integrally to at most one execution level, except possibly one job that may be split as a convex combination of two execution levels. We start from an arbitrary LP solution and shift weight along each job’s Pareto curve, iteratively making jobs integral while preserving the objective value. Given the resulting solution, we either apply the standard bang-per-buck ordering to the selected jobs, or take a single job if its expected utility exceeds a fixed threshold. The threshold is chosen to balance the potential loss when overflowing and the utility when selecting a high-utility job that is not part of the LP solution.
Lower Bound.
We show the existence of an instance such that any non-adaptive policy of the principal is bounded away by a factor of $\Omega(\alpha)$ in terms of utility from the optimal adaptive policy. This establishes a tight $\Theta(\alpha)$ adaptivity gap for the Knapsack Contracts problem (theorem 3.1). In fact, in the given instance, the gap exists between non-adaptive and stopping-adaptive policies.
1.1.2 A Constant Approximation with Limited Adaptivity
Motivated by the dependency of non-adaptive policies on ROI, in Section 4, we consider policies that go beyond non-adaptive policies, but still exhibit limited adaptivity. Our goal is to devise simple near-optimal policies, irrespective of the IOR parameter.
In Section 4.1 we first inspect whether policies that take a fixed, non-adaptive execution order, but can adaptively halt execution when future jobs are non-profitable, approximate the optimal utility. This is arguably the simplest class of policies that goes beyond non-adaptive policies, and can avoid incurring a high cost if the remaining time is insufficient. We show that this does not suffice; namely, we provide an instance where the ratio between the optimal utility and the utility of stopping-adaptive policies is $\Omega(\alpha^{1/3})$. In this instance, a good order of execution is obvious, but an optimal policy can skip over jobs that yield negative utility. This motivates us to study the next class of policies.
In Section 4.2 we give an efficient algorithm to compute a skipping-adaptive policy that gives a constant approximation to the utility of the optimal fully-adaptive policy. The policy is computed via a global time-indexed LP, an approach taken in [20, 24, 3], adapted to our setting.
1.1.3 Non-Stopping Policies
In Appendix F, we study an even more demanding setup, where the policy chooses an ordered set of items in advance, and incurs the cost of all of them, not only items that are inserted before the capacity is met. In the motivating example, this corresponds to a contractor who pre-commits to a fixed slate of subcontractors and must pay all agreed costs, even if earlier overruns make later jobs unprofitable or irrelevant. In the setting without costs (e.g., [10]), non-stopping and non-adaptive policies coincide. With costs, however, we show an $\Omega(\alpha)$ separation between them, and give an $O(\alpha)$-approximation non-stopping algorithm for the fully-adaptive optimum.
1.1.4 Information Gap
Our algorithm for computing a constant-approximation policy uses a substantial amount of distributional information, while our non-adaptive algorithm only uses the first moments of the size distributions and the overflow probability. In Appendix C, we establish an information gap: Consider any policy (adaptive or not) that has limited information about the jobs’ size distributions in the form of finitely many moments, as well as access to the probability of knapsack overflow for a single item. No such policy can guarantee an approximation factor better than $O(\alpha)$ with respect to the utility of the optimal policy with full access to the size distributions (theorem C.2). We note that the $O(\alpha)$ non-adaptive policy only uses the expectation (first moment information) and probability of overflow.
1.1.5 Relaxations.
In Appendix G, we study mild relaxations under which simpler policies achieve a constant-factor approximation. In Section G.2, we allow the IOR parameter to be arbitrarily large, but assume that for each job, the probability of consuming more than $1-\delta$ of the budget is small, for $\delta\in(0,1)$. We give an efficient algorithm that outputs a stopping-adaptive policy with an $O(1/\delta)$-approximation to the optimal fully-adaptive utility. We also show that non-adaptive policies fail to achieve an $\alpha$-independent approximation, even when job sizes are bounded by half the total time.
To prove this, we first study a variant where if a job completes after the deadline, the policy still receives its value and the process stops. For this variant, the algorithm from Section 3.2 gives a constant approximation to the LP value. We then run the same algorithm while treating the deadline as a $\delta$-fraction of the true deadline. In this view, bounded-size jobs correspond exactly to collecting the value of the first job that finishes after the (imagined) deadline, and shrinking the deadline costs only a factor $\delta$ in the objective.
2 Model
In this section we introduce our problem, the Inverse ROI parameter, and our policy hierarchy.
2.1 The Multi-choice Stochastic Knapsack with Costs Problem (MSKC)
We generalize Stochastic Knapsack along two axes: (i) items can have multiple tiers, called choices, and (ii) each choice is associated with a cost. An instance $I=\{v_{i},\{p_{i}^{j},F_{i}^{j}\}_{j\in[m]}\}_{i\in[n]}$ of the multi-choice stochastic knapsack problem with costs (or MSKC for short), consists of $n$ items. Each item $i\in[n]$ has a value $v_{i}$ and $m$ choices. Each choice $j\in[m]$ has a (deterministic) cost $p_{i}^{j}$, and a size distribution $F_{i}^{j}$. There is also a knapsack capacity of size $S$, but without loss of generality, $S=1$.
When item $i$ is selected for the knapsack, a single choice $j$ must be specified. We refer to these as the item-choice pair (sometimes referring just to the item when the choice is clear from context). In this case, the cost $p_{i}^{j}$ is incurred up-front, and the item’s size $s_{i}$ is realized from distribution $F_{i}^{j}$. Item $i$’s realized utility is then $v_{i}-p_{i}^{j}$ (more precisely, this is the “potential” realized utility, i.e., the utility that will be gained if the item fits in the knapsack).
For a given instance $I$, a policy $\varphi$ possibly inspects the knapsack, then either selects an item-choice pair, or terminates. Item-choice pairs are selected according to the policy sequentially, until the total size of selected items exceeds the capacity, or until the policy decides to terminate. In the former case, we refer to the last selected item as the overflow item.
The realized utility of the policy is the total realized utility of items selected before the cumulative size exceeded the capacity, minus the cost of the overflow item, if one exists. Let $\mathrm{utility}(\varphi)$ be a random variable representing the utility obtained by the policy $\varphi$ on instance $I$. We also design algorithms that return policies. When referring to an algorithm $\mathcal{A}$ that is given an instance $I$ and returns a policy $\varphi_{I}$, we denote by $\mathcal{A}(I)=\mathrm{utility}(\varphi_{I})$ the random variable representing the utility obtained by the policy on instance $I$.
We are also interested in the expected utility of a pair $(i,j)$ of item and its choice. This depends on the remaining capacity $t\in[0,1]$, and is equal to $v_{i}\cdot\Pr_{s_{i}\sim F_{i}^{j}}[s_{i}\leq t]\;-\;p_{i}^{j}.$ Let
$w_{i,j}:=v_{i}\cdot\Pr_{s_{i}\sim F_{i}^{j}}[s_{i}\leq 1]-p_{i}^{j}$
be the expected utility of inserting $(i,j)$ when the knapsack is empty ($t=1$).
The lower bounds in this paper critically exploit instances with item-choice pairs $(i,j)$, which have high values and high costs, but only low expected utilities:
$0<w_{i,j}\ll\min\{v_{i},p_{i}^{j}\}.$
In such cases, an item-choice pair may yield a small positive expected utility when the knapsack is empty, yet incur a large negative expected utility, of up to $p_{i}^{j}$, when the capacity becomes limited. We show that this drastic instability property is an inherent obstacle in our model. To do so, we define the Inverse of the Return on Investment parameter:
Definition 2.1 (ROI and $\alpha$ Parameters).
We define
$\mathrm{ROI}(i,j)\ :=\ w_{i,j}/p_{i}^{j},~~~\alpha\ :=\ \max_{i,j}\{1/\mathrm{ROI}(i,j)\}.$
We also refer to $\alpha$ as the Inverse of ROI (IOR) parameter.
Intuitively, a large value of $\alpha$ signifies the presence of at least one item-choice pair where the upfront cost of the item greatly exceeds the utility from successfully inserting the item. Our results show that $\alpha$ is the key parameter governing the adaptivity gaps between different levels of adaptivity, which we now specify.
2.2 Adaptiveness Hierarchy
For a fixed instance, recall that a policy is applied iteratively, possibly observing the state of the knapsack in each iteration, in order to sequentially fill the knapsack. After the formal definitions we give an informal interpretation.
Definition 2.2.
(Non-Adaptive Policy) A non-adaptive policy $\varphi:[n]\rightarrow[n]\times[m]\cup\{\perp\}$ takes as input the iteration number, and returns an item-choice pair, or $\perp$ for termination. Let $\mathrm{NON\_ADAPT}(I)$ be the optimum expected utility obtained by a non-adaptive policy for instance $I$.
Definition 2.3.
(Stopping-Adaptive Policy) A stopping-adaptive policy $\pi:[n]\times[0,1]\rightarrow[n]\times[m]\cup\{\perp\}$ takes as input the iteration number and the remaining knapsack capacity, and returns an item-choice pair, or $\perp$ for termination.
Let $\mathrm{STOP\_ADAPT}(I)$ be the optimum expected utility obtained by a stopping-adaptive policy for instance $I$.
Definition 2.4.
(Skipping-Adaptive Policy) A skipping-adaptive policy $\pi:[n]\times[n]\times[0,1]\rightarrow[n]\times[m]\cup\{\perp\}$ takes as input the iteration number, the last item in the knapsack and the remaining knapsack capacity, and returns an item-choice pair, or $\perp$ for termination. Let $\mathrm{SKIP\_ADAPT}(I)$ be the optimum expected utility obtained by a skipping-adaptive policy for instance $I$.
Definition 2.5.
(Fully-Adaptive Policy). A fully-adaptive policy $\varphi:[n]\times 2^{[n]}\times[0,1]\rightarrow[n]\times[m]\cup\{\perp\}$ takes as input the iteration number, the item set in the knapsack and the remaining knapsack capacity, and returns an item-choice pair, or $\perp$ for termination. Let $\mathrm{ADAPT}(I):=\max_{\varphi}\operatorname{\mathbb{E}}[\mathrm{utility}(\varphi(I))]$ be the optimum expected utility obtained by a fully-adaptive policy for instance $I$.
Intuitively, a non-adaptive policy is an ordering of items. For each $i\in[n]$, it specifies the $i$-th selected item and its associated choice, or $\perp$ for termination. A stopping-adaptive policy, like a non-adaptive one, is an ordering of items, but it can adaptively stop at any point. A skipping-adaptive policy is like a stopping-adaptive policy, with the additional ability to skip an item before stopping. It, too, is an ordering of items, and given the last inserted item it scans the successive ones in the ordering and chooses the next one to insert. [^1] Related to skipping-adaptive policies, we can consider an ordering of items and target capacities — the maximum remaining capacity with which each item will be entered into the knapsack. When given the last item inserted, the policy finds the subsequent item in the ordering, whose target capacity is below the remaining knapsack capacity. [10] study fully-adaptive (and non-adaptive) policies.
We remark that each policy can be represented by a decision tree; a fully-adaptive policy needs an exponentially-sized tree, but the other policies require only poly-sized trees. In fact stopping-adaptive and non-adaptive policies correspond to linear-sized decision trees.
3 Tight $\Theta(\alpha)$ Adaptivity Gap
In this section we establish a tight $\Theta(\alpha)$ gap between simple non-adaptive policies and the optimal fully-adaptive policy. Theorem 3.1 constructs an instance with an $\Omega(\alpha)$ gap between the best non-adaptive policy and the best stopping-adaptive policy. Theorem 3.2 gives a polynomial-time algorithm that outputs a non-adaptive policy whose expected utility is within an $O(\alpha)$ factor of the optimal fully-adaptive utility.
In Appendix F, we study an even more demanding setup, where the policy chooses an ordered set of jobs in advance, and incurs the cost of all of them, not only jobs that are inserted before the capacity is met. In the setting without costs (e.g., [10]), this setup is equivalent to a non-adaptive policy, as all costs are 0. In the MSKC Problem, we show a separation between the two. In Appendix F we define non-stopping policies, and show that there is an instance with an $\Omega(\alpha)$ gap between the optimal non-stopping policy and the optimal non-adaptive policy, and an $O(\alpha)$ approximation non-stopping algorithm to the optimal fully-adaptive policy.
3.1 Lower Bound
We show that there exists an $\Omega(\alpha)$ gap between the optimal Non-Adaptive policy and the optimal Stopping-Adaptive policy. Recall that both models require the algorithm to choose an item ordering at the beginning; the stopping-adaptive policy can then adaptively choose when to halt, while the non-adaptive policy must attempt all items in the chosen order, until all items are inserted or an overflow occurs.
Theorem 3.1.
For every $\alpha$ there exists an instance $I$ of the MSKC problem such that $\frac{\mathrm{STOP\_ADAPT}(I)}{\mathrm{NON\_ADAPT}(I)}=\Omega(\alpha).$
We give a proof sketch here and defer the full proof of theorem 3.1 to section D.1.
Proof sketch.
Fix parameters $\varepsilon,\gamma\in(0,\tfrac{1}{2})$ and let $n=\lceil 1/\varepsilon\rceil$. Consider capacity $1$ and $n$ items with value $v=2$ and cost $p=2-\varepsilon$, so each successful insertion yields net gain $w=v-p=\varepsilon$ and hence $\alpha=\max_{i}p/w=(2-\varepsilon)/\varepsilon=\Theta(1/\varepsilon)$. For item $i\in[n]$, define a two-point size distribution
$s_{i}=\begin{cases}a_{i}&\text{with prob. }1-\varepsilon,\\ b_{i}&\text{with prob. }\varepsilon,\end{cases}\qquad a_{i}=\gamma^{\,n-i+1},\qquad b_{i}=1-\sum_{j<i}a_{j}.$
Since $\gamma<\tfrac{1}{2}$, the sequence $(a_{i})$ grows geometrically and satisfies $a_{i^{\prime}}>\sum_{j<i}a_{j}$ for all $i<i^{\prime}$, implying $b_{i}+a_{i^{\prime}}>1$. Thus, once any earlier item realizes its “big” size $b_{i}$, every subsequent attempt overflows even if it realizes “small”. A stopping-adaptive policy that attempts items in order and halts after the first big realization inserts a geometric number $X$ of items (truncated at $n$) with
$\mathbb{E}[X]=\sum_{t=1}^{n}(1-\varepsilon)^{t-1}=\frac{1-(1-\varepsilon)^{n}}{\varepsilon}\geq\frac{1-e^{-1}}{\varepsilon},$
and therefore achieves expected utility at least $\mathbb{E}[X]\cdot\varepsilon\geq 1-e^{-1}=\Omega(1)$.
In contrast, we show (in the full proof) that the expected utility of any non-adaptive policy that attempts inserting at least $k>1$ items has negative utility. Hence the optimal non-adaptive policy attempts only one item and attains $\mathrm{NON\_ADAPT}=\varepsilon$. Consequently,
$\frac{\mathrm{STOP\_ADAPT}}{\mathrm{NON\_ADAPT}}\geq\frac{1-e^{-1}}{\varepsilon}=\Omega(\alpha),$
as claimed. ∎
3.2 Upper Bound
In this section, we present an $O(\alpha)$-approximate non-adaptive policy for the MSKC problem. Our main theorem is the following:
Theorem 3.2.
There exists a polynomial time algorithm for MSKC that computes for every input $I$ a non-adaptive policy $\pi_{I}:[n]\to[n]\times[m]\cup\{\perp\}$ such that $\frac{\mathrm{ADAPT}(I)}{\operatorname{\mathbb{E}}[\mathrm{utility}(\pi_{I})]}\leq O(\alpha).$
We first define an LP that both guides our solution and gives an upper bound on the performance of every adaptive policy. For every $i\in[n],j\in[m]$ and for a knapsack of size $t$, denote by $w_{ij}=v_{i}\cdot\Pr_{{s_{i}^{j}\sim F_{i}^{j}}}[s_{i}^{j}\leq t]-p_{i}^{j}$ and $\mu_{ij}=\operatorname{\mathbb{E}}_{s_{i}^{j}\sim F_{i}^{j}}[\min\{s_{i}^{j},t\}]$. The first parameter can be viewed as a proxy to the utility from item $i$ inserted with choice $j$ and the second parameter is the expected truncated size.
Consider the following LP, parameterized by the size of the knapsack $t$.
$\Phi(t)=\max\left\{\sum_{ij}w_{ij}\,x_{ij}\;:\;\sum_{ij}\mu_{ij}\,x_{ij}\,\leq\,t,\quad\forall\,i,\;\sum_{j}x_{ij}\,=\,1,\quad x_{ij}\,\geq\,0\right\}.$ (1)
We assume, without loss of generality, that every item $i$ includes a dummy “null” choice $j_{\bot}$ with zero size and zero value; selecting this choice represents the case where none of the actual choices for the item are chosen. In addition, we will use the solution to this LP to guide our algorithm. First, we characterize the structure of optimal solutions:
Lemma 3.3.
There exists an optimal solution to $\Phi(t)$ satisfying the following properties:
  1. 1.
    At most one item $i^{*}$ has exactly two choices $j,j^{\prime}$ such that $x_{i^{*}j},x_{i^{*}j^{\prime}}>0$.
  2. 2.
    Every other item $i\neq i^{*}$ has at most one choice $j$ with $x_{ij}=1$.
Moreover, such a solution can be computed in polynomial time.
Note that every solution returned by the Simplex algorithm satisfies the above properties; the basic feasible solution of the LP has at most $n+1$ non-negative variables, and the constraints ensure that for each item $i$, there is at least one choice $j$ with $x_{ij}>0$, hence at most one item has two non-zero choices. However, the Simplex algorithm may run in exponential time in the worst case. Thus, we describe a polynomial time algorithm to compute such a solution.
First, for each item $i$, we characterize the choices $j$ for which there exists an optimal solution with $x_{ij}=0$.
We then show that there always exists an optimal solution in which at most two choices have positive values.
Definition 3.1 (Single Dominance).
For an item $i$, choice $j^{\prime}$ single dominates choice $j$ if:
$(1)\quad w_{ij^{\prime}}\geq w_{ij}\qquad(2)\quad\mu_{ij^{\prime}}\leq\mu_{ij}$
Definition 3.2 (Double Dominance).
For an item $i$, choices $j^{\prime}$ and $j^{\prime\prime}$ double dominate choice $j$ if:
$(1)$ $\quad w_{ij^{\prime}}<w_{ij}<w_{ij^{\prime\prime}}$
$(2)$ $\quad\mu_{ij^{\prime}}\leq\mu_{ij}\leq\mu_{ij^{\prime\prime}}$
$(3)$ $\quad(\mu_{ij}-\mu_{ij^{\prime}})/(w_{ij}-w_{ij^{\prime}})\geq(\mu_{ij^{\prime\prime}}-\mu_{ij})/(w_{ij^{\prime\prime}}-w_{ij})$
Finally, we say that a choice $j$ is dominated if it is either single-dominated or double-dominated.
Lemma 3.4.
There exists an optimal solution $x$ to $\Phi(t)$ in which every dominated choice is assigned zero weight; namely, for any item $i$ and any choice $j$ that is either single‑dominated or double‑dominated, we have $x_{ij}=0$.
Proof.
Suppose that for some feasible solution $x$, we have $x_{ij}\neq 0$.
First, if $j^{\prime}$ single-dominates $j$, define a new solution by setting $x^{*}_{ij^{\prime}}=x_{ij^{\prime}}+x_{ij},\ x^{*}_{ij}=0.$
This yields a feasible solution whose objective value is at least as large as that of $x$. Hence, there exists an optimal solution in which $x_{ij}=0$.
Second, if $j^{\prime}$ and $j^{\prime\prime}$ double-dominate $j$, define a new solution by setting
$x^{*}_{ij^{\prime}}=x_{ij^{\prime}}\;+\;x_{ij}\cdot\frac{w_{ij^{\prime\prime}}-w_{ij}}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}},\quad x^{*}_{ij^{\prime\prime}}=x_{ij^{\prime\prime}}\;+\;x_{ij}\cdot\frac{w_{ij}-w_{ij^{\prime}}}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}},\quad x^{*}_{ij}=0~.$
This again yields a feasible solution whose objective value is at least as large as that of $x$ (see fig. 2 for illustration, and section D.2 for the full proof). Therefore, in either case, there exists an optimal solution in which $x_{ij}=0$.
[…]
Figure 2: Visualization of the removal of double-dominated points. The hollow circles ($j^{\prime},j^{\prime\prime}$) represent two choices for an item. The red point ($j$) is ”double-dominated” because it lies above the line segment representing the convex combination of $j^{\prime}$ and $j^{\prime\prime}$. In the LP relaxation, this dominated point can be replaced by its vertical projection (the green point) onto the segment, which yields a solution that is no worse in objective value and potentially better in terms of feasibility ($\mu$).
Lemma 3.5.
There exists
an optimal solution $x$ to $\Phi(t)$ in which, for every item $i$, at most two choices satisfy $x_{ij}>0$.
Proof.
Let $x$ be an optimal solution in which no choice is dominated; such a solution exists by Lemma 3.4. Without loss of generality, we may further assume that no additional choices remain and that for every $i\in[n]$ and $j\in[m-1]$, $w_{i,j}<w_{i,j+1}$. Lemma 3.4 then also yields $\mu_{i,j}<\mu_{i,j+1}$.
Consider the pairs $\{\bigl(w_{ij},\mu_{ij}\bigr)\}_{j\in[m]}$ for a fixed index $i$. Connecting consecutive points by straight-line segments (the standard piecewise-linear interpolation) yields a polyline with non-decreasing slopes. Consequently, the boundary is convex.
Suppose for some item $i$ there are strictly more than two choices that satisfy $x_{ij}>0$.
Define $\tilde{\mu}=\sum_{j}x_{ij}\mu_{ij}$. Let $j^{\prime}=\arg\max_{j}\bigl\{\mu_{ij}\leq\tilde{\mu}\bigr\}$ and $j^{\prime\prime}=\arg\min_{j}\bigl\{\mu_{ij}\geq\tilde{\mu}\bigr\}$. By the definition of $\tilde{\mu}$ and by our assumption that at least two choices satisfy $x_{ij}>0$, such indices $j^{\prime}$ and $j^{\prime\prime}$ always exist. Set $\lambda=\frac{\mu_{ij^{\prime\prime}}-\tilde{\mu}}{\mu_{ij^{\prime\prime}}-\mu_{ij^{\prime}}}$. Consider the solution defined by $x^{*}_{ij^{\prime}}=\lambda$, $x^{*}_{ij^{\prime\prime}}=1-\lambda$, and $x^{*}_{ij}=0$ for all $j\notin\{j^{\prime}j^{\prime\prime}\}$. Note that this construction satisfies $\sum_{j}x^{*}_{ij}\,\mu_{ij}=\tilde{\mu}$. Therefore, it yields a feasible solution. Since the resulting point lies on the boundary of the convex set, its objective value is at least as large as that of the original solution $x$, (see fig. 3 for illustration).
[…]
Figure 3: Visualization of the set of undominated points, $\Gamma_{i}$. The blue line represents the efficient (Pareto) frontier formed by undominated choices and their convex combinations. An arbitrary combination of choices that lies above this frontier (the red point) is inefficient. It can be strictly improved by projecting it horizontally onto the frontier (the green point), which achieves the same $\mu$ value for a greater $w$ value.
We are now ready to prove Lemma 3.3.
Proof of Lemma 3.3.
By Lemmas 3.4, 3.5, it follows that there exists an optimal solution $x$ where at most two choices have positive values.
Consider a solution where there exist indices $i_{1},i_{2}$ and $j^{\prime}_{1},j^{\prime\prime}_{1},j^{\prime}_{2},j^{\prime\prime}_{2}$ such that $x_{i_{k},j^{\prime}_{k}}>0$ and $x_{i_{k},j^{\prime\prime}_{k}}>0$ for each $k\in\{1,2\}$. Let
$d_{k}=\frac{w_{i_{k},j^{\prime}_{k}}-w_{i_{k},j^{\prime\prime}_{k}}}{\mu_{i_{k},j^{\prime}_{k}}-\mu_{i_{k},j^{\prime\prime}_{k}}}$
be the current slope of the $k$-th item, and assume without loss of generality that $d_{1}\geq d_{2}$. Define
$\tau=\frac{\mu_{i_{1},j^{\prime}_{1}}-\mu_{i_{1},j^{\prime\prime}_{1}}}{\mu_{i_{2},j^{\prime}_{2}}-\mu_{i_{2},j^{\prime\prime}_{2}}}.$
For $\epsilon>0$, set
$x^{*}_{i_{1},j^{\prime}_{1}}=x_{i_{1},j^{\prime}_{1}}+\epsilon,\quad x^{*}_{i_{1},j^{\prime\prime}_{1}}=x_{i_{1},j^{\prime\prime}_{1}}-\epsilon,$
$x^{*}_{i_{2},j^{\prime}_{2}}=x_{i_{2},j^{\prime}_{2}}-\tau\cdot\epsilon,\quad x^{*}_{i_{2},j^{\prime\prime}_{2}}=x_{i_{2},j^{\prime\prime}_{2}}+\tau\cdot\epsilon.$
By construction, for any $\epsilon>0$ we have
$\sum_{j}x_{i_{1},j}\,\mu_{i_{1},j}+\sum_{j}x_{i_{2},j}\,\mu_{i_{2},j}=\sum_{j}x^{*}_{i_{1},j}\,\mu_{i_{1},j}+\sum_{j}x^{*}_{i_{2},j}\,\mu_{i_{2},j}.$
By choosing the smallest $\epsilon$ for which $x^{*}_{i_{1},j^{\prime}_{1}}=1$, or $x^{*}_{i_{2},j^{\prime}_{2}}=0$, or $x^{*}_{i_{1},j^{\prime\prime}_{1}}=0$, or $x^{*}_{i_{2},j^{\prime\prime}_{2}}=1$, we obtain a feasible solution whose objective value is at least as large as that of the original solution $x$, and whose number of items with two positive choices is strictly smaller. By iteratively applying this procedure, we eventually obtain a solution in which at most a single item has two positive choices.
Finally, we observe that all these transformations are applicable in polynomial time. Therefore, given any solution to $\Phi(t)$, it is possible to maintain a solution satisfying the above properties in polynomial time.
[…]
Figure 4: For an optimal solution $x$, items $i_{1}$ and $i_{2}$ each have two choices with nonzero values. The solution for each item is a convex combination of the two choices, corresponding to the red dots in the graphs $\Gamma_{i_{1}}$ and $\Gamma_{i_{2}}$, which lie on the segments connecting the two choices. As the slope of the first item is larger, the modified solution indicated by the dashed arrows illustrates one possible update that maintains feasibility and increases the objective value.
For a set of item choices $\mathcal{S}\subset[n]\times[m]$, denote $size(\mathcal{S})=\sum_{(i,j)\in\mathcal{S}}s_{i}^{j}$ and $\mu(S)=\sum_{(i,j)\in\mathcal{S}}\mu_{ij}$. The following Lemmas are direct extensions of key lemmas from Dean et al. [10], originally established for the classic Stochastic Knapsack problem with a capacity of 1. We adapt these lemmas to our knapsack variant with multi-choice and costs, and for a general capacity $t$. In this section, we only need the case where $t=1$; however, in appendix G we use the generalized versions. The proofs are presented in Appendix D for completeness.
Lemma 3.6.
For any set of item choices $\mathcal{S}$, $\Pr[size(\mathcal{S})<t]\geq 1-\frac{1}{t}\mu(\mathcal{S}).$
Lemma 3.7.
For any adaptive policy for the MSKC problem with capacity $t$, let $\mathcal{S}$ be the random set of item choices that the policy inserts, then $\operatorname{\mathbb{E}}[\mu(\mathcal{S})]\leq 2t$.
Lemma 3.8.
For any constant $c\geq 2$ and capacity $t$, $\mathrm{ADAPT}\leq\Phi(ct)\leq c\cdot\Phi(t),$ where $\mathrm{ADAPT}$ is the optimal adaptive policy for capacity $t$.
We now present our algorithm, NA-SKC (Non-Adaptive Stochastic Knapsack with Costs):
Input: Instance II with items $i\in[n]$, choices $j\in[m]$, values $v_{i}$, costs $p_{i}^{j}$, distributions $F_{i}^{j}$
Output: Non-adaptive policy $\pi$
// Step 1: LP Solution
Solve linear program $\Phi(1)$ to obtain optimal solution $x$ with the structure from Lemma 3.3 Let $\mathcal{C}=\{(i,j)\mid x_{ij}>0\}$ be the candidate item-choice pairs
// Step 2: Virtual Item Construction
if $\exists$ item $i^{*}$ with two choices $j,j^{\prime}$ such that $x_{i^{*},j},x_{i^{*},j^{\prime}}>0$ then
   Remove $(i^{*},j)$ and $(i^{*},j^{\prime})$ from $\mathcal{C}$ Create virtual item $(i^{*},j^{*})$ with:
 $w_{i^{*},j^{*}}\leftarrow w_{i^{*},j}x_{i^{*},j}+w_{i^{*},j^{\prime}}x_{i^{*},j^{\prime}}$
 $\mu_{i^{*},j^{*}}\leftarrow\mu_{i^{*},j}x_{i^{*},j}+\mu_{i^{*},j^{\prime}}x_{i^{*},j^{\prime}}$
   Add $(i^{*},j^{*})$ to $\mathcal{C}$ When attempting to insert $(i^{*},j^{*})$, choose $j$ with probability $x_{i^{*},j}$ and $j^{\prime}$ with probability $x_{i^{*},j^{\prime}}$
// Step 3: Item Insertion with Threshold Check
$w_{\max}\leftarrow\max_{(i,j)\in\mathcal{C}}w_{ij}$
$\mathcal{T}\leftarrow\frac{\Phi(1)}{4(1+\alpha)}$
if $w_{\max}>\mathcal{T}$ then
   Insert item $(i_{\max},j_{\max})=\arg\max_{(i,j)\in\mathcal{C}}w_{ij}$ return policy that executes only this item
else
 return policy that executes items in $\mathcal{C}$ in decreasing order of density $\frac{w_{ij}}{\mu_{ij}}$
Algorithm 1 NA-SKC (Non-Adaptive Stochastic Knapsack with Costs)
An overview of the algorithm.
We begin by solving $\Phi(1)$ to obtain an optimal solution $x$ with the properties established in Lemma 3.3: each item $i$ has one choice $j$ such that $x_{ij}=1$, except for possibly one item $i^{*}$ with two choices $j,j^{\prime}$ such that $x_{i^{*}j}+x_{i^{*}j^{\prime}}=1$. Let $\mathcal{C}$ denote the set of item-choice pairs selected by the LP, and let $r=|\mathcal{C}|$. If there exists an item with two positive choices, we remove these two choices and replace them with a single virtual choice; whenever the algorithm attempts to insert this virtual item, it randomly selects between the two original choices with probabilities proportional to their values in $x$. Since each item now has a single (possibly virtual) choice, we refer to each item-choice pair $(i,j)$ by the item $i$. Note that from the properties of our solution $x$, it holds that $\sum_{k=1}^{r}{w_{k}}=\Phi(1)$ and $\sum_{k=1}^{r}{\mu_{k}}\leq 1$. We then define a threshold $\mathcal{T}(\alpha)$ according to which we decide whether to insert the item with the maximum utility $w_{\text{max}}$. If the threshold is not met, we insert the items in decreasing order of their density $\frac{w_{i}}{\mu_{i}}$, until the knapsack is full or no more items remain in the candidate set $\mathcal{C}$. Our threshold determines whether or not following the greedy ordering is worth the risk of overflow compared to the alternative of inserting the best item. The bigger $w_{\text{max}},\alpha$ are, the more likely we are to insert the best item.
Note that this algorithm can be easily de-randomized by choosing the choice among the two choices that maximizes the expected gain.
Claim 3.9.
$\mathbb{E}[\textbf{NA-SKC}]\geq\mathcal{T}=\frac{\Phi(1)}{4(1+\alpha)}$
Proof.
First, if $w_{\max}>\mathcal{T}$, the algorithm attempts only item $(i_{\max},j_{\max})$, gaining an expected net utility of $w_{\max}$. Since $w_{\max}>\mathcal{T}$, we have $\mathbb{E}[\textbf{NA-SKC}]=w_{\max}>\mathcal{T}$, as required.
Now, consider the case $w_{\max}\leq\mathcal{T}$. Define $\textbf{NA-SKC}_{k}$ as the contribution of item $k$ to the algorithm’s utility. It holds that $\operatorname{\mathbb{E}}[\textbf{NA-SKC}]=\sum_{k=1}^{r}{\operatorname{\mathbb{E}}[\textbf{NA-SKC}_{k}]}.$ We want to lower bound $\operatorname{\mathbb{E}}[\textbf{NA-SKC}_{k}]$. If item $k$ is inserted and does not overflow, we obtain a utility of $v_{k}-p_{k}\geq w_{k}$. On the other hand, if the item overflows, we pay a cost of $p_{k}$. By the definition of the IOR parameter (Definition 2.1), $\alpha\geq\frac{p_{k}}{v_{k}\Pr[s_{k}\leq 1]-p_{k}}=p_{k}/w_{k}$, which implies that an overflowing item incurs a cost of $p_{k}\leq\alpha w_{k}\leq\alpha w_{\max}$.
We get that
$\operatorname{\mathbb{E}}[\textbf{NA-SKC}_{k}]$ $=(v_{k}-p_{k})\cdot\Pr[k\mbox{ successfully inserted}]-p_{k}\cdot\Pr[k\mbox{ overflows}]$
$\geq(v_{k}-p_{k})\left(1-\sum_{j=1}^{k}{\mu_{j}}\right)-p_{k}\cdot\Pr[k\mbox{ overflows}]$
$\geq w_{k}\left(1-\sum_{j=1}^{k}{\mu_{j}}\right)-\alpha\cdot w_{\max}\cdot\Pr[k\mbox{ overflows}],$ (2)
where the first inequality follows from lemma 3.6 (with $t=1$), the second inequality follows from the definition of $w_{k}\leq v_{k}-p_{k}$ and the argument above that $p_{k}\leq\alpha\,w_{\max}$.
In the above, this bound is also true for the virtual item $i^{*}$: the item’s expected truncated size is $\mu_{i^{*}}$, and its expected utility conditioned on the event that it was successfully inserted is at least $w_{i^{*}}$.
Summing over all $k\in[r]$, we have that
$\operatorname{\mathbb{E}}[\textbf{NA-SKC}]$ $\geq\sum_{k=1}^{r}{w_{k}\left(1-\sum_{j=1}^{k}{\mu_{j}}\right)}-\alpha w_{\max}\sum_{k=1}^{r}{\Pr[k\mbox{ overflows}]}$
$\geq\Phi(1)-\sum_{1\leq j\leq k\leq r}{w_{k}\mu_{j}}-\alpha w_{\max}$ $(*)$
$=\Phi(1)-\sum_{1\leq j<k\leq r}{w_{k}\mu_{j}}-\sum_{k=1}^{r}{w_{k}\mu_{k}}-\alpha w_{\max}$
$\geq\Phi(1)-\frac{1}{2}\sum_{k,j=1}^{r}{w_{k}\mu_{j}}-\sum_{k=1}^{r}{w_{k}\mu_{k}}-\alpha w_{\max}$ $(**)$
$\geq\frac{\Phi(1)}{2}-\sum_{k=1}^{r}{w_{k}\mu_{k}}-\alpha w_{\max}$ $\sum_{k,j=1}^{r}w_{k}\mu_{j}\leq\Phi(1)$
$\geq\frac{\Phi(1)}{2}-w_{\max}(1+\alpha)\geq\frac{\Phi(1)}{4}\geq\mathcal{T}~,$ (3)
In ()(*) we use the fact that $\sum_{k\in[r]}\Pr[k\mbox{ overflows}]\leq 1$ as at most one item overflows, and $\sum_{k=1}^{r}{w_{k}}=\Phi(1)$. In ()(**) we use the decreasing order of density: for every $j<k$ it holds that $w_{k}\mu_{j}\leq w_{j}\mu_{k}$, and thus $w_{k}\mu_{j}\leq\frac{1}{2}\left(w_{k}\mu_{j}+w_{j}\mu_{k}\right)$. In the final two inequalities we use the fact that $\sum_{k=1}^{r}\mu_{k}\leq 1$ and $w_{\max}\leq\mathcal{T}=\frac{\Phi(1)}{4(1+\alpha)}$.
We are ready to prove the main theorem of this section:
Proof of theorem 3.2.
From 3.9 and lemma 3.8, we conclude:
$\frac{\mathrm{ADAPT}}{\operatorname{\mathbb{E}}[\textbf{NA-SKC}]}\leq\frac{2\Phi(1)}{\frac{\Phi(1)}{4(1+\alpha)}}=8(1+\alpha)=O(\alpha)~.$
In our proof, we show that our algorithm finds a policy that approximates the LP value up to a factor $O(\alpha)$. In [10], a similar LP helped guide the solution as its value was a proxy (up to constant factors) of the optimal policy for the stochastic knapsack problem. In our case, this LP provides an upper bound to the optimal policy (see Lemma 3.8), but its value can be larger by a factor $\Omega(\alpha)$ than the utility of the optimal policy (see Section D.3). This shows that we cannot hope for a tighter approximation to the LP’s value.
4 A Constant Approximation with Limited Adaptivity
In this section, we study the power of limited adaptivity, namely skipping-adaptive and stopping-adaptive policies. In Section 4.1, we prove an $\Omega(\alpha^{1/3})$ gap between stopping-adaptive and skipping-adaptive policies. In Section 4.2, we present an $O(1)$-approximation skipping-adaptive algorithm, highlighting the power of adaptively skipping items in a given ordering.
4.1 Lower Bound for Stopping-Adaptive Policies
Recall that the stopping-adaptive model requires the algorithm to fix a non-adaptive ordering of the items at the outset, and it halts if all future sequences of items yield a negative expected utility. This is arguably the simplest mildly-adaptive class of policies. The following theorem shows that allowing the algorithm to adaptively halt is not sufficient to achieve a constant approximation to the optimal fully-adaptive policy in our model. In fact, the gap already exists between stopping-adaptive and skipping-adaptive policies.
Theorem 4.1.
For every $\alpha$ there exists an instance $I$ of the MSKC problem such that
$\frac{\mathrm{SKIP\_ADAPT}(I)}{\mathrm{STOP\_ADAPT}(I)}=\Omega\bigl({\alpha}^{1/3}\bigr).$
In order to prove the above theorem, we define an instance $I$ consisting of an infinite collection of items, each with value $v=15/\epsilon$ and cost $p=15/\epsilon-\epsilon^{2}$. As in Section 3.1, $\epsilon$ is used to determine $\alpha$. Again, as the claim is trivial for bounded $\alpha$, we prove the claim for large values of $\alpha$.
The item set is partitioned into $1/\epsilon$ types of items $L_{1},\ldots,L_{1/\epsilon}$, each containing an infinite number of items. An item of type $L_{i}$ has the size distribution
Probability Size
$1-\epsilon-\epsilon^{2}$ 0
$\epsilon$ $a_{i}$
$\epsilon^{2}$ $b_{i}$
where $a_{i}=\gamma^{\,n-i+1},\qquad b_{i}=1-\displaystyle\sum_{j=1}^{i-1}a_{j}\qquad$ for $\gamma<\tfrac{1}{2}$.
By the definition of the IOR parameter (definition 2.1), $\alpha=\frac{15/\epsilon-\epsilon^{2}}{\epsilon^{2}}=15/\epsilon^{3}-1=O\left(\frac{1}{\epsilon^{3}}\right)$.
For every large enough $\alpha$, we can choose an appropriate $\epsilon$ that sets the IOR to be $\alpha$.
We first analyze the expected utility from a fully-adaptive policy.
Lemma 4.2.
For the instance given above, $\mathrm{SKIP\_ADAPT}(I)\;\geq\;\frac{1}{2e}.$
Proof.
Consider a skipping-policy that selects items by type $L_{1}$, then type $L_{2}$, and so on until type $L_{1/\epsilon}$, according to the following rule:
  • When selecting items of type $L_{i}$, the policy continues to select such items as long as $s_{i}=0$.
  • Once $s_{i}=a_{i}$, the policy moves on to select items of type $L_{i+1}$ if $i<1/\epsilon$.
  • If $s_{i}=b_{i}$, the policy terminates.
Notice that the execution order ensures each inserted item finishes and does not overflow, so the policy obtains a utility of exactly $v-p=\epsilon^{2}$ for every inserted item. Let $X$ be a random variable for the number of items the adaptive policy successfully completes. Since the utility from each item is $\epsilon^{2}$, the policy gets a utility of $\operatorname{\mathbb{E}}[X]\cdot\epsilon^{2}$. Let $X_{i}$ be the number of items of type $L_{i}$ completed by the policy. Notice that
[…]
By the description of the policy, $\operatorname{\mathbb{E}}\bigl[X_{i}\mid\varphi\text{ inserts items of type }L_{i}\bigr]=\frac{1}{\epsilon+\epsilon^{2}}.$
We also have
[…]
The second inequality follows because in order for the policy to reach items of type $L_{1/\epsilon}$, each time an item of type $L_{i}$ (for $i<1/\epsilon$) has its size realized at a value larger than 0, the policy moves on to items of type $L_{i+1}$ only if the item’s size is realized at $a_{i}$, which happens with probability $\frac{\epsilon}{\epsilon+\epsilon^{2}}$. Since this must occur at most $1/\epsilon$ times, the probability is at least $\left(\frac{\epsilon}{\epsilon+\epsilon^{2}}\right)^{1/\epsilon}.$ Therefore, $\operatorname{\mathbb{E}}[X_{i}]\;\geq\;\frac{1}{e(\epsilon+\epsilon^{2})}\;>\;\frac{1}{2e\epsilon},$ which implies $\operatorname{\mathbb{E}}[X]=\sum_{i=1}^{1/\epsilon}\operatorname{\mathbb{E}}[X_{i}]\;>\;\frac{1}{2e\epsilon}\cdot\frac{1}{\epsilon}=\frac{1}{2e\epsilon^{2}}.$ This yields $\mathrm{ADAPT}(I)\;>\;\epsilon^{2}\cdot\frac{1}{2e\epsilon^{2}}\;=\;\frac{1}{2e},$ as desired. ∎
We next bound the expected utility from a stopping-adaptive policy.
Lemma 4.3.
For the instance given above, $\mathrm{STOP\_ADAPT}(I)\;\leq\;4\epsilon.$
Proof.
Consider some sequence $\pi$ of items. Let $Y$ denote the following event: the policy is to insert an item of type $L_{i}$, when some prior item of type $L_{i}$ was processed with a size realized to be larger than 0. We claim that once event $Y$ occurs, then for every sequence of future items, the expected utility is negative. Consider first the expected utility of only inserting the next item, of type $L_{i}$. This is at most
$(1-\epsilon^{2})(v-p)+\epsilon^{2}(-p)=(1-\epsilon^{2})\cdot\epsilon^{2}-\epsilon^{2}(15/\epsilon-\epsilon^{2})=-15\epsilon+\epsilon^{2}<0,$
since if the next item’s size is realized to be $b_{i}$, then the item overflows for sure.
As the utility of any item is at most $v-p=\epsilon^{2},$ this implies that for any sequence to have a positive utility, it has to be of size larger than $(15\epsilon-\epsilon^{2})/\epsilon^{2}$. Consider a partition of the sequence into blocks of size $k=\frac{1}{\epsilon+\epsilon^{2}}+\frac{1}{\epsilon}.$ If the sequence ends at a block of size smaller than $k$, we merge the last two blocks together to form a single block, meaning that the size of each block is between $k$ and $2k$. Denote by $r$ the number of blocks in the sequence. Consider some block $j\in[r]$, and let $n_{j}(i)$ denote the number of items of type $L_{i}$ it contains. From our observation, if $n_{j}(i)>1$, then the expected utility of processing the second item of type $L_{i}$ in the block is at most $15\epsilon+\epsilon^{2}$, in which case the expected utility of the entire block is at most
$15\epsilon+\epsilon^{2}+\epsilon^{2}(2k-1)<-15\epsilon+2\epsilon^{2}+4\epsilon=-11\epsilon+2\epsilon^{2}<0.$
Therefore, to have a positive expected utility, for every $i$, if $n_{j}(i)>1$, then the first $n_{j}(i)-1$ items of type $L_{i}$ must have size 0, and in total, at least $\sum_{i=1}^{1/\epsilon}{\max\{n_{j}(i)-1,0\}}\geq k-\frac{1}{\epsilon}=\frac{1}{\epsilon+\epsilon^{2}}$ items must have size 0. This occurs with probability at most $(1-(\epsilon+\epsilon^{2}))^{\frac{1}{\epsilon+\epsilon^{2}}}<\frac{1}{2},$ in which case the expected utility of the block is at most $2k\epsilon^{2}\leq 4\epsilon$.
Else, with probability at least $1/2$, the expected utility of the block is at most $11\epsilon+2\epsilon^{2}$. Thus, the expected utility of a block is bounded by $\frac{1}{2}(-11\epsilon+2\epsilon^{2})+4\epsilon=-\frac{3}{2}\epsilon+\epsilon^{2}<0.$
Thus, we get that once event $Y$ occurs, since every sequence of items has a negative utility for the policy, it will halt. Let $X$ be a random variable of the number of items inserted before event $Y$ occurs. Since every item has utility at most $\epsilon^{2}$, we have that $\mathrm{STOP\_ADAPT}(I)\leq\operatorname{\mathbb{E}}[X]\epsilon^{2}$. We again break $\pi$ into blocks of size $k$. As analyzed above, for every block, we have that event $Y$ occurs with probability at least $1/2$; therefore, the expected number of blocks processed is bounded by 2, and the expected number of items inserted is bounded by $2k$. Therefore,
$\mathrm{STOP\_ADAPT}(I)\leq\operatorname{\mathbb{E}}[X]\epsilon^{2}\leq 2k\epsilon^{2}\leq 4\epsilon,$
as desired. ∎
Combining Lemmas 4.2 and 4.3 yields Theorem 4.1.
4.2 An $O(1)$ Skipping-Adaptive Policy
theorem 4.1 implies that without skipping items we may not achieve an $\alpha$-independent approximation ratio. In this section we provide an algorithm which finds a skipping-adaptive policy with a constant approximation guarantee. The LP used in Section 3.2 has a $\Omega(\alpha)$ gap to the optimal solution, as shown in Section D.3. In fact, any LP that only uses limited distributional information (e.g. the expected truncated sizes) is insufficient to achieve a constant approximation, as shown in Appendix C. Thus, we turn to a policy that uses more information than just the moments.
The idea is to use a global time-indexed LP as in [20, 24, 3], and design an algorithm that aims to get a constant fraction of its value. In this section, we assume all item sizes (or times) are in $\left\{0,1,\ldots,T\right\}$ and that the knapsack size is $T$ (equivalently, that all items have rational sizes) [^2] In the original problem we assumed that the knapsack size is $1$. Here we rescale all sizes by $T$ since it is more convenient. In section E.3 we discuss the handling of continuous distributions.. In the global time-indexed LP, there is a variable not only for each item-choice $i,j$, but also for each time step $t\in\left\{0,\ldots T\right\}$ (each size is interpreted as a job duration). Thus the decision to insert any item depends also on the time of the decision. For every $i\in[n],\ j\in[m],\ t\in\left\{0,\ldots T\right\}$ let $w_{i,j,t}:=v_{i}\cdot\Pr_{{s_{i}^{j}\sim F_{i}^{j}}}[s_{i}^{j}\leq T-t]-p_{i}^{j}$ be the effective value of choice $i,j$ given remaining capacity of TtT-t and let $\mu_{i,j,t}=\operatorname{\mathbb{E}}_{s_{i}^{j}\sim F_{i}^{j}}[\min\{s_{i}^{j},t\}]$ be the $t$-truncated expected size of choice $i,j$.
For any $b\in\mathbb{N}$, let $\Psi(b)$ be the following linear program:
$\Psi(b)$ $:=\max_{x}\quad\sum_{i\in[n]}\sum_{j\in[m]}\sum_{t\in\left\{0,\ldots b\right\}}w_{i,j,t}\,x_{i,j,t}$
s.t. $\sum_{i\in[n]}\sum_{j\in[m]}\sum_{t^{\prime}\leq t}x_{i,j,t^{\prime}}\,\mu_{i,j,t}\leq 2t,$ $\forall\,t\in\left\{0,\ldots b\right\},$ (4)
$\sum_{j\in[m]}\sum_{t\in\left\{0,\ldots b\right\}}x_{i,j,t}\leq 1,$ $\forall\,i\in[n],$ (5)
$0\leq x_{i,j,t}\leq 1,$ $\forall\,i\in[n],\,j\in[m],\,t\in\left\{0,\ldots b\right\}.$ (6)
While the runtime of computing the optimal solution is pseudo-polynomial in $b$, it is possible to write a compact (approximate) LP and then use its solution to get a $2$-approximate solution for $\Psi(b)$ in polynomial time. All details of the compact (approximate) LP appear in section E.2.
This $LP$ bounds the expected gain of the optimal adaptive policy:
Lemma 4.4.
$\operatorname{\mathbb{E}}[\mathrm{ADAPT}]\leq\Psi(T)$ where $\mathrm{ADAPT}$ is the utility of the optimal adaptive policy.
We defer the full proof of lemma 4.4 to section E.1.
The idea is to define $x_{i,j,t}$ as the probability that $\sigma_{A}$ (an optimal adaptive policy) attempts to insert item-choice $(i,j)$ at time $t$, and show that $x$ is a feasible solution to the $LP$. This is used to show that the sum of contributions of all item-choice pairs to the expected utility of the optimal adaptive policy is upper bounded by the $LP$.
We now give the algorithm that finds an adaptive policy that achieves a constant approximation of the LP $\Psi(T)$.
Input: Instance II with items $i\in[n]$, choices $j\in[m]$, values $v_{i}$, costs $p_{i}^{j}$, distributions $F_{i}^{j}$
Output: Adaptive policy $\pi$
// Step 1: Solve LP
Compute an optimal solution $x^{*}$ to $\Psi(T)$
// Step 2: Per-item randomized start time (at most one $(j,t)$ per item)
[…]
for $i\in[n]$ do
 $\alpha_{i}\leftarrow\sum_{j\in[m],\,t\in\left\{0,\ldots,T\right\}}x^{*}_{i,j,t}$
  Sample $A_{i}\sim\textnormal{{Bernoulli}}\!\left(\frac{\alpha_{i}}{4}\right)$
 if $A_{i}=0$ then continue;
   Draw $(j_{i},t_{i})\sim\textnormal{{Categorical}}\!\left(\frac{x^{*}_{i,j,t}}{\alpha_{i}}\right)$,   $b_{i,j_{i}}\leftarrow t$ // Implicitly $b_{i,j^{\prime}}\leftarrow\infty$ for $j^{\prime}\neq j_{i}$
[…]
// Step 3: Adaptive execution in the random order
[…]
foreach $(i,j,b_{i,j})\in\operatorname{\mathcal{I}}$ in order do
 if $S\leq b_{i,j}$ then
      Insert $(i,j)$; Observe $s_{i}^{j}$ $S\leftarrow S+s_{i}^{j}$
Algorithm 2 SA-SKC (Skipping-Adaptive Stochastic Knapsack with Costs)
The next lemma tells us that given that the randomly assigned start time of item-choice $i,j$ is $b_{i,j}=t$, the probability that all previous item-choices took less than $t$ time is at least half.
Lemma 4.5.
[…]
Proof.
Let $S=\min\left\{\sum_{(i^{\prime},j^{\prime})\in\operatorname{\mathcal{I}}_{i,j}}s_{i^{\prime}}^{j^{\prime}},t\right\}$ denote how much of $[0,t]$ interval is occupied by items inserted before time $t$. Let $x^{*}$ be the optimal solution to $\Psi(T)$. Note that if $(i^{\prime},j^{\prime})\in\operatorname{\mathcal{I}}_{i,j}$ it must be that $b_{i^{\prime},j^{\prime}}\leq t$. Thus:
$\operatorname{\mathbb{E}}[S\mid b_{i,j}=t]$ $\leq\sum_{i^{\prime},j^{\prime}\in\operatorname{\mathcal{I}}_{i,j}}\operatorname{\mathbb{E}}\left[\min\left\{s_{i^{\prime}}^{j^{\prime}},t\right\}\ \middle|\ b_{i,j}=t\right]$
$=\sum_{i^{\prime},j^{\prime}\in\operatorname{\mathcal{I}}_{i,j}}\operatorname{\mathbb{E}}\left[\min\left\{s_{i^{\prime}}^{j^{\prime}},t\right\}\ \middle|\ b_{i,j}=t,b_{i^{\prime},j^{\prime}}\leq t\right]\Pr\left(b_{i^{\prime},j^{\prime}}\leq t\right)$
$\leq\sum_{i^{\prime},j^{\prime}\in[n]\times[m]}\mu_{i^{\prime},j^{\prime},t}\cdot\left(\frac{\alpha_{i^{\prime}}}{4}\cdot\sum_{t^{\prime}\leq t}\frac{x^{*}_{i^{\prime},j^{\prime},t^{\prime}}}{\alpha_{i^{\prime}}}\right)$
$=\frac{1}{4}\sum_{i^{\prime},j^{\prime}\in[n]\times[m]}\sum_{t^{\prime}\leq t}x^{*}_{i^{\prime},j^{\prime},t^{\prime}}\cdot\mu_{i^{\prime},j^{\prime},t}\leq\frac{t}{2},$
where the first inequality follows from the definition of $S$, the second inequality holds since $\operatorname{\mathcal{I}}_{i,j}\subseteq[n]\times[m]$ and the independence between the size of the item $s_{i^{\prime}}^{j^{\prime}}$ and the assigned starting times of the items $b_{i,j},b_{i^{\prime},j^{\prime}}$ and the starting time assignment probability. The last inequality holds due to constraint eq. 4 of the LP (as $x^{*}$ is a solution to $\Psi(T)$).
By Markov inequality: $\Pr\left(S\geq t\mid b_{i,j}=t\right)\leq\frac{\operatorname{\mathbb{E}}[S\mid b_{i,j}=t]}{t}\leq\frac{1}{2}$, implying the desired.
We can now show the constant approximation via an adaptive policy result:
Theorem 4.6.
Algorithm 2 returns an adaptive policy that achieves a $O(1)$ approximation of the optimal adaptive policy.
Proof.
[…]
$\Pr\left(\mathbb{I}_{\left\{i,j,t\right\}}^{ins}\right)\geq\Pr\left(b_{i,j}=t,s(\operatorname{\mathcal{I}}_{i,j})\leq t\right)=\Pr\left(s(\operatorname{\mathcal{I}}_{i,j})\leq t\mid b_{i,j}=t\right)\cdot\Pr\left(b_{i,j}=t\right)\geq\frac{1}{2}\cdot\frac{x^{*}_{i,j,t}}{4}=\frac{x^{*}_{i,j,t}}{8},$ (7)
where the last inequality is due to lemma 4.5 and the fact that
$\Pr\left(b_{i,j}=t\right)=\Pr\left(A_{i}=1\right)\cdot\Pr\left((j_{i},t_{i})=(j,t)\mid A_{i}=1\right)=\frac{\alpha_{i}}{4}\cdot\frac{x^{*}_{i,j,t}}{\alpha_{i}}=\frac{x^{*}_{i,j,t}}{4}.$
For any $i\in[n],j\in[m]$:
$\operatorname{\mathbb{E}}[W_{i,j}]$ $=\sum_{t\in\left\{0,\ldots,T\right\}}\operatorname{\mathbb{E}}[W_{i,j,t}]=\sum_{t\in\left\{0,\ldots,T\right\}}\operatorname{\mathbb{E}}[W_{i,j,t}\mid\mathbb{I}_{\left\{i,j,t\right\}}^{ins}]\cdot\Pr\left(\mathbb{I}_{\left\{i,j,t\right\}}^{ins}\right)\geq\sum_{t\in\left\{0,\ldots,T\right\}}w_{i,j,t}\cdot\Pr\left(\mathbb{I}_{\left\{i,j,t\right\}}^{ins}\right)$
$\geq\frac{1}{8}\sum_{t\in\left\{0,\ldots,T\right\}}x^{*}_{i,j,t}\cdot w_{i,j,t},$ (8)
where the last inequality holds due to eq. 7.
From linearity of expectation we get:
$\operatorname{\mathbb{E}}[\textbf{SA-SKC}]$ $=\sum_{i,j}\operatorname{\mathbb{E}}[W_{i,j}]\geq\sum_{i,j}\frac{1}{8}\sum_{t\in\left\{0,\ldots,T\right\}}x^{*}_{i,j,t}\cdot w_{i,j,t}=\frac{1}{8}\Psi(T)\geq\frac{1}{8}\mathrm{ADAPT},$
where the first inequality follows from eq. 7 and the second inequality is due to lemma $4.$ ∎
5 Conclusion and Future Directions
In this work, we study an economically motivated extension of the canonical Stochastic Knapsack problem through the lens of simple-vs-optimal policies. By introducing costs, we show that the adaptivity gap can be unbounded, and that this gap is governed by a natural return-on-investment (ROI) parameter. We also introduce a hierarchy of increasingly adaptive policies, and show that limited adaptivity can still achieve near-optimal utility.
An open question concerns the power of stopping-adaptive policies: while we show an $\Omega(\alpha^{1/3})$ gap relative to skipping-adaptive policies, the best known upper bound remains the $O(\alpha)$ guarantee inherited from non-adaptive policies. Closing this gap would complete the picture of our proposed hierarchy.
A broader direction is to revisit other classic decision-making problems where introducing costs is well motivated, and to study how this changes the optimization landscape. We expect simple-vs-optimal approaches to remain useful in such settings, along with the ROI parameter.
Finally, we suggest re-examining other optimization problems with mixed-sign objectives, and exploring whether our approach can yield new insights there as well.
Appendix A Further Related Work
Limited adaptivity.
Different limited adaptivity notions have been studied in the literature. Instead of studying the size of the decision tree, other works study the number of adaptive rounds/choices. Several recent works in stochastic optimization explore the power of $k$ adaptive queries compared to full adaptivity: [17, 1] study Stochastic Submodular Cover, while [26] study Informative Path Planning from an information-theoretic perspective (putting aside computational considerations). Esfandiari et al. [14] consider both Stochastic Submodular Maximization and Stochastic Minimum Cost Coverage. The studied problems and techniques are distinct from ours, and focus on the number of adaptive decision rounds/choices; to our knowledge our notions of limited adaptivity have not been previously studied.
At a high level, our work is related to free-order prophet inequalities and Pandora’s box problems [4, 21], which study adaptive ordering under uncertainty, though the objectives and cost structures differ substantially from ours.
The power of limited adaptivity has also been studied in the context of non-stochastic optimization, in particular, in maximization of submodular and other complement-free set functions: see e.g. the works of [8, 6, 7, 9, 23].
Algorithmic contract theory
The study of sequential contracts in [15] investigates the problem of incentivizing a single agent to perform multiple jobs in a sequential manner. In contrast, our work considers a sequence of contracts in which the jobs are executed sequentially by different agents, with no interdependence between the jobs. The works of [25, 16, 2, 18] consider contract design subject to a limited monetary budget constraining the transfers from principal to agent. In our work, we study a multi-agent setting under a time budget constraint on the agents’ actions, where time is limited and the principal must incentivize the agents to complete their jobs on time.
In recent years, a series of studies has emerged that study contracts through the lens of the theory of computation, initiated by [5, 13]; a recent survey paper [12] provides an overview of developments in algorithmic contract theory. The authors of [13] raise concerns regarding the complexity and unintuitiveness of optimal contracts, while exploring the approximation ratios of simpler contract forms, such as linear contracts. The multi-agent setting was initially explored by Babaioff et al. [5] and further developed by [11], but in their setting the agents only choose between effort and no-effort, and the outcome is only success or failure of the project; we allow multiple actions and outcomes. Several works consider agent decisions under costs and incentives, including sequential exploration settings studied in Hoefer et al. [22], where optimal contracts are designed for a principal–agent exploration task. The work of Hoefer et al. [22] is similar in flavor to ours as it combines stochastic optimization with contract design and delegation.
LP-based approach.
Our approach builds on LP-based upper bounds for adaptive policies introduced in Dean et al. [10], extending them to settings with costs and mixed-sign objectives. We use convexity and dominance arguments to obtain well-structured LP solutions. In the context of limited-adaptivity, we adapt global time-indexed LP techniques from [3, 20, 24] to design constant-approximation skipping-adaptive policies. Unlike [20, 24], our setting also captures multiple choices for each item and negative item utility.
Appendix B Knapsack Contracts: An Application to Contract Design
In this section we show the formal connection between MSKC and between a natural multi-agent contract setting, in which a principal sequentially delegates jobs to agents up to a deadline. Intuitively, this setting is very similar to MSKC, since the principal has a value for each completed job, there are multiple choices of effort levels at which each job can be executed, and incentivizing different choices incurs different costs on the principal (the cost of paying the agent enough to incentivize the chosen effort level; this cost is well-understood based on existing contract theory). There is however one main difference: in the contract setting (unlike MSKC), the principal’s costs are random variables, since the principal pays each agent according to the realized size of each job. In section B.1 we set up the contract problem, and in section B.2 we show how to reduce it to MSKC despite the aforementioned difference.
B.1 Setup
An instance $I$ of the Knapsack Contracts problem consists of a principal, $n$ agents, and a knapsack of size $S$. Without loss of generality, let $S=1$, where $S$ represents a budget of some limited resource. As our running example, we assume $S$ is the time remaining until a set deadline. Each agent $i$ has a job which they can carry out to the benefit of the principal. Agent $i$ is represented by the following:
  • $v_{i}$: The value which the principal obtains if agent $i$ completes their job before the deadline.
  • $\{(F_{i}^{j},c_{i}^{j})\}_{j\in[m]}$: The possible effort levels of agent ii. [^3] All agents have the same number of effort levels without loss of generality, since we can add “dummy” effort levels that would never be incentivized by the principal. Each effort level $j\in[m]$ determines a distribution $F_{i}^{j}$ over the time it takes the agent to complete the job, as well as an agent-cost $c_{i}^{j}\geq 0$. If the agent chooses to exert effort level $j$, they incur agent-cost $c_{i}^{j}$, and complete their job within time $s_{i}$ drawn (independently) from distribution $F_{i}^{j}$. We assume non-decreasing agent-costs $c_{i}^{1}\leq\dots\leq c_{i}^{m}$. Let $\mathbf{F}_{i}=\left(F_{i}^{1},\ldots,F_{i}^{m}\right)$. Let $\mathrm{supp}(F_{i}^{j})$ denote the support of $F_{i}^{j}$, and $\mathrm{supp}(\mathbf{F}_{i})=\cup_{j\in[m]}\mathrm{supp}(F_{i}^{j}).$ We assume the support of each distribution is finite and is explicitly given.
The principal knows $I=(v_{i},\{F_{i}^{j},c_{i}^{j}\}_{j\in[m]})_{i\in[n]}$ in advance. After an agent completes a job, the principal observes the time it took the agent to complete it, but it does not observe the effort level chosen by the agent. The principal offers each agent $i$ a contract $t_{i}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, mapping $i$’s observed processing time to a transfer paid by the principal to the agent. Given contract $t_{i}$, agent $i$ chooses an effort level
$j\in\arg\max_{\ell}\operatorname{\mathbb{E}}_{s_{i}\sim F_{i}^{\ell}}[t_{i}(s_{i})]-c_{i}^{\ell}.$
Definition B.1 (Implementable effort level).
For agent $i$, an effort level $j$ is implementable if there exists a contract $t_{i}$ for which this action maximizes the agent’s expected utility among all actions given the contract. In this case, we say contract $t_{i}$ incentivizes action $j$.
Any implementable effort level $j$ for agent $i$ has an optimal contract that incentivizes it, which can be obtained by solving the following LP:
$\min\quad$ $\sum_{s\in\mathrm{supp}(F_{i}^{j})}t_{i}(s)\cdot F_{i}^{j}(s)$
s.t. $\sum_{s\in\mathrm{supp}(F_{i}^{j})}t_{i}(s)\cdot F_{i}^{j}(s)-c_{i}^{j}\geq\sum_{s\in\mathrm{supp}(F_{i}^{j^{\prime}})}t_{i}(s)\cdot F_{i}^{j^{\prime}}(s)-c_{i}^{j^{\prime}}$ $\forall j^{\prime}\in[m]$
$t_{i}(s)\geq t_{i}(s^{\prime})$ $\forall s<s^{\prime}\in\mathrm{supp}(\mathbf{F}_{i})$ (9)
$t_{i}(s)\geq 0$ $\forall s\in\mathrm{supp}(\mathbf{F}_{i})$
Constraint (9) ensures monotonicity of payments. This is required as without it, an agent might have an incentive to hover around after completing the job and get a higher payment.
The principal’s problem can then be described as: when hiring agent $i$, the principal chooses an effort level $j$ to incentivize, and offers the optimal contract that incentivizes that effort level. The principal’s collected value is the sum of the values of the jobs that the agents get to complete with a total processing time less than the budget $S$. The principal aims to maximize their utility: the expected total value minus the expected sum of transfers made to the agents.
B.2 Reduction
The following proposition shows that in order to find an approximately optimal policy for the Knapsack Contracts problem, it suffices to find an approximately optimal policy for the MSKC problem.
See 1.1
Proof.
Given a Knapsack Contracts instance, for each agent $i\in[n]$, we create a corresponding item $i$ in the MSKC instance. For each implementable effort level $j$ of agent $i$, we create a corresponding choice $j$ for item $i$. The parameters are mapped as follows:
  • The value of item $i$ is set to the principal’s value $v_{i}$.
  • The size of choice $(i,j)$ is a random variable with the distribution $F_{i}^{j}$.
  • The cost of choice $(i,j)$ is set to $p_{i}^{j}=\operatorname{\mathbb{E}}_{s\sim F_{i}^{j}}[t_{i}(s)]$, where $t_{i}$ is the contract that minimizes the expected payment, as given by the LP in Definition B.1.
Now, consider any policy $\varphi$ for the contracts problem, and denote by $\varphi^{\prime}$ the corresponding policy for the knapsack problem. $\varphi$ specifies the choice of (agent, effort level) pairs for the contracts problem, and $\varphi^{\prime}$ specifies the corresponding item-choice pairs for the knapsack problem.
It is left to show that the expected profits are identical. Denote by $S$ the set of successfully completed jobs in the contracts and knapsack problems, and by $A$ the set of agents/items attempted in the contracts and knapsack problems. Notice that $S$ and $A$ have exactly the same distribution under $\varphi$ and $\varphi^{\prime}$, as both policies make the same choices under the same randomization.
The utility for policy $\varphi$ is a random variable that depends on the (possibly random) choices of the policy and the random sizes of the attempted jobs. Its expectation is given by
$\operatorname{\mathbb{E}}[\mathrm{utility}(\varphi)]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in S}{v_{i}}-\sum_{(i,j)\in A}{t_{i}(s_{i})}\right]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in S}{v_{i}}\right]-\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in A}{t_{i}(s_{i})}\right].$ (10)
Expanding the transfer term, we get:
$\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in A}{t_{i}(s_{i})}\right]$ $=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in A}{t_{i}(s_{i})}\right]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)}{\mathbbm{1}_{(i,j)\in A}\cdot t_{i}(s_{i})}\right]=\sum_{(i,j)}\operatorname{\mathbb{E}}\left[{\mathbbm{1}_{(i,j)\in A}\cdot t_{i}(s_{i})}\right]$
$\stackrel{{\scriptstyle(\star)}}{{=}}\sum_{(i,j)}\operatorname{\mathbb{E}}\left[\mathbbm{1}_{(i,j)\in A}\right]\cdot\operatorname{\mathbb{E}}_{s_{i}\sim F_{i}^{j}}\left[t_{i}(s_{i})\right]\stackrel{{\scriptstyle(\star\star)}}{{=}}\sum_{(i,j)}\operatorname{\mathbb{E}}\left[\mathbbm{1}_{(i,j)\in A}\right]\cdot p_{i}^{j}$
$=\operatorname{\mathbb{E}}\left[\sum_{(i,j)}\mathbbm{1}_{(i,j)\in A}\cdot p_{i}^{j}\right]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in A}p_{i}^{j}\right],$
where equality $(\star)$ is due to the independence between $\mathbbm{1}_{(i,j)\in A}$ and $s_{i}^{j}$, and equality $(\star\star)$ is due to the definition of $p_{i}^{j}$. Plugging this back in eq. 10 yields:
$\operatorname{\mathbb{E}}[\mathrm{utility}(\varphi)]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in S}{v_{i}}-\sum_{(i,j)\in A}p_{i}^{j}\right]=\operatorname{\mathbb{E}}[\mathrm{utility}(\varphi^{\prime})]~,$
which implies the desired.
Thus, given a $\varphi^{\prime}$ that $c$-approximate the optimal policy in the constructed instance of the MSKC problem, the corresponding $\varphi$ $c$-approximates the optimal policy in the original Knapsack Contracts instance. ∎
Implications of the reduction to the Knapsack Contracts Problem.
The above implies: $(i)$ there exists a skipping-adaptive policy that gives an $O(1)$ approximation to the optimal principal’s utility (by Theorem 4.6); and $(ii)$ one can naturally adapt the ROI parameter to the Knapsack Contracts setting, treating expected transfer that the principal makes when incentivizing agent $i$ to exert effort level $j$ as $p_{i}^{j}$ in Definition 2.1. This implies that an instance with $\alpha$ IOR for the Knapsack Contracts problem is reduced to an MSKC instance with $\alpha$ IOR. Thus, Theorem 3.2 implies an $O(\alpha)$-approximate non-adaptive policy for the Knapsack Contracts problem. As all of our lower bound instances use items with a single choice, the lower bounds also trivially extend to the Knapsack Contracts problem, where each agent has a single effort level, and the principal need only cover the agent-cost exactly. Therefore, all results obtained for the MSKC problem, summarized by Figure 1, also apply to the contract problem.
Appendix C $\Omega(\alpha)$ Informational Gap
In this section we establish a lower bound that exposes a fundamental information gap in the problem of Stochastic Knapsack with Costs. Previous constant-factor approximation algorithms for Stochastic Knapsack (without costs) have used limited distributional information, specifically the expected (or truncated) sizes and the probability of fitting within the current capacity. In Proposition C.1 we demonstrate that this information is insufficient. We show that any algorithm, even one that finds a fully-adaptive policy, which relies solely on such limited data, cannot achieve a constant-factor approximation in the presence of costs. More generally, we prove in Theorem C.2 that even providing the first $k$ moments of each distribution for any constant $k$ is insufficient.
C.1 Warm-Up: First Moment
Proposition C.1.
Let $\mathcal{A}$ be any algorithm whose knowledge of the item size distributions is restricted to the expected values $\operatorname{\mathbb{E}}[s_{i}]$, the truncated expected values $\mu_{i}=\operatorname{\mathbb{E}}[\min\{s_{i},1\}]$, and oracle access to the function $q_{i}(B)=\Pr[s_{i}\leq B]$ for each $i\in[n]$, where $B$ is the current remaining knapsack capacity. For every $\alpha$ there exists an input instance $I$ on which the algorithm’s performance is bounded by:
$\frac{\mathrm{ADAPT}(I)}{\operatorname{\mathbb{E}}[\mathcal{A}(I)]}\geq\Omega(\alpha).$
The proof relies on constructing two types of items, one “good” and one “bad”. The optimal strategy would be to insert only the “good” items, and any algorithm that starts by inserting a “bad” item will result in a solution with a value that is only a $\frac{1}{\alpha}$-fraction of OPT. We design the size distributions to be indistinguishable to algorithm $\mathcal{A}$, forcing $\mathcal{A}$ to arbitrarily pick an item to insert into the knapsack. By creating an instance where the number of “bad” items far exceeds the number of “good” items, algorithm $\mathcal{A}$ will almost surely pick a “bad” item first, resulting in our lower bound.
Proof of Proposition C.1.
Fix $\epsilon\in(0,1)$. The two item types are defined as follows:
Property Type 1 Item (“good”) Type 2 Item (“bad”)
Value $v_{i}$ 22 22
Cost $p_{i}$ $2-\epsilon$ $2-\epsilon$
Size $s_{i}$ dist. $s_{1}=\begin{cases}0,&\text{w.p. }1-\epsilon,\\ 1,&\text{w.p. }\epsilon;\end{cases}$ $s_{2}=\begin{cases}\epsilon/2,&\text{w.p. }1-\frac{\epsilon}{2-\epsilon},\\ 1,&\text{w.p. }\frac{\epsilon}{2-\epsilon}.\end{cases}$
Note that in this instance the expected values and the truncated expected values are identical: $\operatorname{\mathbb{E}}[s_{1}]=\operatorname{\mathbb{E}}[s_{2}]=\mu_{1}=\mu_{2}=\epsilon$ and $q_{1}(1)=q_{2}(1)=1$, hence, algorithm $\mathcal{A}$ cannot distinguish between the two items. Inserting one item of type 2 results in a net utility of $\epsilon$, and reduces the remaining capacity $B$ by at least $\epsilon/2$. Inserting any further items (of either types) results in a negative expected utility:
$\operatorname{\mathbb{E}}[\text{inserting item of type $i$}\;|\;B\leq 1-\epsilon/2]$ $\leq\Pr[s_{i}\neq 1]\cdot v_{i}-p_{i}$
$\leq(1-\frac{\epsilon}{2-\epsilon})\cdot 2-(2-\epsilon)$
$=\epsilon(1-\frac{2}{2-\epsilon})\;<\;0.$
Now, consider an input instance $I$ with $n>\frac{1}{\epsilon}$ items of type 1 and $n^{2}-n$ items of type 2. The optimal policy will attempt to insert items of type 1, and stop if an item’s size instantiates $s_{i}=1$, or if no more items of type 1 remain. The expected number of items successfully inserted by OPT is $\frac{1}{\epsilon}$, each item contributing $\epsilon$ net utility, resulting in
$\mathrm{ADAPT}(I)\geq\frac{1}{\epsilon}\cdot\epsilon=1.$
Since $\mathcal{A}$’s only strategy is to randomly select an item to insert, with probability $1-\frac{1}{n}$ it begins by inserting a “bad” item. By taking $n\to\infty$, the expected utility of $\mathcal{A}$ approaches $\epsilon$:
$\lim_{n\to\infty}\operatorname{\mathbb{E}}[\mathcal{A}(I)]=\epsilon.$
Thus for a large enough $n$, the performance ratio of the algorithm can be bounded:
$\frac{\mathrm{ADAPT}(I)}{\operatorname{\mathbb{E}}[\mathcal{A}(I)]}\geq\frac{1}{2\epsilon}\geq\frac{1}{4}\alpha.$
C.2 $k$ Moments
We generalize the previously shown distributional informational gap to show that, even an algorithm with knowledge of the first $k$ (or truncated) moments of the distribution, has an $\Omega(\alpha)$ gap compared to an algorithm with knowledge of the exact distribution.
Theorem C.2.
Fix $k\geq 1$. Let $\mathcal{A}$ be any algorithm whose knowledge of the item size distributions is restricted to the moment $\operatorname{\mathbb{E}}[s_{i}^{j}]$, and truncated moment $\mu_{i}^{j}=\operatorname{\mathbb{E}}[(\min\{s_{i},1\})^{j}]$ for every item $i\in[n]$ and moment $j\in[k]$, and oracle access to the function $q_{i}(B)=\Pr[s_{i}\leq B]$, where $B$ is the current remaining knapsack capacity. For every $\alpha$ there exists an input instance $I$ on which the algorithm’s performance is bounded by:
$\frac{\mathrm{ADAPT}(I)}{\operatorname{\mathbb{E}}[\mathcal{A}(I)]}\geq\Omega(\alpha).$
Similarly to the previous proof, we construct two types of items, one “good” and one “bad”, which are indistinguishable to $\mathcal{A}$ yet are critically different for the optimal algorithm.
Proof of Theorem C.2.
Fix $\epsilon\in(0,1)$. Fix distinct sizes
$t_{j}:=1-2^{-j},\quad j=1,2,\dots,k,$
$t_{k+1}:=1.$
The two item types are defined as follows. For sufficiently small $\delta>0$ and for $p_{j}\in[0,1]$ to be determined later:
Property Type 1 Item (“good”) Type 2 Item (“bad”)
Value $v_{i}$ 22 22
Cost $c_{i}$ $2-\epsilon$ $2-\epsilon$
Size $s_{i}$ dist. $s_{1}=\begin{cases}0,&\text{w.p. }1-2\epsilon\\ t_{j},&\text{w.p. }\tfrac{\epsilon}{k+1},\;j=1,\dots,k+1\\ 1,&\text{w.p. }\epsilon\end{cases}$ $s_{2}=\begin{cases}\delta,&\text{w.p. }1-2\epsilon\\ t_{j},&\text{w.p. }p_{j},\;j=1,\dots,k+1\\ 1,&\text{w.p. }\epsilon\end{cases}$
Note that since $\Pr[s_{i}\leq 1]=1$, the truncated moments are identical to the original moments. We now focus on finding the correct parameters such that the two distributions have the same moments; We seek $p_{j}$ with $\sum_{j}p_{j}=\epsilon$, satisfying for each $r=1,\dots,k$:
$\operatorname{\mathbb{E}}[s_{1}^{r}]=\operatorname{\mathbb{E}}[s_{2}^{r}].$
That is, for $r=1,\dots,k$:
$\sum_{j=1}^{k+1}\frac{\epsilon}{k+1}\,t_{j}^{r}+\epsilon$ $=(1-2\epsilon)\,\delta^{r}+\sum_{j=1}^{k+1}p_{j}\,t_{j}^{r}+\epsilon$
$\iff\sum_{j=1}^{k+1}t_{j}^{r}\,\left(\frac{\epsilon}{k+1}-p_{j}\right)$ $=(1-2\epsilon)\,\delta^{r}.$
Denote $x_{j}:=\frac{\epsilon}{k+1}-p_{j}$. Then $\sum_{j}x_{j}=0$ (since $\sum p_{j}=\epsilon$ and $\sum\frac{\epsilon}{k+1}=\epsilon$). We get the following system of equations:
$\sum_{j=1}^{k+1}x_{j}$ $=0;$ (11)
$\sum_{j=1}^{k+1}t_{j}^{r}\,x_{j}$ $=(1-2\epsilon)\,\delta^{r},\quad\text{for }r=1,\dots,k.$ (12)
We now rewrite (11) and (12) in matrix form $Ax=b$, where
$A=\begin{pmatrix}1&1&\cdots&1\\ t_{1}&t_{2}&\cdots&t_{k+1}\\ t_{1}^{2}&t_{2}^{2}&\cdots&t_{k+1}^{2}\\ \vdots&\vdots&&\vdots\\ t_{1}^{k}&t_{2}^{k}&\cdots&t_{k+1}^{k}\end{pmatrix},x=\begin{pmatrix}x_{1}\\ \vdots\\ x_{k+1}\end{pmatrix},b=\begin{pmatrix}0\\ (1-2\epsilon)\delta\\ \vdots\\ (1-2\epsilon)\delta^{k}\end{pmatrix}$.
Note that $A$ is a $(k+1)\times(k+1)$ matrix where the first row corresponds to the zero-sum constraint (11), and the subsequent $k$ rows correspond to the moment-matching conditions (12). This is a transposed Vandermonde matrix [19]. Since the $t_{j}$ are distinct, it yields a unique solution $x=A^{-1}b$. Hence
$p_{j}=\frac{\epsilon}{k+1}-x_{j}.$
It is left to prove that $p_{j}\geq 0$ to conclude that $s_{2}$ is indeed a valid distribution. Note that since $A$ and $A^{-1}$ are functions of $k$ (and the fixed $t_{j}$ values), then after fixing $k$, $x_{j}$ is a function of $\epsilon$ and $\delta$:
$x_{j}(\epsilon,\delta)=\sum_{r=1}^{k}(A^{-1})_{j,r+1}[(1-2\epsilon)\delta^{r}].$
For a fixed $\epsilon$, we have that $|x_{j}(\delta)|=O(\delta)$. Thus, for $p_{j}=\frac{\epsilon}{k+1}-x_{j}$ to be non-negative, we can choose $\delta$ sufficiently small such that $\frac{\epsilon}{k+1}$ dominates the $x_{j}$ terms, ensuring $p_{j}\geq 0$. Hence, the two distributions are indistinguishable to $\mathcal{A}$.
Inserting a single item of type 2 results in a net utility of $\epsilon$, and reduces the remaining capacity $B$ by at least $\delta$. Inserting any further items (of either types) results in a negative expected utility:
$\operatorname{\mathbb{E}}[\text{inserting item of type $i$}\;|\;B\leq 1-\delta]$ $\leq\Pr[s_{i}\neq 1]\cdot v_{i}-p_{i}$
$=(1-\epsilon)\cdot 2-(2-\epsilon)$
$=-\epsilon<0.$
We conclude the proof by considering an input instance $I$ with $n>\frac{1}{\epsilon}$ items of type 1 and $n^{2}-n$ items of type 2. The optimal policy does as well as the following policy: it will attempt to insert items of type 1, and stop if an item’s size instantiates $s_{i}\neq 0$, or if no more items of type 1 remain. The expected number of items successfully inserted by OPT is $\frac{1}{2\epsilon}$, each item contributing $\epsilon$ net utility, resulting in
$\mathrm{ADAPT}(I)\geq\frac{1}{2\epsilon}\cdot\epsilon=\frac{1}{2}.$
Since $\mathcal{A}$’s only strategy is to randomly select an item to insert, with probability $1-\frac{1}{n}$ it begins by inserting a “bad” item. By taking $n\to\infty$, the expected utility of $\mathcal{A}$ approaches $\epsilon$:
$\lim_{n\to\infty}\operatorname{\mathbb{E}}[\mathcal{A}(I)]=\epsilon.$
Thus for a large enough $n$, the performance ratio of the algorithm can be bounded:
$\frac{\mathrm{ADAPT}(I)}{\operatorname{\mathbb{E}}[\mathcal{A}(I)]}\geq\frac{\frac{1}{2}}{2\epsilon}\geq\frac{1}{8}\alpha.$
Appendix D Omitted Proofs from Section 3
D.1 Proof of theorem 3.1
Proof.
For a constant $\alpha$, the claim is trivially true. We prove the claim for every $\alpha>3$. We build an instance $I$ as follows. Consider some $\epsilon<\frac{1}{2}$ and any $\gamma\in(0,\tfrac{1}{2})$, where $\epsilon$ will be used to determine $\alpha$. Our instance has $n$ items for $n=\lceil\frac{1}{\epsilon}\rceil$. Each item has value $v=2$ and cost $p=2-\epsilon$.
The size distribution of item $i\in[n]$ is
$s_{i}\;=\;\begin{cases}a_{i},&\text{with probability }1-\epsilon,\\[4.0pt] b_{i},&\text{with probability }\epsilon,\end{cases}$ where $\qquad a_{i}=\gamma^{\,n-i+1},\qquad b_{i}=1-\displaystyle\sum_{j=1}^{i-1}a_{j}$
By Definition 2.1, we have $\alpha=\frac{2-\epsilon}{\epsilon}\;=\;\Theta\!\left(\frac{1}{\epsilon}\right).$ Note that for every $\alpha>3$, we can choose an $\epsilon<1/2$ in order to set the IOR of our instance to be $\alpha$.
Since for every $i$, $\sum_{j=1}^{i-1}a_{j}+b_{i}=1,$ a stopping-adaptive policy can safely insert the items sequentially until it encounters the first item $i$ whose realized size equals $b_{i}$, and then halt.
We show that since $b_{i}+a_{i^{\prime}}>1$ for every pair $i<i^{\prime}$, attempting to insert item $i^{\prime}$ without considering the realization of earlier items’ size yields negative expected utility. Consequently, the optimal $\mathrm{NON\_ADAPT}$ policy executes at most a single item. Our proof combines these two facts to yield the desired bound.
Formally, consider a stopping-adaptive policy that inserts items in increasing index order ($i=1,2,\ldots$) until it encounters the first item whose size realizes as $s_{i}=b_{i}$, and then stops. Let $X$ denote the number of items the policy inserts. We have
$\operatorname{\mathbb{E}}[X]\;=\;\sum_{i=1}^{n}\Pr[X\geq i]\;=\;\sum_{i=1}^{n}(1-\epsilon)^{i-1}\;=\;\frac{1-(1-\epsilon)^{n}}{\epsilon}.$
Since $n\geq 1/\epsilon$, we get that $\operatorname{\mathbb{E}}[X]\geq\frac{1-e^{-1}}{\epsilon}$. Therefore the expected utility is at least
$\mathrm{STOP\_ADAPT}(I)\;=\;\operatorname{\mathbb{E}}[X]\,(v-p)\;\geq\;\frac{1-e^{-1}}{\epsilon}\,\cdot\,\epsilon\;=\;1-e^{-1}.$
Consider now a non‑adaptive policy. This policy fixes the number of items to insert, $k$, and an ordering $\pi=i_{1},i_{2},\ldots,i_{k}$. For each $\ell\in[k]$, let $\mathrm{utility}(\ell)$ denote the utility from the $\ell$‑th item in the sequence. Clearly, $\operatorname{\mathbb{E}}[\mathrm{utility}(1)]=\epsilon$ for every non‑empty sequence of items. We will show that for every sequence with $k>1$ we have $\operatorname{\mathbb{E}}[\mathrm{utility}(k)]<0$, which implies that it is always better to execute at most a single item, since we can iteratively remove items and only improve the expected utility of the policy.
Given a sequence $\pi=i_{1},i_{2},\ldots,i_{k}$ of items, consider first the case in which $i_{\ell}>i_{k}$ for some earlier item in the sequence ($\ell<k$). If $i_{k}$ is reached and its size realises as $s_{i_{k}}=b_{i_{k}}$, the item overflows, because $s_{i_{\ell}}+s_{i_{k}}\geq a_{i_{\ell}}+b_{i_{k}}>1$. Hence
[…]
$\leq\Pr[i_{k}\text{ is reached}]\Bigl(\Pr[s_{i_{k}}=b_{i_{k}}](-p)+\Pr[s_{i_{k}}=a_{i_{k}}](v-p)\Bigr)$
$=\Pr[i_{k}\text{ is reached}]\bigl(\epsilon(-p)+(1-\epsilon)(v-p)\bigr)$
$=\Pr[i_{k}\text{ is reached}]\bigl(-2\epsilon+\epsilon^{2}+\epsilon-\epsilon^{2}\bigr)$
$=-\,\epsilon\,\Pr[i_{k}\text{ is reached}]\;<\;0.$
Now consider the case where $i_{\ell}<i_{k}$ for all $\ell<k$. Let $j=\arg\max_{\ell<k}i_{\ell}$, the largest‑index item among the first k1k-1 inserted. Note that, $i_{j}<i_{k}$ by our assumption. We have that
$\operatorname{\mathbb{E}}[\mathrm{utility}(k)]$ $=$ $\operatorname{\mathbb{E}}[\mathrm{utility}(k)|i_{k}\mbox{ is reached}]\cdot\Pr[i_{k}\mbox{ is reached}]$
$=$ $\operatorname{\mathbb{E}}[\mathrm{utility}(k)|s_{i_{\ell}}=a_{i_{\ell}}\quad\forall\ell<k]\cdot\Pr[s_{i_{\ell}}=a_{i_{\ell}}\quad\forall\ell<k]$
$+\operatorname{\mathbb{E}}[\mathrm{utility}(k)|i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]$
$\ \ \cdot\Pr[i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]$
$=$ $(v-p)\cdot(1-\epsilon)^{k-1}+(-p)\cdot\Pr[i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]$
$=$ $\epsilon\cdot(1-\epsilon)^{k-1}+(-p)\cdot\Pr[i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]$
$\leq$ $\epsilon\cdot(1-\epsilon)^{k-1}+(-p)\cdot\Pr[s_{i_{j}}=b_{i_{j}}\ \wedge\ (s_{i_{\ell}}=a_{i_{\ell}}\quad\forall\ell<k,\ell\neq j)]$
$=$ $\epsilon\cdot(1-\epsilon)^{k-1}+(-2+\epsilon)\cdot(1-\epsilon)^{k-2}\cdot\epsilon$
$=$ $\epsilon\cdot(1-\epsilon)^{k-2}(1-\epsilon-2+\epsilon)$
$=$ $-\epsilon\cdot(1-\epsilon)^{k-2}<0.$
In the above, the third equality follows since $i_{k}$ always overflows in case some $s_{i_{\ell}}=b_{i_{\ell}}$ for some $i_{\ell}<i_{k}$, as $b_{i_{\ell}}+a_{i_{k}}>1$. The first inequality follows from the fact that p-p is negative, and $\Pr[i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]\geq\Pr[s_{i_{j}}=b_{i_{j}}\ \wedge\ (s_{i_{\ell}}=a_{i_{\ell}}\quad\forall\ell<k,\ell\neq j)]$ since for the event $s_{i_{j}}=b_{i_{j}}$ and for all other $\ell<k$, $s_{i_{\ell}}=a_{i_{\ell}},$ we have that $i_{k}$ is reached. Since $\sum_{\ell<k,\ell\neq j}a_{i_{\ell}}\leq\sum_{\ell=1}^{i_{j}-1}a_{\ell}=1-b_{i_{j}},$ there will be no overflowing item before $i_{k}$ is executed.
We conclude that the optimal non-adaptive policy only inserts a single item, and obtains a utility of $\epsilon$, while the optimal stopping-adaptive policy obtains a utility of 1, as argued above. Therefore, $\frac{\mathrm{STOP\_ADAPT}(I)}{\mathrm{NON\_ADAPT}(I)}\geq\frac{1-e^{-1}}{\epsilon}=\Omega(\alpha).$
D.2 Proof of lemma 3.4
Full proof of lemma 3.4.
For two solutions $x,x^{*}$, denote by $\Delta_{w}$ and $\Delta_{\mu}$ the difference in the total values and the total expected truncated sizes, respectively, between $x^{*}$ and $x$. We want to show that our constructed solution $x^{*}$ has at least as much value ($\Delta_{w}\geq 0$) and at most as much expected truncated size ($\Delta_{\mu}\leq 0$) as the original solution $x$, to conclude that $x^{*}$ is a feasible optimal solution.
For the single dominance case, we define
$x^{*}_{ij^{\prime}}=x_{ij^{\prime}}+x_{ij},\quad x^{*}_{ij}=0~.$
We have that
$\Delta_{w}$ $=(x_{ij^{\prime}}^{*}-x_{ij^{\prime}})w_{ij^{\prime}}+(x_{ij}^{*}-x_{ij})w_{ij}$
$=x_{ij}w_{ij^{\prime}}-x_{ij}w_{ij}$
$=x_{ij}(w_{ij^{\prime}}-w_{ij})$
$\geq 0$ $\text{(since $w_{ij^{\prime}}\geq w_{ij}$)}~,$
$\Delta_{\mu}$ $=(x_{ij^{\prime}}^{*}-x_{ij^{\prime}})\mu_{ij^{\prime}}+(x_{ij}^{*}-x_{ij})\mu_{ij}$
$=x_{ij}\mu_{ij^{\prime}}-x_{ij}\mu_{ij}$
$=x_{ij}(\mu_{ij^{\prime}}-\mu_{ij})$
$\leq 0$ $\text{(since $\mu_{ij^{\prime}}\leq\mu_{ij}$)}~.$
For the double dominance case, we define
$x^{*}_{ij^{\prime}}=x_{ij^{\prime}}\;+\;x_{ij}\cdot\frac{w_{ij^{\prime\prime}}-w_{ij}}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}},\quad x^{*}_{ij^{\prime\prime}}=x_{ij^{\prime\prime}}\;+\;x_{ij}\cdot\frac{w_{ij}-w_{ij^{\prime}}}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}},\quad x^{*}_{ij}=0~.$
We have that
$\Delta_{w}$ $=(x_{ij^{\prime}}^{*}-x_{ij^{\prime}})w_{ij^{\prime}}+(x_{ij^{\prime\prime}}^{*}-x_{ij^{\prime\prime}})w_{ij^{\prime\prime}}+(x_{ij}^{*}-x_{ij})w_{ij}$
$=x_{ij}w_{ij^{\prime}}\cdot\frac{w_{ij^{\prime\prime}}-w_{ij}}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}}+x_{ij}w_{ij^{\prime\prime}}\cdot\frac{w_{ij}-w_{ij^{\prime}}}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}}-x_{ij}w_{ij}$
$=x_{ij}\left(\frac{w_{ij^{\prime}}(w_{ij^{\prime\prime}}-w_{ij})+w_{ij^{\prime\prime}}(w_{ij}-w_{ij^{\prime}})-w_{ij}(w_{ij^{\prime\prime}}-w_{ij^{\prime}})}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}}\right)$
$=0~,$
$\Delta_{\mu}$ $=(x_{ij^{\prime}}^{*}-x_{ij^{\prime}})\mu_{ij^{\prime}}+(x_{ij^{\prime\prime}}^{*}-x_{ij^{\prime\prime}})\mu_{ij^{\prime\prime}}+(x_{ij}^{*}-x_{ij})\mu_{ij}$
$=x_{ij}\mu_{ij^{\prime}}\cdot\frac{w_{ij^{\prime\prime}}-w_{ij}}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}}+x_{ij}\mu_{ij^{\prime\prime}}\cdot\frac{w_{ij}-w_{ij^{\prime}}}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}}-x_{ij}\mu_{ij}$
$=x_{ij}\left(\frac{\mu_{ij^{\prime}}(w_{ij^{\prime\prime}}-w_{ij})+\mu_{ij^{\prime\prime}}(w_{ij}-w_{ij^{\prime}})-\mu_{ij}(w_{ij^{\prime\prime}}-w_{ij^{\prime}})}{w_{ij^{\prime\prime}}-w_{ij^{\prime}}}\right)~.$
From double dominance we have that
$\frac{\mu_{ij}-\mu_{ij^{\prime}}}{w_{ij}-w_{ij^{\prime}}}\geq\frac{\mu_{ij^{\prime\prime}}-\mu_{ij}}{w_{ij^{\prime\prime}}-w_{ij}}$
Since $w_{ij}-w_{ij^{\prime}}>0$ and $w_{ij^{\prime\prime}}-w_{ij}>0$, we can cross-multiply:
$(\mu_{ij}-\mu_{ij^{\prime}})(w_{ij^{\prime\prime}}-w_{ij})\geq(\mu_{ij^{\prime\prime}}-\mu_{ij})(w_{ij}-w_{ij^{\prime}})$
Expanding both sides gives:
$\mu_{ij}w_{ij^{\prime\prime}}-\mu_{ij}w_{ij}-\mu_{ij^{\prime}}w_{ij^{\prime\prime}}+\mu_{ij^{\prime}}w_{ij}\geq\mu_{ij^{\prime\prime}}w_{ij}-\mu_{ij^{\prime\prime}}w_{ij^{\prime}}-\mu_{ij}w_{ij}+\mu_{ij}w_{ij^{\prime}}$
We can cancel the $\mu_{ij}w_{ij}$ term from both sides. Rearranging the terms to match the numerator of our expression for $\Delta_{\mu}$ yields:
$\mu_{ij^{\prime}}(w_{ij^{\prime\prime}}-w_{ij})+\mu_{ij^{\prime\prime}}(w_{ij}-w_{ij^{\prime}})-\mu_{ij}(w_{ij^{\prime\prime}}-w_{ij^{\prime}})\leq 0~.$
This confirms the numerator is non-positive. Since $w_{ij^{\prime\prime}}>w_{ij^{\prime}}$ we conclude that $\Delta_{\mu}\leq 0$. ∎
Proof of lemma 3.6.
$\Pr[size(\mathcal{S})\geq t]$ $=\Pr[\min\{size(\mathcal{S}),t\}\geq t]$
$\leq\frac{1}{t}\operatorname{\mathbb{E}}[\min\{size(S),t\}]$
$=\frac{1}{t}\operatorname{\mathbb{E}}[{\min\{\sum_{(i,j)\in\mathcal{S}}{s_{i}^{j}},t\}}]$
$\leq\frac{1}{t}\operatorname{\mathbb{E}}[\sum_{(i,j)\in\mathcal{S}}{\min\{s_{i}^{j},t\}}]$
$=\frac{1}{t}\sum_{(i,j)\in\mathcal{S}}{\operatorname{\mathbb{E}}[\min\{s_{i}^{j},t\}]}$
$=\frac{1}{t}\sum_{(i,j)\in\mathcal{S}}{\mu_{ij}}=\frac{1}{t}\mu(S)$
Proof of lemma 3.7.
Let $\mathcal{S^{-}}$ be the (random) set of item choices that the policy inserts except the final item choice. By the stopping condition
$size(\mathcal{S}^{-})<t~.$
It follows that for every $(i,j)\in\mathcal{S}^{-}$, $s_{i}^{j}<t$ and therefore $\mu_{ij}=\operatorname{\mathbb{E}}[\min\{s_{i}^{j},t\}]=\operatorname{\mathbb{E}}[s_{i}^{j}]$. Denote by $(i^{*},j^{*})$ the final item choice made by the policy. Then,
$\operatorname{\mathbb{E}}[\mu(\mathcal{S})]$ $=\operatorname{\mathbb{E}}[\sum_{(i,j)\in\mathcal{S}}{\mu_{ij}}]=\operatorname{\mathbb{E}}[\sum_{(i,j)\in\mathcal{S}^{-}}{\mu_{ij}}]+\operatorname{\mathbb{E}}[\mu_{i^{*}j^{*}}]$
$=\operatorname{\mathbb{E}}[\sum_{(i,j)\in\mathcal{S}^{-}}{s_{i}^{j}}]+\operatorname{\mathbb{E}}[\min\{t,s_{i^{*}}^{j^{*}}\}]=\operatorname{\mathbb{E}}[size(\mathcal{S}^{-})]+\operatorname{\mathbb{E}}[\min\{t,s_{i^{*}}^{j^{*}}\}]$
$\leq t+t=2t~.$
Proof of lemma 3.8.
For any adaptive policy and capacity $t$, denote by $\mathcal{S}$ the random set of pairs $(i,j)$ that the policy inserts. Set $x_{ij}=\Pr[(i,j)\in\mathcal{S}]$. From the previous lemma, the expected mass that $\mathrm{ADAPT}$ attempts to insert is $\operatorname{\mathbb{E}}(\mu(\mathcal{S}))=\sum_{ij}\mu_{ij}x_{ij}\leq 2t$, therefore, it is a feasible solution to $\Phi(2t)$ with $\sum_{ij}w_{ij}x_{ij}\leq\Phi(2t)$. Since $w_{ij}x_{ij}$ is an upper bound on the expected utility of inserting the item choice $(i,j)$, we have that
$\mathrm{ADAPT}\leq\sum_{ij}{w_{ij}x_{ij}}\leq\Phi(2t)\leq\Phi(ct)~,$
for any $c\geq 2$.
Now, given an optimal solution $x$ to $\Phi(ct)$, we can construct a feasible solution $x^{\prime}$ to $\Phi(t)$ by setting $x^{\prime}_{ij}=\frac{x_{ij}}{c}$ for all $i,j$, and then increasing the value of $x^{\prime}_{ij_{\bot}}$ to ensure that $\sum_{j}x^{\prime}_{i,j}=1$ for all $i$. This results in a feasible solution to $\Phi(t)$ with $\sum_{ij}w_{ij}x^{\prime}_{ij}=\frac{1}{c}\sum_{ij}w_{ij}x_{ij}$. We conclude that
$\mathrm{ADAPT}\leq\Phi(ct)\leq c\cdot\Phi(t)$
D.3 LP Overconfidence: An $\Omega(\alpha)$ Overestimation of the True Utility
In this section, we provide an instance of the MSKC problem where the value of the LP introduced in Eq. (1) is larger by a factor $\Omega(\alpha)$ from the obtainable value of any policy. This is in stark contrast to the case without costs, where Dean et al. [10] show a non-adaptive policy whose utility is a constant factor away from the value of the analogous LP. As our non-adaptive policy in Section 3.2 gives $O(\alpha)$-approximation to the value of the LP in Eq. (1), this approximation is the best possible with respect to the value of the LP.
In our instance, we have an infinite number of identical items with value $1$ and cost $1-\epsilon$. The size-distribution of the items is
$s=\begin{cases}\epsilon^{n},&\text{w.p. }1-\epsilon,\\ 1,&\text{w.p. }\epsilon;\end{cases}.$
By Definition 2.1, the value of the IOR parameter is $\alpha=\frac{1-\epsilon}{\epsilon}=O(1/\epsilon)$. For every item $i$, we have $\mu_{i}=(1-\epsilon)\cdot\epsilon^{n}+\epsilon\cdot 1=\epsilon^{n}+\epsilon-\epsilon^{n+1}\leq 2\epsilon$ and $w_{i}=\epsilon$. Setting $x_{i}=1$ for $\ell=\lfloor\frac{1}{2\epsilon}\rfloor$ items, and for an additional item $i^{\prime}$ set $x_{i^{\prime}}=\frac{1}{2\epsilon}$-$\ell$, implies that
$\sum_{i}x_{i}\cdot\mu_{i}\leq 2\epsilon\cdot\sum_{i}x_{i}=2\epsilon\cdot\frac{1}{2\epsilon}=1,$
which is a feasible assignment to the LP. Therefore, $\Phi(1)\geq\frac{1}{2\epsilon}\cdot\epsilon=1/2.$ We next bound the utility of any policy for this instance.
For any policy, after inserting the first item, we have that any item fits with probability at most $1-\epsilon$; thus, we have that the benefit from trying to insert any item after the first one is at most $(1-\epsilon)v-p=0.$ Therefore, the policy that only inserts one item is optimal, and it obtains a utility of $\epsilon\leq 2\epsilon\cdot\Phi(1)=\Phi(1)/\Omega(\alpha).$
Appendix E Appendix for Section 4
E.1 Proof of lemma 4.4
Proof.
Let $\sigma_{A}$ be an optimal adaptive policy. Let $x_{i,j,t}$ be the probability that $\sigma_{A}$ attempts to insert $(i,j)$ at time $t$. We claim that $x$ is a feasible solution to the LP $\Psi(T)$. It is immediate that eq. 6, eq. 5 hold (by the definition of $x_{i,j,t}$, each item and choice are inserted at most once in the run of $\sigma_{A}$). It remains to show that eq. 4 holds. For any $i\in[n]$, $j\in[m]$, $t^{\prime}\in\left\{0,\ldots,T\right\}$: let $\mathbb{I}_{\left\{i,j,t^{\prime}\right\}}^{ins}$ be the indicator for the event where item-choice $(i,j)$ is inserted at time $t^{\prime}$ and let $\mathbb{I}_{\left\{i,j,s\right\}}^{size}$ be the indicator for the event where item-choice $(i,j)$ realizes to size $s$; that is: $s_{i}^{j}=s$.
For any time $t$, let $(i_{t},j_{t})$ be the last item-choice pair that $\sigma_{A}$ inserts up until or at time $t$, and let $P_{t}$ be the set of all item-choices successfully inserted by $\sigma_{A}$ before inserting $(i_{t},j_{t})$. So the sum of all inserted item sizes is at most $t$.
$\sum_{i,j\in P_{t}}\sum_{t^{\prime}\leq t}\sum_{s\leq T}\mathbb{I}_{\left\{i,j,s\right\}}^{size}\cdot s\leq t.$
Let $P^{\prime}_{t}=P_{t}\cup\left\{(i_{t},j_{t})\right\}$. We get $\sum_{i,j\in P^{\prime}_{t}}s_{i}^{j}=\sum_{i,j\in P_{t}}s_{i}^{j}+s_{i_{t}}^{j_{t}}$ and thus:
$\sum_{i,j\in P^{\prime}_{t}}\sum_{t^{\prime}\leq t}\sum_{s\leq T}\mathbb{I}_{\left\{i,j,s\right\}}^{size}\cdot\min(s,t)\leq 2t.$
For any $t^{\prime}\leq t$ and $i\in[n],j\in[m]$: $\mathbb{I}_{\left\{i,j,t^{\prime}\right\}}^{ins}=1\iff{i,j}\in P^{\prime}_{t}$ and thus we have:
$\sum_{i,j\in[n]\times[m]}\ \sum_{t^{\prime}\leq t}\ \sum_{s\leq T}\mathbb{I}_{\left\{i,j,t^{\prime}\right\}}^{ins}\cdot\mathbb{I}_{\left\{i,j,s\right\}}^{size}\cdot\min(s,t)\leq 2t.$
By taking expectation and using linearity of expectation we have:
$\sum_{i,j\in[n]\times[m]}\ \sum_{t^{\prime}\leq t}\ \sum_{s\leq T}\operatorname{\mathbb{E}}\left[\mathbb{I}_{\left\{i,j,t^{\prime}\right\}}^{ins}\cdot\mathbb{I}_{\left\{i,j,s\right\}}^{size}\right]\cdot\min(s,t)=$
$\qquad=\sum_{i,j\in[n]\times[m]}\ \sum_{t^{\prime}\leq t}\Pr\left(\mathbb{I}_{\left\{i,j,t^{\prime}\right\}}^{ins}\right)\left(\sum_{s\leq T}\Pr\left(\mathbb{I}_{\left\{i,j,s\right\}}^{size}\right)\min(s,t)\right)$
$\qquad=\sum_{i,j\in[n]\times[m]}\ \sum_{t^{\prime}\leq t}x_{i,j,t^{\prime}}\ \mu_{i,j,t},$
where the first equality holds due to the independence between $\mathbb{I}_{\left\{i,j,t^{\prime}\right\}}^{ins}$, $\mathbb{I}_{\left\{i,j,s\right\}}^{size}$ (the events are independent as the policy inserts item-choice $(i,j)$ before seeing the realized value of $s_{i}^{j}$). Hence
$\sum_{i,j\in[n]\times[m]}\ \sum_{t^{\prime}\leq t}x_{i,j,t^{\prime}}\ \mu_{i,j,t}\leq 2t.$
We’ve shown that $x$ is indeed a feasible solution to the LP.
We now bound $E[\mathrm{ADAPT}]$. Let $W_{i,j}$ be the utility $\sigma_{A}$ gets from item-choice $i,j$ (if $\sigma_{A}$ did not attempt inserting the item-choice then the corresponding utility is 0). From linearity of expectation we get:
$\operatorname{\mathbb{E}}[\mathrm{ADAPT}]$ $=\sum_{i,j}\operatorname{\mathbb{E}}[W_{i,j}]=\sum_{i,j}\sum_{t\in\left\{0,\ldots,T\right\}}\operatorname{\mathbb{E}}[W_{i,j}\mid\sigma_{A}\text{ inserts }(i,j)\text{ at time }t^{\prime}]\cdot\Pr\left(\mathbb{I}_{\left\{i,j,t\right\}}^{ins}\right)$
$=\sum_{i,j}\sum_{t\in\left\{0,\ldots,T\right\}}\operatorname{\mathbb{E}}[W_{i,j}\mid\sigma_{A}\text{ inserts }(i,j)\text{ at time }t]\cdot x_{i,j,t}$
$\leq\sum_{i,j}\sum_{t\in\left\{0,\ldots,T\right\}}x_{i,j,t}\cdot w_{i,j,t}\leq\sum_{i,j}\sum_{t\in\left\{0,\ldots,T\right\}}x^{*}_{i,j,t}\cdot w_{i,j,t}=\Psi(T),$
where the last inequality holds due to the fact that $x$ is a feasible solution to the LP. ∎
E.2 From Pseudo-Polytime to Truly-Polytime Algorithm
In this section, we consider the $O(1)$ skipping-adaptive algorithm presented in section 4.2 (see algorithm 2), showing how to turn the algorithm from pseudo-polynomial to fully-polynomial. We do so by giving a compact LP that may be solved in polynomial time yet gives a $2$ approximation to the original LP.
We formulate an LP $\Psi_{P}(b)$ - a compact version of $\Psi(b)$ (defined in section 4.2) with $O(nm\log(b))$ constraints.
We first show that $\Psi_{P}(b)\geq\frac{1}{2}\Psi(b)$ (lemma E.1), and then give a way to create an approximate solution to $\Psi(b)$ given a solution to $\Psi_{P}(b)$ (lemma E.2). Thus, it is enough to use $\Psi_{P}(b)$ in our algorithm. We finally discuss how to modify the algorithm to use the compact LP.
For any $b\in\mathbb{N}$, let $\Psi_{P}(b)$ be the following linear program:
$\Psi_{P}(b)$ $:=\max_{x}\quad\sum_{i\in[n]}\sum_{j\in[m]}\sum_{t=0}^{\log(b)}w_{i,j,2^{t+1}}\,x_{i,j,2^{t}}$
s.t. $\sum_{i\in[n]}\sum_{j\in[m]}\sum_{t^{\prime}\leq t}x_{i,j,2^{t^{\prime}}}\,\mu_{i,j,2^{t+1}}\leq 2\cdot 2^{t},$ $\forall\,t\in[0,\log(b)],$ (13)
$\sum_{j\in[m]}\sum_{t\in[0,\log(b)]}x_{i,j,2^{t}}\leq 1,$ $\forall\,i\in[n],$ (14)
$0\leq x_{i,j,2^{t}}\leq 1,$ $\forall\,i\in[n],\,j\in[m],\,t\in[0,\log(b)].$ (15)
Lemma E.1.
$\Psi_{P}(b)\geq\frac{1}{2}\Psi(b)$.
Proof.
Let $x$ be the optimal solution to $\Psi(b)$.
Let $\hat{x}_{i,j,1}=\frac{1}{2}\sum_{t\in\left\{1,2,3\right\}}x_{i,j,t}$, and $\hat{x}_{i,j,2^{t}}=\frac{1}{2}\sum_{t\in[2^{t+1},2^{t+2})}x_{i,j,t}$ for any $t\in(1,\log(b)]$.
Let us show that all constraints of $\Psi_{P}(b)$ hold for $\hat{x}$.
For any $i\in[n]$, by constraint 5 of $\Psi(b)$:
$\sum_{j\in[m]}\sum_{t\in[0,\log(b)]}\hat{x}_{i,j,2^{t}}\leq\frac{1}{2}\sum_{j\in[m]}\sum_{t\geq 0}x_{i,j,t}\leq\frac{1}{2}.$
Also,
$\sum_{i\in[n]}\sum_{j\in[m]}\sum_{t^{\prime}\leq t}\hat{x}_{i,j,2^{t^{\prime}}}\,\mu_{i,j,2^{t+1}}$ $=\sum_{i\in[n]}\sum_{j\in[m]}\sum_{t^{\prime}=1}^{2^{t+2}-1}\frac{1}{2}x_{i,j,t}\cdot\mu_{i,j,2^{t+1}}$
$\leq\frac{1}{2}\sum_{i\in[n]}\sum_{j\in[m]}\sum_{t^{\prime}=1}^{2^{t+2}-1}x_{i,j,t}\cdot\mu_{i,j,2^{t+2}-1}\leq\frac{1}{2}\left(2^{t+2}-1\right)\leq 2\cdot 2^{t},$
where the first inequality is due to the fact that $\mu_{i,j,2^{t+1}}\leq\mu_{i,j,2^{t+2}-1}$ by the definition of $\mu$, and the second inequality is due to the fact that $x$ is a feasible solution to $\Psi(b)$ and thus constraint 4 holds.
So all constraints of $\Psi_{P}(b)$ hold and thus $\hat{x}$ is a feasible solution. Thus, the value obtained from this solution is a lower bound for $\Psi_{P}(b)$ and therefore:
$\Psi_{P}(b)\geq\sum_{i\in[n]}\sum_{j\in[m]}\sum_{t=0}^{\log(b)}w_{i,j,2^{t+1}}\,\hat{x}_{i,j,2^{t}}\geq\frac{1}{2}\sum_{i\in[n]}\sum_{j\in[m]}\sum_{t\in\left\{0,\ldots,b\right\}}w_{i,j,t}\,x_{i,j,t}=\frac{1}{2}\Psi(b).$
Let $x^{*}$ be the optimal solution of $\Psi_{P}(b)$. Let $y$ be defined as follows: $y_{i,j,t}=\frac{x^{*}_{i,j,2^{l}}}{2^{l}}$ for any $i,j\in[n]\times[m]$ and $l=\lfloor{\log(t)}\rfloor$. In the following lemma we show that $y$ is a feasible solution of $\Psi(b)$ which yields the same value as $x^{*}$.
Lemma E.2.
Let $x^{*}$ be the optimal solution of $\Psi_{P}(b)$, and let $y\in[n]\times[m]\times\left\{0,\ldots,b\right\}$ be defined by $y_{i,j,t}=\frac{x^{*}_{i,j,2^{l}}}{2^{l}}$ for any $i,j\in[n]\times[m]$ and $l=\lfloor{\log(t)}\rfloor$. Then $y$ is a feasible solution of $\Psi(b)$ yielding a value of at least half the optimal LP value.
Proof.
Since $x^{*}$ is a feasible solution of $\Psi_{P}(b)$ then constraints 5, 6 hold (due to constraints 14, 15 of $\Psi_{P}(b)$).
Let $t\in[2^{l},2^{l+1})$ for some $l\in[\log(b)]$.
$\sum_{i\in[n]}\sum_{j\in[m]}\sum_{t^{\prime}\leq t}y_{i,j,t^{\prime}}\,\mu_{i,j,t}$ $=\sum_{i\in[n]}\sum_{j\in[m]}\sum_{t^{\prime}\leq t}\frac{x^{*}_{i,j,2^{\lfloor\log(t^{\prime})\rfloor}}}{2^{\lfloor\log(t^{\prime})\rfloor}}\,\mu_{i,j,t}\leq\sum_{i\in[n]}\sum_{j\in[m]}\sum_{t^{\prime}\leq t}\frac{x^{*}_{i,j,2^{\lfloor\log(t^{\prime})\rfloor}}}{2^{\lfloor\log(t^{\prime})\rfloor}}\,\mu_{i,j,2^{l+1}}$
$\leq\sum_{i\in[n]}\sum_{j\in[m]}\sum_{l^{\prime}\leq l}\sum_{t\in[2^{l^{\prime}},2^{l^{\prime}+1}-1]}\frac{x^{*}_{i,j,2^{l^{\prime}}}}{2^{l^{\prime}}}\,\mu_{i,j,2^{l+1}}$
$=\sum_{i\in[n]}\sum_{j\in[m]}\sum_{l^{\prime}\leq l}x^{*}_{i,j,2^{l^{\prime}}}\,\mu_{i,j,2^{l+1}}\leq 2\cdot 2^{l}\leq 2\cdot t,$
where the first inequality is due to the monotonicity of $\mu_{i,j,t}$ in $t$, the second inequality is due to summing over more elements and the last inequality is due to constraint 13 of $\Psi_{P}(b)$.
We deduce that constraint 4 holds, and thus $y$ is a feasible solution to $\Psi(b)$.
Finally:
$\sum_{i\in[n]}\sum_{j\in[m]}\sum_{l\in[0,\log(b)]}\sum_{t\in[2^{l},2^{l+1}-1]}w_{i,j,t}\,\frac{x^{*}_{i,j,2^{l}}}{2^{l}}$ $\geq\sum_{i\in[n]}\sum_{j\in[m]}\sum_{l\in[0,\log(b)]}\sum_{t\in[2^{l},2^{l+1}-1]}w_{i,j,2^{l+1}}\,\frac{x^{*}_{i,j,2^{l}}}{2^{l}}$
$=\sum_{i\in[n]}\sum_{j\in[m]}\sum_{l\in[0,\log(b)]}w_{i,j,2^{l+1}}\,x^{*}_{i,j,2^{l}}=\Psi_{P}(b),$
where the first inequality holds due to the monotonicity of $w_{i,j,t}$ in $t$. We thus get the desired by lemma E.1. ∎
The complete fully polynomial algorithm can simply solve the $\Psi_{P}(T)$ LP, and use its solution $x^{*}$ such that for each $x^{*}_{i,j,2^{l}}$ we add an additional step of choosing $t$ in a uniform way: $t\sim Unif{[2^{l},2^{l+1}-1]}$. This way, the probability to choose item choice $(i,j)$ with $b_{i,j_{i}}=t$ is $\frac{x^{*}_{i,j,2^{l}}}{2^{l}}$ which is indeed $y$ (and so the proof works as before).
Since this time $y$ is not the optimal solution, but a $2$-approximate solution, we pay an additional constant factor and still get a $O(1)$ approximation ratio.
E.3 Handling Continuous Distributions
In algorithm 2 we assume that (similar to [20]) the size distributions are discrete; namely, that all the items’ sizes are integers between $1$ and $T$ (or, equivalently, that the sizes are $\left\{0,\frac{1}{T},\frac{2}{T},\cdots,1\right\}$). In this section, we describe how to apply our algorithm to instances with continuous distributions, using discretization and $\epsilon$-resource augmentation.
Fix $\epsilon>0$. We set the discretization parameter $\delta=\epsilon/n$. Given an instance $I$ with continuous size distributions, we construct a rounded instance $\hat{I}$ where all item sizes $s_{i}^{j}$ are rounded down to the nearest multiple of $\delta$: $\hat{s}_{i}^{j}=\delta\lfloor s_{i}^{j}/\delta\rfloor$.
Let $OPT(I)$ and OPT(I^)OPT($\hat${I}) be the optimal expected utilities for the original and rounded instances, respectively, with capacity 1. Since reducing item sizes can only increase the probability of successfully fitting any set of items, the optimal value for the rounded problem is an upper bound on the optimal value for the original problem. Thus, $OPT(\hat{I})\geq OPT(I)$.
algorithm 2 returns a policy $\varphi$ that achieves a $O(1)$ approximate policy for $\hat{I}$.
We claim that given an augmented capacity of $1+\epsilon$, the expected utility of $\varphi$ on $I$ is the same as the expected utility of $\varphi$ running on $\hat{I}$.
For any execution path (realization) of $\varphi$, let $\mathcal{S}$ be the set of item-choice pairs successfully inserted by $\varphi$. By the knapsack constraint: $\sum_{(i,j)\in\mathcal{S}}\hat{s}_{i}^{j}\leq 1$.
In the original instance, the true size of each item is $s_{i}^{j}=\hat{s}_{i}^{j}+\eta_{i}^{j}$, where $0\leq\eta_{i}^{j}<\delta$. Since at most $n$ items can be placed in the knapsack, the total rounding error is bounded by $n\delta=n(\epsilon/n)=\epsilon$. Thus, the total size in the original instance is:
$\sum_{(i,j)\in\mathcal{S}}s_{i}^{j}<\sum_{(i,j)\in\mathcal{S}}\hat{s}_{i}^{j}+\epsilon\leq 1+\epsilon.$
We deduce that the same policy $\varphi$ applied to instance $I$ given an $\epsilon$ additional capacity may get at least the same utility as $\varphi$ applied to instance $\hat{I}$ with no additional capacity.
Appendix F Non-Stopping Policies
In this section, we consider the non-stopping model, where the policy must first commit to a bundle of items alongside their order of insertion, and then insert all the items according to the order, without stopping at the first overflow; therefore, such a policy incurs the costs of all items in the bundle regardless of how many of these actually fit in the knapsack. This is clearly a restriction of the non-adaptive model, where the policy stops after the first overflow.
In the original work of stochastic knapsack without costs by Dean et al. [10], non-adaptive and non-stopping policies are equivalent, since inserting additional items after the first overflow does not increase the expected value and incurs no additional cost. However, in our setting with costs, an overflow may result in payments for many items without obtaining their corresponding values. This model describes real-world scenarios where payments for jobs are made in advance, before any jobs have been executed and their durations are realized. We formally define:
Definition F.1 (Non-Stopping Policy).
A non-stopping policy is an ordering of a subset of items and choices $\pi$. The policy inserts all items in the given order, paying all their associated costs, and collecting the values of all items that fit in the knapsack before overflow. Let $\Pi$ be the set of non-stopping policies. We denote by $\mathrm{NON\_STOP}(I)=\max_{\pi\in\Pi}\operatorname{\mathbb{E}}[\mathrm{utility}(\pi)]$ the optimum expected utility obtained by a non-stopping policy for instance $I$.
In Section F.1 we show that in our model there is a $\Omega(\alpha)$-gap between the optimal non-stopping policy and the optimal non-adaptive policy. In Section F.2, we present an $O(\alpha)$-approximate non-stopping policy to the optimal fully-adaptive policy.
F.1 Non-Stopping vs. Non-Adaptive Policies
Unlike the no-costs case, in our setting, non-stopping policies can be much worse than non-adaptive policies.
Theorem F.1.
For every $\alpha$, there exists an instance $I$ of the MSKC problem such that
$\frac{\mathrm{NON\_ADAPT}(I)}{\mathrm{NON\_STOP}(I)}=\Omega(\alpha).$
Proof.
Consider $n$ items with value $v_{i}=1$, cost $p_{i}=1-\epsilon$, and size distribution
$s_{i}=\begin{cases}0,&\text{w.p. }1-\epsilon^{2},\\ 2,&\text{w.p. }\epsilon^{2}.\end{cases}$
Note that $\alpha=\frac{1}{\epsilon}$. Denote by $\pi_{NA}$ a non-adaptive policy that inserts all items until the first overflow (i.e., until an item realizes to $s_{i}=2$). The expected number of items inserted before the first overflow is $\frac{1}{\epsilon^{2}}$, each contributing a utility of $\epsilon$, except the last (overflowing) item, incurring a cost of $1-\epsilon$. This results in a total expected utility of
$\mathrm{utility}(\pi_{NA})\geq 1/\epsilon-1.$
For a non-stopping policy, since all items are identical, the only degree of freedom for the algorithm is choosing how many items it would attempt to insert. Consider the non-stopping policy $\pi_{NS-k}$ that inserts $k$ items, for some $1\leq k\leq n$. We prove that the expected utility of this policy is at most 1. The fixed total cost of inserting $k$ items is $k(1-\epsilon)$. The expected value from the $i$th item is $(1-\epsilon^{2})^{i}$, as it is successfully inserted only if all the first $i$ items fit. Hence, the expected value overall is
$\sum_{i=1}^{k}(1-\epsilon^{2})^{i}=(1-\epsilon^{2})\frac{1-(1-\epsilon^{2})^{k}}{1-(1-\epsilon^{2})}=\frac{1-\epsilon^{2}}{\epsilon^{2}}[1-(1-\epsilon^{2})^{k}].$
Therefore, the expected utility of $\pi_{NS-k}$ is
$\mathrm{utility}(\pi_{NS-k})=\frac{1-\epsilon^{2}}{\epsilon^{2}}[1-(1-\epsilon^{2})^{k}]-k(1-\epsilon).$
To maximize this expression over $k$, consider the marginal utility of inserting an additional item:
$\Delta p(k)=\mathrm{utility}(\pi_{NS-k+1})-\mathrm{utility}(\pi_{NS-k})=(1-\epsilon^{2})^{k+1}-(1-\epsilon).$
$\Delta p(k)$ is decreasing in $k$, thus, the optimal $k$ is the largest $k$ such that $\Delta p(k)>0$. Solving for $k$, we get
$\Delta p(k)>0$
$\iff$ $(1-\epsilon^{2})^{k+1}>1-\epsilon$
$\iff$ $(k+1)\ln(1-\epsilon^{2})>\ln(1-\epsilon)$
$\iff$ $k<\frac{\ln(1-\epsilon)}{\ln(1-\epsilon^{2})}-1$
We rewrite the RHS:
$\frac{\ln(1-\epsilon)}{\ln(1-\epsilon^{2})}-1$ $=\frac{\ln(1-\epsilon)-\ln(1-\epsilon^{2})}{\ln(1-\epsilon^{2})}$
$=\frac{\ln(1-\epsilon)-\left[\ln(1-\epsilon)+\ln(1+\epsilon)\right]}{\ln(1-\epsilon^{2})}$
$=\frac{-\ln(1+\epsilon)}{\ln(1-\epsilon^{2})}$
$\leq\frac{-\epsilon}{-\epsilon^{2}}=\frac{1}{\epsilon}$
Where the final inequality uses the fact that $\ln(1+x)\leq x$ for x>1x>-1, for both the numerator and denominator. Therefore, the optimal $k$ is at most $\frac{1}{\epsilon}$. Even if all $k$ items fit, the algorithm’s utility is at most
$\mathrm{utility}(\pi_{NS-k})\leq k\cdot\epsilon\leq\frac{1}{\epsilon}\cdot\epsilon=1.$
We conclude that for this instance,
$\frac{\mathrm{NON\_ADAPT}(I)}{\mathrm{NON\_STOP}(I)}\geq\frac{\mathrm{utility}(\pi_{NA})}{\max_{k}[\mathrm{utility}(\pi_{NS-k})]}\geq\frac{1/\epsilon-1}{1}=\Omega(\alpha).$
F.2 An $O(\alpha)$ Non-Stopping Policy
In this section, we present an $O(\alpha)$-approximate non-stopping algorithm for the MSKC problem.
Theorem F.2.
There exists a polynomial time algorithm for MSKC that computes for every input $I$ a non-stopping policy $\pi_{I}$ such that $\frac{\mathrm{ADAPT}(I)}{\operatorname{\mathbb{E}}[\mathrm{utility}(\pi_{I})]}\leq O(\alpha).$
The algorithm is a variation of Algorithm 1: We start by solving the LP $\Phi(1)$ and consider items sorted by their density. However, in contrast to Algorithm 1, we then insert only a small ($\approx 1/\alpha$) mass of items, to ensure that with high probability no overflowing occurs.
An overview of the algorithm.
First, we solve the LP $\Phi(1)$ to obtain an optimal solution $x$ of item-choices. As shown in Section 3.2, we can calculate a solution such that every item has one choice with $x_{ij}=1$ except for at most one item with two choices $x_{ij}+x_{ij^{\prime}}=1$, which can be converted into a single choice. Also as in Section 3.2, since each item now has a single (possibly virtual) choice, we refer to each item-choice pair $(i,j)$ by the item $i$ for convenience. We then sort the items by decreasing order of density $d_{i}=\frac{w_{i}}{\mu_{i}}$, and define a threshold $U=\frac{1}{3(1+\alpha)}$. Denote by $S$ the maximal prefix of items $\{1,\dots,k\}$ such that their cumulative truncated size satisfies
$\mu(S)=\sum_{i=1}^{k}\mu_{i}\leq U<\sum_{i=1}^{k+1}\mu_{i}.$
The non-stopping algorithm NS-SKC either inserts the items in $S$, or the single most profitable item-choice $(i_{\max},j_{\max})$ with $w_{\max}=\max_{i}\{w_{i}\}$. Specifically, it inserts $S$ if $\frac{2}{3}\sum_{i=1}^{k}w_{i}\geq w_{\max}$, and inserts $(i_{\max},j_{\max})$ otherwise. Denote by $\pi_{NS}(S)$ and by $\pi_{NS}(w_{\max})$ the policies that insert $S$ and $(i_{\max},j_{\max})$, respectively. The following lemma lower bounds the expected utility obtained when inserting $S$.
Lemma F.3.
$\operatorname{\mathbb{E}}[\mathrm{utility}(\pi_{NS}(S))]\geq\frac{2}{3}\sum_{i=1}^{k}w_{i}$
Proof.
Denote by $E_{i}$ the event that all items $1,\dots,i$ fit in the knapsack. The expected utility of the policy is
$\operatorname{\mathbb{E}}[\mathrm{utility}(\pi_{NS}(S))]$ $=\sum_{i=1}^{k}v_{i}\cdot\Pr[E_{i}]-p_{i}$
$\geq\sum_{i=1}^{k}v_{i}\cdot(1-\mu(S))-p_{i}$ (by lemma 3.6)
$\geq\sum_{i=1}^{k}(w_{i}+p_{i})\cdot(1-\mu(S))-p_{i}$
$=(1-\mu(S))\cdot\sum_{i=1}^{k}w_{i}-\mu(S)\cdot\sum_{i=1}^{k}p_{i}$
$\geq(1-\mu(S))\cdot\sum_{i=1}^{k}w_{i}-\mu(S)\cdot\alpha\sum_{i=1}^{k}w_{i}$ (since $\forall i:p_{i}\leq\alpha w_{i}$)
$=\sum_{i=1}^{k}w_{i}\cdot(1-(1+\alpha)\mu(S))$
$\geq\sum_{i=1}^{k}w_{i}\cdot\left(1-\frac{1+\alpha}{3(1+\alpha)}\right)$ (since $\mu(S)\leq U$)
$=\frac{2}{3}\sum_{i=1}^{k}w_{i}.$
We can now conclude with the main result of this section.
Proof of Theorem F.2.
Since $\sum_{i=1}^{k+1}\mu_{i}>U$ and the items are ordered by density, we have that
$\sum_{i=1}^{k}w_{i}+w_{\max}\geq\sum_{i=1}^{k+1}w_{i}>U\cdot\Phi(1)=\frac{1}{3(1+\alpha)}\Phi(1).$
It follows that
$\max\left\{\sum_{i=1}^{k}w_{i},w_{\max}\right\}\geq\frac{1}{2}\cdot\left(\sum_{i=1}^{k}w_{i}+w_{\max}\right)\geq\frac{1}{6(1+\alpha)}\Phi(1).$
Our non-stopping policy inserts $S$ if $\frac{2}{3}\sum_{i=1}^{k}w_{i}\geq w_{\max}$, and inserts $(i_{\max},j_{\max})$ otherwise. Therefore, by Lemma F.3 the expected utility of NS-SKC is bounded by
$\operatorname{\mathbb{E}}[\textbf{NS-SKC}]$ $\geq\max\left\{\operatorname{\mathbb{E}}[\mathrm{utility}(\pi_{NS}(S))],\operatorname{\mathbb{E}}[\mathrm{utility}(\pi_{NS}(w_{\max}))]\right\}$
$\geq\max\left\{\frac{2}{3}\sum_{i=1}^{k}w_{i},w_{\max}\right\}$
[…]
We remark that [10] studies a policy class termed fixed-set, in which the policy must choose an unordered subset of items to insert, and receives the cumulative values of the items if and only if the entire set is successfully inserted into the knapsack; otherwise, it obtains zero utility. In the classic stochastic knapsack problem without costs, it is shown that there exists a fixed-set policy that achieves a constant approximation of the fully-adaptive optimal policy.
Note that a fixed-set policy can only perform worse than a non-stopping policy, since an overflow results in no received value at all (and possibly a large negative utility in our model due to the presence of costs). This is in contrast to a non-stopping policy that, at the very least, receives the value of the items inserted prior to the overflow. However, our non-stopping policy can be easily converted into a fixed-set policy: either insert $S$ in its entirety, or insert $(i_{\max},j_{\max})$. It can be similarly shown that such a policy achieves an $O(\alpha)$ approximation of the optimal fully-adaptive policy.
Appendix G Relaxations
In this section, we consider several relaxations of the original model. In particular, we focus on the setting where items are size-bounded with high probability; specifically, for some $\delta\in(0,1)$, for all $i,j$,
$\Pr_{s_{i}^{j}\sim F_{i}^{j}}\!\left[s_{i}^{j}>1-\delta\right]\leq\frac{1}{2n\alpha}.$
To facilitate the analysis, we first study a relaxed variant in which there is no explicit bound on item sizes. In this variant, the policy receives the value of each item inserted before the capacity is met, and the process terminates upon overflow.
We show that this policy approximates the utility of the optimal policy in the original model, where no value is collected from an overflowing item. We then demonstrate how to transform this policy into one that operates under the bounded-size assumption and does not collect value from overflowing items.
G.1 Start-by-Deadline Model
We define SKC-OF to be the algorithm that accepts all items inserted before reaching the capacity. SKC-OF runs NA-SKC (Algorithm 1) with a different threshold $\mathcal{T}=\Phi(\delta)/4$. Note that, unlike the NA-SKC algorithm, this threshold is independent of $\alpha$. We denote by $\textbf{SKC-OF}(t)$, the algorithm that runs NA-SKC with a capacity $t$, and a threshold $\mathcal{T}=\Phi(t)/4$. In particular, $\textbf{SKC-OF}=\textbf{SKC-OF}(\delta)$.
Lemma G.1.
For any constant $c\geq 1$,
$\operatorname{\mathbb{E}}[\textbf{SKC-OF}(t)]\geq\frac{1}{c}\Phi(ct)/4.$
Proof.
In case $w_{\max}\geq\Phi(t)/4$, we only take $w_{\max}$ and obtain a utility of
$v_{i_{\max}}-p_{i_{\max}}^{j_{\max}}\geq w_{\max}\geq\Phi(t)/4\geq\frac{1}{c}\Phi(ct)/4,$
where the last inequality follows Lemma 3.8. Hence, assume that $w_{\max}<\Phi(t)/4$.
Recall that $\textbf{SKC-OF}_{k}$ is the utility of item $k$ in the SKC-OF algorithm, as defined in Section 3.2. In our case, the utility from item $k$ will be as large as in the MSKC instance, and will not incur a cost in case of an overflow. Therefore, by adapting Eq. (2), our bound on $\textbf{SKC-OF}_{k}$ will be
$\operatorname{\mathbb{E}}[\textbf{SKC-OF}_{k}]\geq w_{k}\left(1-\frac{1}{t}\sum_{j=1}^{k}{\mu_{j}}\right).$
We adapt the bound on SKC-OF from Eq. (3) accordingly to be
$\operatorname{\mathbb{E}}[\textbf{SKC-OF}]\geq\Phi(t)/2-w_{\max}\geq\Phi(t)/4\geq\frac{1}{c}\Phi(ct)/8,$ (16)
where the second inequality follows our assumption that $w_{\max}\leq\Phi(t)/4$ and the last inequality follows Lemma 3.8. ∎
If the policy gets the value of any item inserted when there is still some capacity left (even if it overflows), then the adaptivity gap becomes constant, as formalized below.
Lemma G.2.
For every instance of the MSKC problem $I$, $\mathrm{ADAPT}_{OF}(1)\leq 2\Phi(1)+w_{max}$.
Proof.
let $\mathcal{S}$ be the random set of item choices selected by the policy, and recall that $w_{ij}=v_{i}\cdot\Pr_{{s_{i}^{j}\sim F_{i}^{j}}}[s_{i}^{j}\leq 1]-p_{i}^{j}$ (like in section 3.2, and w.l.o.g we assume that: $\Pr_{{s_{i}^{j}\sim F_{i}^{j}}}[s_{i}^{j}\leq 1]=1$), $w_{\max}=\max_{(i,j)\in\mathcal{C}}w_{ij}$ (Algorithm 1). Using $\Phi(t)$ (Eq. (1)), we obtain:
$\mathrm{ADAPT}_{OF}(1)$ $=$ $\sum_{ij}w_{ij}\cdot\Pr[(i,j)\in\mathcal{S}]$ (17)
$=$ $\sum_{ij}w_{ij}\cdot\Pr[(i,j)\in\mathcal{S}\wedge s_{i}\leq 1]+\sum_{ij}w_{ij}\cdot\Pr[(i,j)\in\mathcal{S}\wedge s_{i}>1]$
$\leq$ $\sum_{ij}w_{ij}\cdot\Pr[s_{i}\leq 1]\cdot\Pr[(i,j)\in\mathcal{S}]+w_{max}\sum_{ij}\Pr[(i,j)\in\mathcal{S}\wedge s_{i}>1]$
$\leq$ $\sum_{ij\in\mathcal{S}}w_{ij}\cdot x_{ij}+w_{max}$
$\leq$ $\Phi(2)+w_{max}$
$\leq$ $2\Phi(1)+w_{max}.$
The first inequality holds because $\Pr[s_{i}>1]$ is independent of $(i,j)\in\mathcal{S}$. The final inequality follows since at most one element per run has size greater than 1, implying the summation of those probabilities is at most one. By Lemma 3.7, we know $\operatorname{\mathbb{E}}[\mu(\mathcal{S})]\leq 2$, giving $\sum_{ij\in A}w_{ij}x_{ij}\leq\Phi(2)$. ∎
Theorem G.3.
SKC-OF is a 9-approximation algorithm to $\mathrm{ADAPT}_{OF}$.
Proof.
By Equations (16),(17):
$\frac{\mathrm{ADAPT}_{OF}}{\textbf{SKC-OF}}$ $\leq$ $\frac{2\Phi(1)+w_{max}}{\Phi(1)/4}.$
If $w_{max}>\Phi(1)/4$, then $\textbf{SKC-OF}=w_{max}$ giving:
$\frac{\mathrm{ADAPT}_{OF}}{\textbf{SKC-OF}}$ $\leq$ $\frac{8w_{max}+w_{max}}{w_{max}}=9.$
Else:
$\frac{\mathrm{ADAPT}_{OF}}{\textbf{SKC-OF}}$ $\leq$ $\frac{2\Phi(1)+\Phi(1)/4}{\Phi(1)/4}=9.$
G.2 Bounded-Size Items
We now turn to the bounded item size case. Our algorithm, SKC-BOUND, runs SKC-OF, using a capacity of $\delta$ (instead of $1$), i.e., it terminates once the total size of accepted items exceeds $\delta$.
Theorem G.4.
$\frac{\mathrm{ADAPT}}{\operatorname{\mathbb{E}}[\textbf{SKC-BOUND}]}\leq O\left(\frac{1}{\delta}\right)$
Proof.
If $w_{\max}\geq\Phi(\delta)/4$, SKC-OF selects only the item with $w_{\max}$ and we obtain that:
$\operatorname{\mathbb{E}}[\textbf{SKC-BOUND}]\geq\Phi(\delta)/4$. Otherwise: If all items have sizes $s_{i}^{j}<1-\delta$, then the policy never overflows (since it stops at total size $\delta$), achieving the same expected value as SKC-OF. However, if $s_{i}^{j}>1-\delta$ for some $i,j$, the process may terminate early, losing at most $\alpha w_{\max}$.
Hence:
$\operatorname{\mathbb{E}}[\textbf{SKC-BOUND}]$ $\geq\operatorname{\mathbb{E}}[\textbf{SKC-OF}]-\alpha w_{\max}\cdot\Pr[\exists\text{ item $i$ with }s_{i}^{j}>1-\delta]$
$\geq\Phi(\delta)/4-\alpha w_{\max}\cdot\sum_{i=1}^{n}\Pr[s_{i}^{j}>1-\delta]$
$\geq\Phi(\delta)/4-\alpha w_{\max}\cdot n\cdot\frac{1}{2n\alpha}$
$\geq\Phi(\delta)/4-\Phi(\delta)/8$
$=\Phi(\delta)/8.$
The theorem follows from Lemma 3.8,
$\frac{\mathrm{ADAPT}}{\operatorname{\mathbb{E}}[\textbf{SKC-BOUND}]}\leq\frac{\Phi(2)}{\Phi(\delta/8)}\leq\frac{\frac{16}{\delta}\Phi(\delta/8)}{\Phi(\delta/8)}=O\left(\frac{1}{\delta}\right)~.$
We note that SKC-BOUND determines a non-adaptive order of item choices (as in the other algorithms), but it uses an adaptive stopping rule, stopping once the total realized sizes first exceed $\delta$. In Section G.3, we show that even for bounded-sized items, any non-adaptive policy is bounded away by a factor $\Omega(\sqrt{\alpha})$ from the optimal policy (Theorem G.5).
G.3 $\Omega(\sqrt{\alpha})$ Adaptivity Gap of Bounded-Size Items
Theorem G.5.
There exists an instance $I$ of the MSKC-BOUND problem such that
$\frac{\mathrm{STOP\_ADAPT}(I)}{\mathrm{NON\_ADAPT}(I)}=\Omega\bigl(\sqrt{\alpha}\bigr).$
Proof.
Consider an item with the following parameters:
Property Item
Value $v$ 22
Cost $p$ $2-\epsilon^{2}$
[…]
We consider the case where $S^{\prime}<\frac{1}{2}$, and compute the expected utility per item.
$\operatorname{\mathbb{E}}[w|S^{\prime}<\frac{1}{2}]=v\cdot\Pr[s\leq S^{\prime}]-p=2\cdot(1-\epsilon)-(2-\epsilon^{2})=2-2\epsilon-2+\epsilon^{2}=-2\epsilon+\epsilon^{2}<0$
In expectation, the adaptive policy can include approximately $1/\epsilon$ items before reaching $S^{\prime}<1/2$, yielding expected total utility
$\frac{1}{\epsilon}\cdot\epsilon^{2}=\epsilon.$
For a non-adaptive policy, the number of attempted insertions is fixed in advance. If the policy inserts only one item, its expected utility is $v-p=\epsilon^{2}$. Let $\mathrm{utility}(i)$ denote the expected utility from the $i$th item. For $i>2$:
$\mathrm{utility}(i)$ $=$ $\operatorname{\mathbb{E}}[\mathrm{utility}(i)|s_{j}=\epsilon^{n}\ \forall j<i]\cdot\Pr[s_{j}=\epsilon^{n}\quad\forall j<i]$
$+\operatorname{\mathbb{E}}[\mathrm{utility}(i)|\exists!j<i:\ s_{j}>\epsilon^{n}]\cdot\Pr[\exists!j<i:\ s_{j}>\epsilon^{n}]$
$=$ $(v-p)\cdot(1-\epsilon)^{i-1}+((v-p)\cdot(1-\epsilon)-p\cdot\epsilon)\cdot\Pr[\exists!j<i:\ s_{j}>0]$
$=$ $\epsilon^{2}\cdot(1-\epsilon)^{i-1}+(\epsilon^{2}-2\epsilon)\cdot\Pr[\exists!j<i:\ s_{j}>0]$
$\leq$ $\epsilon^{2}\cdot(1-\epsilon)^{i-1}+(\epsilon^{2}-2\epsilon)\cdot\Pr[s_{i-1}>0\ \wedge\ s_{j}=0\quad\forall j<i-1]$
$=$ $\epsilon\cdot(1-\epsilon)^{i-2}(\epsilon\cdot(1-\epsilon)+\epsilon^{2}-2\epsilon)$
$\leq$ $\epsilon\cdot(1-\epsilon)^{i-2}(\epsilon^{2}-\epsilon)<0.$
In the above, the first inequality follows from the fact that $\epsilon^{2}-2\epsilon$ is negative, and that the probability that exactly one of the first i1i-1 items has size greater than $\epsilon^{n}$ is larger than the probability that only item i1i-1 has size greater than $\epsilon^{n}$. By Definition 2.1, the value of the IOR parameter is $\alpha=\frac{2-\epsilon^{2}}{\epsilon^{2}}=O(1/\epsilon^{2})$.
Therefore, we have that a non-adaptive policy only loses from having a sequence larger than 2, and the optimal non-adaptive policy has a utility of $2\epsilon^{2}$. We conclude that for this instance, we have
$\frac{\mathrm{STOP\_ADAPT}(I)}{\mathrm{NON\_ADAPT}(I)}=\frac{1}{2\epsilon}=O(\sqrt{\alpha}).$